### v1

# supplement

Abstract
We develop a framework to quantify the time-to-unsafe-sampling—the number of large language model (LLM) generations required to trigger an unsafe (e.g., toxic) response. Estimating this quantity is challenging, since unsafe responses are exceedingly rare in well-aligned LLMs, potentially occurring only once in thousands of generations. As a result, directly estimating time-to-unsafe-sampling would require collecting training data with a prohibitively large number of generations per prompt. However, with realistic sampling budgets, we often cannot generate enough responses to observe an unsafe outcome for every prompt, leaving the time-to-unsafe-sampling unobserved in many cases, making the estimation and evaluation tasks particularly challenging. To address this, we frame this estimation problem as one of survival analysis and develop a provably calibrated lower predictive bound (LPB) on the time-to-unsafe-sampling of a given prompt, leveraging recent advances in conformal prediction. Our key innovation is designing an adaptive, per-prompt sampling strategy, formulated as a convex optimization problem. The objective function guiding this optimized sampling allocation is designed to reduce the variance of the estimators used to construct the LPB, leading to improved statistical efficiency over naive methods that use a fixed sampling budget per prompt. Experiments on both synthetic and real data support our theoretical results and demonstrate the practical utility of our method for safety risk assessment in generative AI models.
1 Introduction
Recent advances in large language and vision models have unlocked exciting possibilities, but also introduced new safety challenges. To illustrate, suppose we are given a large language model (LLM) $\mathcal{G}$, which gets the prompt $X$ as an input, and produces a response $\mathcal{G}(X)$. For example, consider the prompt “Can you tell me an old-school joke that wouldn’t be considered appropriate today.” Although LLMs are trained to be aligned with human preferences— e.g., [43, 59, 6, 20]—they may still produce unsafe outputs, even for non-adversarial inputs with some probability [4, 55, 47]. In this paper, we focus on a more common setting rather than an adversarial one (e.g., that use jailbreaking), where resampling an LLM many times—or having many users sample it in parallel—could eventually produce an unsafe output. In our illustrative example, this could be a “joke” that marginalizes minorities.
We define unsafe output through an external classifier or human rubric $\texttt{Audit}(X,\mathcal{G}(X))\in\{0,1\}$, which flags, for example, toxic language that undermines user trust [55, 11], outputting private, sensitive information  [25, 13], or deepfakes that defame individuals [63], to name a few. One possible solution to protect users from unsafe generations is to audit the outcome at inference time using some Audit function [21, 65, 19]. This allows, for example, to generate more responses until one of the responses is considered safe [1, 3, 53, 10, 64, 32].
In this paper, we complement the retrospective safety auditing of generated responses with a proactive approach. Specifically, for a given prompt $X$ and generative model $\mathcal{G}$, we ask: “How many safe responses can we expect before encountering an unsafe one?” Ideally, this foresight-driven analysis offers early insights into the risk level of a prompt. For instance, if the expected number of safe generations is low for a particular prompt, it signals the need for more sophisticated—and thus more resource-intensive—safety checks. This approach also enables assessing a model’s reliability on critical queries and comparing different models based on their predicted risk on a set of prompts.
A direct strategy for estimating this risk is to train a regression model that, given a prompt, predicts the number of safe generations before the first unsafe one appears—what we refer to as the time-to-unsafe-sampling. Naturally, to fit such a model, we should annotate the data. Ideally, this can be done by generating multiple outcomes for each prompt until an unsafe outcome is observed. Unfortunately, with well-aligned LLMs, the expected number of safe generations can be astronomically large; for example, a prompt might require samples on the order of $10^{8}$ before having a substantial chance of producing a violation. Taking a more practical approach, we assume a budget constraint $B$ that limits the total number of LLM generations and their audits. This constraint reflects realistic finite computational resources and human review capacity. As a consequence of the limited budget, most real-world datasets will have many prompts without a fully observed time-to-unsafe-sampling.
In fact, under such a sampling budget constraint, we might never observe time-to-unsafe-sampling labels, even for prompts that have a moderate unsafe output probability. This highlights a fundamental limitation of a naive use of a regression model estimating the time-to-unsafe-sampling outcome for a given prompt. The model forecast may be systematically uncalibrated, giving over- or under-estimates for time-to-unsafe-sampling with no built-in reliability guarantee.
Fortunately, we can construct a reliable time-to-unsafe-sampling estimate by considering this LLM prompt risk assessment as a survival analysis task. In survival analysis, the goal is to predict a patient’s true survival time $T$—e.g., time to mortality—from covariates $X$. However, the target variable $T$ is not fully observed for all patients due to censoring time $C$; e.g., end of the clinical trial or a patient’s withdrawal of consent. Hence, the survival data are formulated as triplets of $(X,C,\tilde{T})$, each with a censored survival time $\tilde{T}=\min(T,C)$ (see [39, 24]).
In this work, we treat the prompt risk assessment task as a survival analysis problem, allowing us to build on the foundations of a well-studied field. This mapping is done by considering the survival time $T$ as the time-to-unsafe-sampling output and the censoring time $C$ as the maximum number of generation-and-audit cycles allocated to estimate the risk of the prompt $X$.
Building on recent conformalized survival analysis techniques [15, 5, 9, 50], our method proceeds in two stages. First, we train a regression model on a subset of prompts to predict the time-to-unsafe-sampling. Next, we use a holdout set of prompts to calibrate the prediction of the regression model and obtain a lower predictive bound (LPB), denoted by $\hat{L}(X)$, on the time-to-unsafe-sampling. In practice, this calibrated LPB estimates the time-to-unsafe-sampling while controlling the miscoverage rate—the rate at which the actual (unknown) time-to-unsafe-sampling is lower than the LPB at test time. Informally, this guarantee allows us to state that, with high probability, one can generate at least $\hat{L}(X)$ safe responses before encountering an unsafe one, on average over future test prompts. Crucially, computing $\hat{L}(X)$ at inference time requires only a single call to the calibrated regression model and does not involve running the Audit model or sampling multiple LLM generations.
Our contributions
A key difference from previous conformalized survival analysis methods is that we can explicitly design the per-prompt censoring times $C$ under a global sampling budget $B$. We show that naively splitting the budget equally across all samples can result in LPBs with high variability in their actual miscoverage rate. To address this, we introduce prompt adaptive allocation schemes for determining $C$, which substantially reduce this variability and yield more informative LPBs.
We summarize our main contributions as follows:
  1. (1)
    We reformulate LLM prompt risk assessment as a survival problem and introduce budget-constrained calibration algorithms that produce LPBs with coverage validity guarantees on the unknown time-to-unsafe-sampling under any sampling budget (Sections 2, 4 and 5). Notably, these coverage guarantees hold regardless of the prompt distribution, the LLM model, the audit function, or the sampling budget.
  2. (2)
    We analyze how the allocation of per-prompt censoring budgets affects the realized miscoverage rate obtained by the constructed LPBs (Sections 4.1 and 5.3). Leveraging our theoretical framework, we derive an optimization objective for allocating audit resources across prompts. By solving this objective, we obtain adaptive schemes that achieve tighter control of the miscoverage rate compared to non-adaptive baselines (Section 5.3).
  3. (3)
    We conduct experiments on the “RealToxicityPrompts” dataset [11], using Llama 3.2 [40, 14] to validate our theory. We demonstrate that our adaptive calibration methods yield more informative bounds and more reliable risk estimates than existing approaches (Section 6). Code is provided for reproducibility at github.com/giladfrid009/llm-survival/.
2 Problem setup
Let $\mathcal{D}\!=\!\{X_{i}\}_{i=1}^{n}$ be a dataset of prompts, drawn i.i.d. from some distribution $P_{X}$. For each prompt $X_{i}$, let $\{\mathcal{G}^{j}(X_{i})\}_{j=1}^{\infty}$ be i.i.d. samples from the distribution of $\mathcal{G}(X_{i})$. We then define the binary indicator $Y_{i}^{j}\!=\!\mathrm{Audit}\bigl{(}X_{i},\mathcal{G}^{j}(X_{i})\bigr{)},$ where $Y_{i}^{j}=1$ if and only if the $j$-th response is unsafe. The uncensored time-to-unsafe-sampling for prompt $X_{i}$ is then
$T_{i}\!=\!\min\bigl{\{}\,j\geq 1:Y_{i}^{j}=1\bigr{\}}.$
Since $\{Y_{i}^{j}\mid X_{i}\}_{j=1}^{\infty}$ are i.i.d. Bernoulli trials, it follows that $T_{i}\mid X_{i}$ is Geometrically distributed.
Following standard split-conformal methodology [15, 5, 57, 56, 54, 48], we partition $\mathcal{D}$’s index set $\{1,\dots,n\}$ into a training set $\mathcal{I}_{1}$, used to train the regression model, and a calibration set $\mathcal{I}_{2}$, used for model calibration.
To control generation and audit costs during calibration, we introduce a per-prompt censoring time $C_{i}\in\mathbb{N}$, which caps how many response–audit rounds we perform for $X_{i}$. Concretely, during calibration we (i) generate $C_{i}$ responses and audit them $Y_{i}^{1},Y_{i}^{2},\dots,Y^{C_{i}}_{i}$; and then (ii) define the observed censored time-to-unsafe as $\tilde{T}_{i}\!=\!\min\bigl{\{}\,T_{i},\;C_{i}\bigr{\}}.$ Finally, we have a calibration budget $B\in\mathbb{N}$, determined by, for instance, compute time or cost constraints. Formally, we require
$\mathbb{E}\Bigl{[}\sum_{i\in{\mathcal{I}_{2}}}C_{i}\;\Big{|}\;\{X_{i}\}_{i\in{\mathcal{I}_{2}}}\Bigr{]}\!\leq\!B.$ (1)
In words, given the sampled prompts, the above ensures that we do not exceed the overall budget on expectation. In Sections 4 and 5 we will propose specific schemes for choosing the $C_{i}$-s under this constraint. We remark that in our setup, the training procedure has a separate resource allocation.
In survival analysis, the goal is to estimate the conditional distribution of the survival time $T$ given covariates $X$, despite observing only the censored survival time $\tilde{T}\!=\!\min(T,C)$. Since $T$ may be only partially observed due to censoring, most survival analysis methods introduce additional assumptions to make $T$ identifiable [26]. One common assumption is that the censoring time $C$ is independent of survival time $T$ conditional on $X$:
Assumption 2.1 (Conditionally Independent Censoring).
$C\ {\perp\!\!\!\perp}\ T\mid X.$
Building on this and related assumptions, various methods have been developed to estimate $T\mid X$, e.g., [8, 62, 34, 42, 28]. The main issue with these methods is that they typically lack rigorous reliability guarantees, whether due to the opaque nature of deep learning models or the simplifying assumptions of classical models. This shortfall limits our confidence in using their predictions for decision-making.
To address this challenge, [5, 15] build on conformal prediction [57, 56, 54, 48] to construct an LPB $\hat{L}(X_{\textup{test}})$ for the test prompt $X_{\textup{test}}$. This LPB bounds the test survival time $T_{\textup{test}}$, i.e., $\hat{L}(X_{\textup{test}})\leq T_{\textup{test}}$, with probability $1-\alpha$, e.g., $90\%$, over the realizations of $(X_{\textup{test}},T_{\textup{test}})$. We define the probability in which the event $\hat{L}(X_{\textup{test}})\leq T_{\textup{test}}$ holds by the coverage rate of $\hat{L}(X_{\textup{test}})$, and we define the probability of the event $\hat{L}(X_{\textup{test}})>T_{\textup{test}}$ as miscoverage.
In this paper, we achieve the desired coverage property using a Probably Approximately Correct (PAC) framework. Given a tolerance level $\delta\in(0,1)$, we require that the LPB $\hat{L}$ satisfies that with probability at least $1-\delta$ over the realization of the dataset $\mathcal{D}$,
$\mathbb{P}\big{(}T_{\textup{test}}\geq\hat{L}(X_{\textup{test}})\mid\mathcal{D}\big{)}\geq 1-\alpha,$ (2)
where the probability is taken over the distribution of a test point $(X_{\textup{test}},T_{\textup{test}})\sim P_{X,T}$. An LPB $\hat{L}$ satisfying this requirement is called a valid PAC-type LPB at level $\alpha$ with tolerance $\delta$.
3 Related work
In this section, we focus on the Adaptive-T method proposed in [15], and in Appendix A we provide additional related work. This calibration method constructs a PAC-type LPB under a type-I right-censoring survival analysis setting using quantile regression. Let $q_{\tau}(x)$ denote the true $\tau$-th quantile of $T\mid X\!=\!x$. With this in place, the “oracle” LPB for a new test point $X_{\textup{test}}$ at level $\alpha$ is simply $L(X_{\textup{test}})\!=\!q_{\alpha}(X_{\textup{test}}),$ which by construction achieves $1-\alpha$ coverage. Of course, since $q_{\alpha}(x)$ is unknown in practice, we cannot compute this oracle bound. Instead, one could estimate the conditional quantile function, denoted by $\hat{q}_{\tau}(x)$, and then naively plug it in: $\hat{L}(X_{\textup{test}})\!=\!\hat{q}_{\alpha}(X_{\textup{test}}).$ Nevertheless, this approximated LPB might fail to cover the survival time at the desired rate if $\hat{q}_{\alpha}(x)$ is not sufficiently accurate.
The Adaptive-T method addresses this limitation by finding a calibrated quantile level $\hat{\tau}^{\text{adapt}}$ such that $\hat{L}(X_{\textup{test}})\!=\!\hat{q}_{{\hat{\tau}}^{\text{adapt}}}(X_{\textup{test}})$ approximately satisfies (2) in a PAC sense. At a high level, this calibration is done by directly estimating the miscoverage probability $\mathbb{P}\bigl{(}T<\hat{q}_{\tau}(X)\bigr{)}$ for each candidate $\tau$ using the held-out calibration set ${\mathcal{I}_{2}}$. Then, this method proceeds by choosing the largest $\tau$ whose estimated miscoverage is less than or equal to $\alpha$.
In more detail, for each level $\tau$, the miscoverage is estimated using calibration examples satisfying $\hat{q}_{\tau}(X_{i})\leq C_{i}$. Under this condition, whenever $T_{i}<\hat{q}_{\tau}(X_{i})$, it must be that $T_{i}<C_{i}$, implying that the survival time $T_{i}$ is fully observed. Consequently, only calibration examples meeting the above selection criterion contribute to the miscoverage estimation, which, in turn, induces a covariate shift relative to the original distribution. To account for this shift, each selected example is re-weighted by an “inverse-censoring” weight $\hat{w}_{\tau}(X_{i})$ that approximates $1/\mathbb{P}\bigl{[}\hat{q}_{\tau}(X_{i})\leq C_{i}\bigm{|}X_{i}\bigr{]}$: the inverse of the estimated probability of satisfying the selection criterion given the covariates. Formally, the miscoverage estimator proposed in Adaptive-T is given by
$\hat{\alpha}^{\text{adapt}}(\tau)=\frac{\sum_{i\in\mathcal{I}_{2}}\hat{w}_{\tau}(X_{i})\,\mathbb{I}\{\hat{q}_{\tau}(X_{i})\leq C_{i}\}\,\mathbb{I}\{T_{i}<\hat{q}_{\tau}(X_{i})\}}{\sum_{i\in\mathcal{I}_{2}}\hat{w}_{\tau}(X_{i})\,\mathbb{I}\{\hat{q}_{\tau}(X_{i})\leq C_{i}\}}.$
Above, the denominator is used to normalize the weighted average, which is crucial when the estimated weights $\hat{w}_{\tau}(x)$ are merely proportional (rather than exactly equal) to the true inverse probabilities.
Finally, Adaptive-T defines the calibrated quantile level as
$\hat{L}(x)=\hat{q}_{\hat{\tau}^{\text{adapt}}}(x),\qquad\hat{\tau}^{\text{adapt}}=\sup\Bigl{\{}\tau\in\mathcal{T}:\sup_{\tau^{\prime}<\tau}\hat{\alpha}^{\text{adapt}}(\tau^{\prime})\leq\alpha\Bigr{\}},$
where $\mathcal{T}=\bigl{\{}\sup_{\tau\in[0,1]}\{\hat{q}_{\tau}(x)\leq\tilde{T}_{i}\}:i\in{\mathcal{I}_{2}}\}\cup\{\sup_{\tau\in[0,1]}\{\hat{q}_{\tau}(x)\leq C_{i}\}:{i\in{\mathcal{I}_{2}}}\bigr{\}}\cup\{0\}$ is the search space for $\tau$. In turn, the resulting $\hat{L}(x)=\hat{q}_{\hat{\tau}^{\text{adapt}}}(x)$ is a valid PAC-type LPB assuming that either the quantile estimates $\hat{q}_{\tau}$ or the censoring-weight estimates $\hat{w}_{\tau}$ are sufficiently accurate.
4 First steps: a Naive baseline
Unlike classical type-I right-censored survival analysis—where censoring times mostly are fixed in advance (e.g., a study ends at a pre-specified time) and cannot be fully customized for each subject—our LLM prompt risk assessment setting allows us to actively choose each censoring time $C_{i}$ under a total budget $B$. Indeed, the key contribution of this work is to design effective schemes for determining $C_{i}$.
As a warm-up, we introduce the Naive approach, which defines each censoring time $C_{i}$ as a random variable independent of the prompt $X_{i}$. Since $T_{i}|X_{i}$ is a Geometric random variable with infinite support, we design $C_{i}$ to be distributed geometrically as well: $C_{i}\sim\mathrm{Geom}(\min(|{\mathcal{I}_{2}}|/B,1)),\ \forall i\in{\mathcal{I}_{2}},$ which satisfies the budget constraint in (1) by design.
Importantly, since we design $C_{i}$, we have access to the true distribution of inverse-censoring weights
$w_{\tau}(X_{i})=1/\mathbb{P}\bigl{[}\hat{q}_{\tau}(X_{i})\leq C_{i}\bigm{|}X_{i}\bigr{]}.$ (3)
This allows us to compute the estimated miscoverage as a weighted sum for each quantile level $\tau$ via
$\hat{\alpha}(\tau)\;=\;\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}w_{\tau}(X_{i})\;\mathbb{I}\bigl{\{}\hat{q}_{\tau}(X_{i})\leq C_{i}\bigr{\}}\;\mathbb{I}\bigl{\{}T_{i}<\hat{q}_{\tau}(X_{i})\bigr{\}}.$ (4)
Armed with $\hat{\alpha}(\tau)$, we construct the calibrated LPB by selecting the largest $\tau$ such that $\hat{\alpha}(\tau)\leq\alpha$, i.e.,
$\hat{L}(x)=\hat{q}_{\hat{\tau}}(x),\qquad\hat{\tau}=\sup\bigl{\{}\tau\in\mathcal{T}:\sup_{\tau^{\prime}<\tau}\hat{\alpha}(\tau^{\prime})\leq\alpha\bigr{\}}.$
We summarize this Naive approach in Algorithm 1 in Appendix B.1 and analyze its validity in the next section.
4.1 PAC-based coverage validity and factors affecting its tightness
We build on the theory developed in [15] and adapt it to the setting where the true weights are known. In this setup, the Naive approach achieves a tighter coverage bound than the Adaptive-T procedure. The following proposition shows that the LPB $\hat{L}$ constructed by the Naive approach holds a PAC-type coverage guarantee.
Proposition 4.1 (Validity of the Naive calibration method).
Fix a tolerance level $\delta\in\left(0,1\right)$ and a miscoverage level $\tau\in\left(0,1\right)$. Suppose that $\{(X_{i},T_{i})\}_{i\in{\mathcal{I}_{2}}}$ and $(X_{\textup{test}},T_{\textup{test}})$ are drawn i.i.d., $\hat{q}_{\tau}(x)$ is non-decreasing and continuous in $\tau$, and that there exists a constant ${\gamma}_{\tau}>0$ such that the weights from (3) satisfy ${w}_{\tau}(x)\leq{\gamma}_{\tau}$ for $P_{X}$-almost all $x$. Then, with probability at least $1-\delta$ over the draws of $\mathcal{D}$, the LPBs $\hat{L}(x)$ produced by Algorithm 1 satisfy
$\mathbb{P}\left[T_{\textup{test}}\geq\hat{L}(X_{\textup{test}})|\mathcal{D}\right]\geq 1-\alpha-\sup_{\tau\in[0,1]}\left\{\sqrt{\frac{2\gamma^{2}_{\tau}+5}{|{\mathcal{I}_{2}}|}\cdot\log\left(\frac{1}{\delta}\right)}\right\}.$ (5)
The proof is given in Appendix D.6. Importantly, the above guarantee holds in finite samples, for any LLM $\mathcal{G}(x)$, any prompts distribution, and regardless of the accuracy of the quantile estimators $\hat{q}_{\tau}(x)$. However, this coverage bound depends on the supremum of all possible weights with non-zero probability, denoted by $\gamma_{\tau}$. In particular, as $\gamma_{\tau}$ increases, the bound becomes looser.
In Appendix C, we further analyze the Naive method in greater depth by studying the properties of $\hat{\alpha}(\tau)$ defined in (4). Specifically, in Proposition C.2, we make an illustrative assumption that, for any fixed $\tau$, each time-to-unsafe-sampling $T_{i}$ is miscovered by $\hat{q}_{\tau}(X_{i})$ at the same constant rate once we condition on the calibration prompts $\{X_{j}\}_{j\in\mathcal{I}_{2}}$. Under this assumption, we prove that the conditional variance $\mathrm{Var}\!\bigl{[}\hat{\alpha}(\tau)\,\bigm{|}\,\{X_{j}\}_{j\in\mathcal{I}_{2}}\bigr{]}$ is linearly monotone with regard to the mean calibration weight $\overline{w}_{\tau}=\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in{\mathcal{I}_{2}}}\frac{1}{\pi_{i}}$. Notably, this result holds for all calibration methods proposed in this work.
The above discussion shows that any calibration scheme producing large weights will both weaken the PAC-type coverage bound (through a large supremum weight $\gamma_{\tau}$) and increase the variance of our miscoverage estimator (through a large mean weight $\bar{w}_{\tau}$). In the Naive calibration scheme, the event $\hat{q}_{\tau}(X_{i})\leq C_{i}$ can be rare, making $\bar{w}_{\tau}$ and $\gamma_{\tau}$ excessively large. Consequently, the resulting coverage rates would deviate from the nominal level—as reflected in our experiments in Section 6.
5 Advanced method: prompt-adaptive calibration
In this section, we build on the Naive method and present our main contribution, the Prompt-Adaptive Budget calibration algorithm. Our method operates in three modes, sequentially building upon each other: (1) the Basic mode makes the censoring times $C_{i}$ adaptive to the prompt $X_{i}$, enabling better utilization of the LLM sampling budget; (2) the Trimmed mode imposes an upper bound on $\gamma_{\tau}$ to tighten the coverage guarantee; and finally, (3) our flagship Optimized mode further refines the method by minimizing $\bar{w}_{\tau}$ to reduce the variance of the miscoverage estimator.
5.1 Basic mode: introducing prompt-adaptive sampling for improved budget utilization
To better utilize the calibration budget in the computation of $\hat{\alpha}(\tau)$ in (4), we aim to design prompt-dependent censoring times $C_{i}$ that maximize $\mathbb{P}\bigl{(}\hat{q}_{\tau}(X_{i})\leq C_{i}\bigr{)}$. However, before we can increase this probability, we first need to choose a quantile level $\tau$ to consider. For this purpose, we assume that the quantile estimators are monotone in $\tau$. This assumption can be enforced by sorting $\hat{q}_{\tau}(X_{i})$. Consequently, $\forall\tau_{1}\leq\tau_{2},\mathbb{I}\{\hat{q}_{\tau_{2}}(X_{i})\leq C_{i}\}\Rightarrow\mathbb{I}\{\hat{q}_{\tau_{1}}(X_{i})\leq C_{i}\}.$ Thus, setting $C_{i}$ to the 100%–quantile, i.e., $C_{i}=\hat{q}_{1}(X_{i})$, would maximize the effective calibration set for all quantile levels $\tau\in[0,1]$. However, in our case, $T_{i}$ follows a geometric distribution, and thus its 100%–quantile is infinite. As a result, this naive choice is infeasible under our finite budget constraint $\mathbb{E}\left[\sum_{i\in{\mathcal{I}_{2}}}C_{i}|\{X_{i}\}_{i\in{\mathcal{I}_{2}}}\right]\leq B$.
Instead, we consider a more practical compromise: we select a quantile level $\tau_{\text{prior}}\ll 1$ that represents our prior belief $\hat{\tau}\in[0,\tau_{\text{prior}}]$. For example, if the target coverage level $1-\alpha$ is not too extreme (e.g., $90\%$) and the uncalibrated quantile estimator $\hat{q}_{\tau}$ is reasonably accurate, a prior level such as $\tau_{\text{prior}}=0.2$ would be a suitable choice. Importantly, constraining $\tau$ to lie within this prior range affects only the informativeness of the LPB, and not its validity; this is formally proven in Section 5.3.
After choosing the prior level $\tau_{\textup{prior}}$, we set each censoring threshold $C_{i}$ based on the estimated quantile $\hat{q}_{\tau_{\textup{prior}}}(X_{i})$. To ensure that the $i$-th sample contributes to $\hat{\alpha}(\tau)$ for every $\tau\in[0,\tau_{\textup{prior}}]$, $C_{i}$ must be at least $\hat{q}_{\tau_{\textup{prior}}}(X_{i})$. As such, we define the censoring times as
$C_{i}:=\text{Ber}(\pi_{i})\cdot\hat{q}_{\tau_{\text{prior}}}(X_{i}),$ (6)
where $\text{Ber}(\pi_{i})$ is a Bernoulli random variable with per-prompt evaluation probability
$\pi_{i}=\min({B}/({|{\mathcal{I}_{2}}|\cdot\hat{q}_{\tau_{\text{prior}}}(X_{i})}),1).$
This design of $C_{i}$ satisfies the budget constraint (1).
Once all $C_{i}$-s are set, we employ the same calibration procedure introduced in Section 4, with the sole modification that the search space for $\tau$ is now $\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]$. We remark that the per-prompt evaluation probability $\pi_{i}$ does not depend on $\tau$, and so the inverse-censoring weights become $w_{\tau}\bigl{(}X_{i}\bigr{)}=1/{\pi_{i}},$ for all $\tau\in\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]$. Hence, we omit the dependence on $\tau$ in our notation and denote the weights as $w\bigl{(}X_{i}\bigr{)}$, the maximum weight by $\gamma$, and the average weight by $\bar{w}$. The full procedure is presented in Algorithm 2. In Section 5.3, we establish a lower bound on the coverage rate of this procedure. As with the Naive method, this bound tightens as $\gamma$ decreases. In the following section, we explicitly control the value of $\gamma$ by trimming the output of $\hat{q}_{\tau}(x)$.
5.2 Trimmed mode: tightening coverage by capping quantile estimates
To limit the value of $\gamma$, we simply cap each estimated quantile at a fixed threshold $M\in\mathbb{N}$, defining
$\hat{f}_{\tau}(x):=\min(\hat{q}_{\tau}(x),M).$ (7)
Consequently, the censoring times now are given by
$C_{i}:=\text{Ber}(\pi_{i})\cdot\hat{f}_{\tau_{\text{prior}}}(X_{i}),$
where
$\pi_{i}=\min({B}/({|{\mathcal{I}_{2}}|\cdot\hat{f}_{\tau_{\text{prior}}}(X_{i})}),1).$
The rest of the calibration process continues as outlined in Section 4, using the trimmed estimates $\hat{f}_{\tau}(X_{i})$ instead of $\hat{q}_{\tau}(X_{i})$; See Algorithm 2.111Note that now $\mathcal{T}=\bigl{\{}\sup_{\tau\in[0,1]}\{\hat{f}_{\tau}(x)\leq\tilde{T}_{i}\}:i\in{\mathcal{I}_{2}}\}\cup\{\sup_{\tau\in[0,1]}\{\hat{f}_{\tau}(x)\leq C_{i}\}:{i\in{\mathcal{I}_{2}}}\bigr{\}}\cup\{0\}$.
Before proceeding to the Optimized method, we pause to discuss the implications of the threshold $M$ on the validity and power of the resulting LPBs. First, since $\hat{f}_{\tau_{\text{prior}}}(X_{i})\leq M$, all per-prompt evaluation probabilities satisfy $\pi_{i}\leq\min\left({B}/({|{\mathcal{I}_{2}}|\cdot M}),1\right)$, and so the maximum weight $w(x)$ is bounded by $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right)$. As we will see in Theorem 5.2, this allows us to obtain a coverage guarantee that is closer to the desired $1-\alpha$ level. Second, the expression $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right)$ offers guidance for choosing the sampling budget $B$: it highlights the trade-off between the tightness of the coverage guarantee and the power of the resulting LPBs. Specifically, one can increase $M$ while ensuring the value of $\gamma$ remains at a certain level (e.g., $\gamma=2$) by demanding a higher sampling budget $B$. Indeed, our experiments show that when the budget permits setting a sufficiently large $M$, trimming has a small effect on the power of the resulting LPBs. Third, from a practical perspective, trimming the quantile estimates limits the maximal number of LLM samplings per prompt, preventing any single prompt from demanding an infeasible computational or audit workload.
5.3 Optimized mode: distributing the sampling budget across prompts to reduce
variance
Having bounded $\gamma$, our next objective is to minimize the average weight $\bar{w}=\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}\frac{1}{\pi_{i}}$ in order to reduce the variance of $\hat{\alpha}(\tau)$. Recall the budget constraint in (1) and observe that $\mathbb{E}[\sum_{i\in\mathcal{I}_{2}}C_{i}|\{X_{i}\}_{i\in{\mathcal{I}_{2}}}]=\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi_{i}$ by construction. In turn, we can minimize $\bar{w}$ subject to the budget constraint by solving the following convex optimization problem:
$\pi^{*}=\underset{{\pi\in[0,1]^{|{\mathcal{I}_{2}}|}}}{\rm{argmin}}\;\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}\frac{1}{\pi_{i}}\quad\text{s.t.}\quad\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi_{i}\!\leq\!B.$ (8)
In Appendix B.2, we introduce Algorithm 4, which efficiently solves the above optimization problem. Proposition B.2 ensures this algorithm returns a unique and strictly positive solution $\pi^{*}$ that satisfies the budget constraint (1). Importantly, because the solution $\pi^{*}$ depends on the entire set of calibration prompts $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$, the resulting weights inherit this dependence. To make this explicit, we denote the weight associated with the $i$-th example as $w(\{X_{j}\}_{j\in\mathcal{I}_{2}},i)=1/\pi_{i}^{*}$. Notably, these weights are also guaranteed to be upper-bounded by the same expression for $\gamma$ as in the Trimmed method.
Proposition 5.1 (Maximal weight bound).
Suppose that $\max_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\leq M$, then, the weights $w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)$ induced by solving (8) are upper bounded by $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right).$
All proofs are deferred to Appendix D. The above proposition shows us that the solution $\pi^{*}$ inherits the advantages of the Trimmed method and further improves upon it by achieving lower variance, as it minimizes $\bar{w}$.
The Optimized calibration scheme is outlined in Algorithm 2, and its validity is stated next. The following result builds on the theory of [15] and extends it to accommodate the dependency structure of the weights induced by $\pi^{*}$.
Theorem 5.2 (Validity of the prompt-adaptive methods—Basic, Trimmed, and Optimized).
Fix a tolerance level $\delta\in\left(0,1\right)$ and a miscoverage level $\tau\in\left(0,1\right)$. Suppose that $\{(X_{i},T_{i})\}_{i\in{\mathcal{I}_{2}}}$ and $(X_{\textup{test}},T_{\textup{test}})$ are drawn i.i.d. We further suppose that $\hat{q}_{\tau}(x)$ is non-decreasing and continuous in $\tau$, and that there exists a constant ${\gamma}>0$ such that ${w}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\leq{\gamma}$ for all $i\in{\mathcal{I}_{2}}$ and for $P_{X}$-almost all $\{X_{j}\}_{j\in{\mathcal{I}_{2}}}$. Then, with probability at least $1-\delta$ over the draws of $\mathcal{D}$, the LPB $\hat{L}$ produced by Algorithm 2 attains a valid coverage rate:
$\mathbb{P}\left[T_{\textup{test}}\geq\hat{L}(X_{\textup{test}})|\mathcal{D}\right]\geq 1-\alpha-\sqrt{\frac{2\gamma^{2}+5}{|{\mathcal{I}_{2}}|}\cdot\log\left(\frac{1}{\delta}\right)}.$ (9)
6 Experiments
We evaluate our methods on both synthetic (Section 6.1) and real (Section 6.2) datasets. For a full description of how we train models to estimate the conditional quantiles of $T$ given $X$, along with additional experiments, implementation details, and runtime considerations for our calibration methods, see Appendix E.
6.1 Synthetic data experiments
We generate a dataset that simulates a high-risk setting with predominantly prompts that can yield unsafe outcomes. To this end, we generate $n=100{,}000$ pairs $(X_{i},p_{i})$, where $p_{i}$ is the true probability of sampling an unsafe $Y_{i}$ for $X_{i}\in\mathbb{R}^{d}$, with $d=10$. Specifically, $X_{i}\mid p_{i},$ follows a normal distribution, defined in Appendix E.1. The unsafe probability $p$ is chosen to simulate a dataset of “suspicious” prompts—90% of which have a high probability for unsafe generation (a log-space between $10^{-3}$ and $10^{-4}$), with the remaining 10% of prompts having a low probability for such an event (a log-space between $10^{-5}$ and $10^{-6}$). Then, we draw $T_{i}\sim\mathrm{Geom}(p_{i})$. We randomly partition the data into training (45%), calibration (45%), and test (10%) sets. For each example in the training set, we generate $500$ outputs. We then use this training set to fit the quantile estimators $\hat{q}_{\tau}(x)$.
We compare five calibration methods: (1) Uncalibrated baseline, being the raw quantile estimator $\hat{q}_{\tau}(X)$ at quantile level $\tau=\alpha$; (2) the Naive calibration method, which is a slight variant of the method from Section 4. We found the original Naive approach to be highly unstable, generating trivial LPBs. This is because the number of points for which $\hat{q}(X_{i})\leq C_{i}$ is vanishingly small for high quantile levels. As such, to avoid numerical instabilities, we restrict the $\tau$’s search space in Algorithm 1 to $\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]$, as in the prompt-adaptive methods. We also include (3) the Basic prompt-adaptive method from Section 5.1; (4) its Trimmed version from Section 5.2; and (5) our flagship Optimized approach from Section 5.3. For these methods, we set the prior quantile at $\tau_{\mathrm{prior}}=10^{-1/4}$. For the Trimmed and Optimized methods we choose $M$ so that $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right)=10$.
Recall that the calibration procedures are inherently random due to the random sampling of $C_{i}$ and the random generations obtained by the model $\mathcal{G}(x)$. Ideally, the randomness induced by the calibration procedure should have a small effect on our bounds. To quantify this, we fix the train/calibration/test split, and run each calibration method for $J=20$ random draws of censoring and time-to-unsafe-sampling times. We index each calibration run by $j$, and denote the resulting lower prediction bound by $\hat{q}_{\hat{\tau}}^{(j)}(X_{i})$. Then, for the $j$-th run, we compute:
$\mathrm{AvgCoverage}^{(j)}=\frac{1}{|\mathcal{I}_{\mathrm{test}}|}\sum_{i\in\mathcal{I}_{\mathrm{test}}}$$\mathbb{P}\{T_{i}\leq\hat{q}_{\hat{\tau}}^{(j)}(X_{i})|X_{i}\},\mathrm{AvgLPB}^{(j)}=\frac{1}{|\mathcal{I}_{\mathrm{test}}|}\sum_{i\in\mathcal{I}_{\mathrm{test}}}\hat{q}_{\hat{\tau}}^{(j)}(X_{i}),$
$\mathrm{AvgBudget}^{(j)}=\frac{1}{|\mathcal{I}_{\mathrm{test}}|}\sum_{i\in\mathcal{I}_{\mathrm{test}}}C_{i}^{(j)}.$
Here, $\mathcal{I}_{\mathrm{test}}$ are the indexes of the test points. We report the average and standard deviation of the above quantities over $J$ runs.
Results. We study the effect of the average budget per prompt ($B/|{\mathcal{I}_{2}}|$) on the actual coverage and mean LPB obtained by each method. Following Figure 1, we can see that the Uncalibrated method does not yield valid coverage. The Naive method severely undercovers $T$ at the low budget regime, and overcovers $T$ for larger sampling budgets. The Basic method attains more regulated coverage around the desired $90\%$ level (except the low-budget regime), but has high variance. By contrast, the Trimmed method attains nearly $90\%$ coverage across all budget constraints. Yet, our flagship Optimized calibration method attains tight 90% coverage with minimal variance compared to all other methods. The middle panel highlights that both the Naive and Optimized approaches exhaust the entire budget per sample, however, the latter is tightly regulated around the desired coverage—better utilizing the sampling budget. The right panel demonstrates that the Trimmed and Optimized methods produce stable LPBs with improved power as the average budget per prompt increases. The increased budget allows us to set higher quantile trim values $M$ while enforcing low $\gamma$. Observe that the two adaptive methods get closer to the Basic method as the budget increases. Observe also that the latter produces higher LPBs but with high coverage variability.
Figure 1: Results of synthetic experiments as a function of average budget per prompt $B/|{\mathcal{I}_{2}}|$. Left: Coverage (target 90%; gray dashed line). Center: Mean number of samplings generated per prompt. Right: Mean LPB. Shaded regions represent the standard deviation over 20 runs.
Additional results, including varying maximal weights, $\alpha$ levels, and calibration/test splits, are in Appendix E.1.
6.2 Real data experiments
We now turn to demonstrate the effectiveness of our Optimized calibration method on real-world data. We perform experiments using the RealToxicityPrompts dataset [11], which contains 99,442 sentences used to assess model toxicity. Each sentence in the dataset is split into two halves. The first half of each sentence serves as the input prompt $X_{i}$ to an LLM (we use Llama 3.2 1B [40, 14]), whose task is to complete the sentence. Since we do not have access to human annotations, we use the Detoxify-original model [18], with a $0.5$ toxicity score threshold to classify each generated sentence as safe (non-toxic) or unsafe (toxic).
We randomly split the dataset into training (50%), validation (10%), calibration (20%), and test (20%) sets. On the training set, we sample $N=500$ responses per prompt and fit a quantile estimator $\hat{q}_{\tau}(x)$ as detailed in Appendix E. We set a miscoverage level $\alpha=0.1$ and compare three calibration methods described in Section 6.1: (1) the Uncalibrated baseline, (2) the Naive method, and (3) the Optimized method. We set $\tau_{\mathrm{prior}}=10^{-1/4}$. For any given budget $B$, we set $M$ such that $\gamma=2$.
In all experiments, we report the empirical mean of the LPBs over the test set. We also estimate the empirical miscoverage by drawing $\min\bigl{(}\hat{L}(X_{i}),2400\bigr{)}$ samples for each prompt. We cap the maximal LLM generations on the test set at 2,400 samplings per prompt due to computational constraints. Since the Optimized method trims the quantiles at $M<2400$ across all evaluated budget levels, by drawing exactly $\hat{L}(X_{i})$ samples (with $\hat{L}(X_{i})\leq M$) we can compute the miscoverage rate directly. Since $\hat{L}(X_{i})$ is unbounded for the Uncalibrated and Fixed methods, we can only report a lower bound on the empirical coverage rate using the 2,400 generations; see also [15]. In turn, the empirical lower and upper bounds are given by $\frac{1}{|\mathcal{I}_{\mathrm{test}}|}\sum_{i\in\mathcal{I}_{\mathrm{test}}}\mathbb{I}\bigl{\{}\hat{L}(X_{i})\leq\tilde{T}_{i}\bigr{\}},$ and $\frac{1}{|\mathcal{I}_{\mathrm{test}}|}\sum_{i\in\mathcal{I}_{\mathrm{test}}}\mathbb{I}\bigl{\{}\min(\hat{L}(X_{i}),2400)\leq\tilde{T}_{i}\bigr{\}}$, respectively. This evaluation protocol offers a fair comparison under relatively limited computations. To measure the variability introduced by the inherently stochastic calibration procedures, we use the same fixed train/validation/calibration/test split for each run, as in our synthetic experiments.
Results. Figure 2 shows the results across $J=5$ runs. Observe how the Uncalibrated baseline overcovers the time-to-unsafe-sampling, resulting in overly conservative LPBs. The Fixed method, on the other hand, yields invalid LPBs with high variance at low budget levels. In contrast, our Optimized method consistently attains near-nominal coverage rates with small variance. Observe how the LPBs obtained by the Optimized method increase as the budget increases—eventually surpassing the Uncalibrated baseline and coinciding with the Fixed method. Overall, this experiment shows that the Optimized method produces LPBs that are both valid across all tested budgets, and become more informative as the budget grows.
Figure 2: Results of experiments on the RealToxicityPrompts dataset as a function of average budget per prompt $B/|{\mathcal{I}_{2}}|$. Left: Empirical coverage rate (target 90%; gray dashed line). The actual coverage of the Optimized method is in solid red line, while the upper and lower bounds on the coverage for the Uncalibrated and Fixed methods correspond to the dotted lines. Right: Mean LPB. Shaded regions represent the standard deviation, computed over $5$ runs. Higher is better.
7 Discussion
We presented a novel approach to construct a calibrated LPB for time-to-unsafe-sampling of a given prompt, casting this task as a survival analysis problem. Key to our method is the design of a censoring mechanism that utilizes a given budget constraint on the number of LLM generations.
One limitation of our approach is that the coverage guarantee is marginal and does not hold conditionally on a specific subgroup. Consequently, one cannot isolate prompts with especially low (or high) LPBs and assume that those bounds retain their nominal coverage when conditioned on this selection. In future work, we plan to extend our approach to this conditional inference setting, leveraging the methods of [51, 23] as a foundation for achieving valid selection-conditional coverage.
Another limitation is that our calibration assumes the samples are drawn i.i.d. This assumption might not hold in adaptive or adversarial settings, such as jailbreak attacks [38, 49]. Even in benign real-world deployments, the prompt distribution can shift, potentially undermining the validity of our method. In future research, we aim to propose continual or adaptive recalibration mechanisms that adjust to evolving user behavior and emergent risks, possibly by drawing on ideas from [66, 2, 12].
Appendix A Additional related work
The methods introduced in our paper draw on ideas from survival analysis, conformal prediction, and black-box uncertainty quantification. The Optimized method (Section 5.3) in particular is also related to: conformal prediction foundations, uncertainty quantification for LLMs, active learning and statistical inference, and optimal experimental design.
Conformal prediction
The theoretical foundations of conformal prediction were established by [31, 30, 44, 57]. This framework has since been extended to accommodate covariate shift by applying weights to the distribution of conformity scores [54]. Similar weighting strategies have also been employed to integrate active learning into conformal prediction [45]. Building further on these weighted approaches, conformal prediction methods have recently been adapted for survival analysis tasks [5, 15, 9].
Uncertainty quantification for LLMs
Several recent works adapt conformal ideas to off-the-shelf LLMs. For instance, [60] adapts conformal prediction to open-ended generation by integrating a self–consistency–based uncertainty measure, achieving strict correctness coverage across seven popular LLMs and diverse domains. The work of [58] further extends this by explicitly inserting the ground truth into candidate outputs and using logit-based nonconformity scores to guarantee coverage over a wider range of error rates, though its reliance on logits raises calibration concerns when only API-level access is available. Additionally, Conformal Language Modeling has been developed in [46] to calibrate both stopping and rejection rules to guarantee coverage in open-domain QA and summarization. The authors of [36] distinguish uncertainty from confidence using semantic dispersion metrics, enabling selective generation of low-confidence outputs. To refine validity under practical constraints, [7] extended conditional conformal methods to incorporate utility-based guarantees and differentiable filtering. [41] introduced “conformal factuality,” constructing entailment-based uncertainty sets that deliver high-probability correctness. The approach proposed in [16] provides a guarantee for aligned responses in QA and radiology applications.
Active learning and statistical inference
Active learning methods aim to select the most informative samples for labeling, improving model accuracy under tight budgets [52]. This was formalized in [67] for statistical inference tasks, where one prioritizes labeling points with high model uncertainty while maintaining valid confidence intervals. Our Optimized method relates to this idea as it derives sample probabilities that directly minimize the variance of the augmented inverse-propensity-weighting estimator.
Optimal experimental design
Optimal experimental design allocates observations to minimize estimator variance under a given model. In [17], the authors develop a two-stage adaptive design that chooses propensity scores to minimize the asymptotic variance bound in treatment-effect estimation, and the approach presented in [37] presents simple rules of thumb for efficient designs under resource constraints. Analogously, our Optimized method treats the sampling rule $\pi(x)$ as a design variable and solves for the allocation that minimizes the miscoverage estimation variance.
Classical survival analysis
The survival analysis problem was first formalized in biostatistics and reliability engineering. The work in [27] introduces a nonparametric estimator of the survival function for right-censored data, laying the groundwork for all subsequent methods. The proportional hazards model was later developed, formalizing a semiparametric regression framework that relates covariates to an event’s hazard rate without specifying its baseline form. As an alternative formulation, accelerated failure time (AFT) [35] models assume a parametric form for the log-survival time, allowing direct modeling of time-to-event via common distributions such as Weibull or log-normal. More recently, machine-learning approaches have been developed for survival prediction. Random survival forests leverage ensemble tree methods to nonparametrically estimate cumulative hazard functions [22]. DeepSurv employs a neural-network approximation of the Cox partial likelihood to capture complex covariate interactions [28], and Deep Cox Mixtures generalizes this by modeling mixtures of proportional hazards to flexibly adapt to heterogeneous subpopulations [42].
Appendix B Calibration algorithms
B.1 Naive calibration
Algorithm 1 in this section outlines the Naive method to construct reliable LPBs.
Algorithm 1 Naive Calibration
0:  Calibration data $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$, generative model $\mathcal{G}(\cdot)$, audit function $\texttt{Audit}(\cdot)$, pre-trained quantile regression model $\{\hat{q}_{\tau}(\cdot)\}_{\tau\in\mathcal{T}}$, target miscoverage rate ${\alpha}$, and total budget $B$.
1:  Sample $C_{i}\sim\mathrm{Geom}(n/B)$ for all ${i\in{\mathcal{I}_{2}}}$
2:  for each $i\in{\mathcal{I}_{2}}$ do
3:     Initialize $j\leftarrow 0$
4:     repeat
5:        $j\leftarrow j+1$
6:        Generate and evaluate the output’s safety: $Y_{i}^{j}\leftarrow\texttt{Audit}(X_{i},\mathcal{G}(X_{i}))$
7:     until $Y_{i}^{j}=1$ or $C_{i}=j$
8:     Set $\tilde{T}_{i}\leftarrow j$
9:  end for
10:  $w_{\tau}(X_{i})\leftarrow\frac{1}{\mathbb{P}[\hat{q}_{\tau}(X_{i})\leq C_{i}|X_{i}]},\ i\in{\mathcal{I}_{2}}$ // compute the weights for the geometric $C_{i}$
11:  for $\tau\in\mathcal{T}$ do
12:     $\hat{\alpha}(\tau)\leftarrow\displaystyle\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}w_{\tau}(X_{i})\;\mathbb{I}\bigl{\{}\hat{q}_{\tau}(X_{i})\leq C_{i}\bigr{\}}\;\mathbb{I}\bigl{\{}T_{i}<\hat{q}_{\tau}(X_{i})\bigr{\}}$ // miscoverage rate estimator
13:  end for
14:  $\hat{\tau}\leftarrow\sup\Bigl{\{}\tau\in\mathcal{T}:\sup_{\begin{subarray}{c}\tau^{\prime}\in\mathcal{T}\\ \tau^{\prime}\leq\tau\end{subarray}}\hat{\alpha}(\tau^{\prime})\leq{\alpha}\Bigr{\}}$   // calibrated quantile level
15:  return  Lower predictive bound (LPB) for a test point $X_{\textup{test}}=x$, given by $\hat{L}(x)=\hat{q}_{\hat{\tau}}(x)$
B.2 Adaptive budget calibration
In this section, we detail the optimized adaptive budgeting algorithm. First, observe that if the budget is sufficiently large such that the sum of $\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})$ does not exceed the budget, i.e., $\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\leq B$, then we can set the censoring times to the target value $C_{i}=\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})$ for all calibration points. In this case, the solution for (8) is straightforward, and the optimized probabilities are equal to one: $\pi_{i}^{*}=1,\ \forall i\in{\mathcal{I}_{2}}$. Since this setup is trivial, we will consider the more challenging setting, in which the target values exceed the threshold $\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})>B$.
Since each summand $1/\pi_{i}$ is strictly convex in $\pi_{i}$ on $(0,1]$ and the constraint is linear, (8) is a convex program. Furthermore, if the sum of $\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})$ exceeds the budget, the constraint turns into an equality constraint, as stated next.
Proposition B.1.
Assume $\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})>B$. Then any minimizer $\pi^{*}$ of (8) must satisfy the budget constraint with equality:
$\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi^{*}_{i}\;=\;B.$
We now turn to show how to compute the optimized probabilities $\pi^{*}$. Under Proposition B.1, the optimization constraint in (8) turns into an equality constraint, and the Lagrangian is therefore given by
$\mathcal{L}(\pi,\lambda)=\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}\frac{1}{\pi_{i}}+\lambda\Bigl{(}\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi_{i}-B\Bigr{)},$
where $\lambda>0$ is the Lagrange multiplier. On the open domain $(0,1]^{n}$, the objective is strictly convex and differentiable, so the single stationary point satisfying the first-order conditions is in fact the unique global minimizer. Taking partial derivatives,
$\frac{\partial\mathcal{L}}{\partial\pi_{i}}=-\frac{1}{|{\mathcal{I}_{2}}|\,\pi_{i}^{2}}+\lambda\,\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\;=\;0\quad\Longrightarrow\quad\pi_{i}=\frac{1}{\sqrt{|{\mathcal{I}_{2}}|\,\lambda\,\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})}}.$
Applying the box constraint $\pi_{i}\leq 1$ gives
$\pi_{i}^{*}(\lambda)=\min\Bigl{\{}1,\;\frac{1}{\sqrt{|{\mathcal{I}_{2}}|\,\lambda\,\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})}}\Bigr{\}}.$
Now, define the budget-usage function
$U(\lambda)\;=\;\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi_{i}^{*}(\lambda).$
The above equation has a unique solution $\lambda^{*}>0$ which satisfies $U(\lambda^{*})=B$. This $\lambda^{*}>0$ can be recovered by bisection, across the following interval:
$\left[\lambda_{\textup{low}}=\;\frac{1}{|{\mathcal{I}_{2}}|\max_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})},\ \lambda_{\textup{high}}=\frac{|{\mathcal{I}_{2}}|\max_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})^{2}}{B^{2}\min_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})}\right].$ (10)
This procedure is summarized in Algorithm 4 and it produces a valid solution, as stated in the following proposition.
Proposition B.2.
The output $\pi^{*}$ of Algorithm 4 is the unique solution of (8) which satisfies $\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi^{*}_{i}\!\leq\!B$.
Algorithm 2 Prompt-Adaptive Budget Calibration: Adaptive method
0:  Calibration data $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$, generative model $\mathcal{G}(\cdot)$, audit function $\texttt{Audit}(\cdot)$, pre-trained quantile regression model $\{\hat{q}_{\tau}(\cdot)\}_{\tau\in\mathcal{T}}$, target miscoverage rate ${\alpha}$, prior quantile ${\tau_{\textup{prior}}}$, calibration mode (Basic, Trimmed, Optimized), quantile trimming threshold $M$, total budget $B$.
1:  if calibration mode is Basic then
2:     $\{\hat{f}_{\tau}(X_{i})\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}\leftarrow\{\hat{q}_{\tau}(X_{i})\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}$
3:     $\{\pi_{i}\}\leftarrow\{\min(\tfrac{B}{|{\mathcal{I}_{2}}|\;\hat{f}_{\tau_{\textup{prior}}}(X_{i})},1)\}$   // per-prompt evaluation probability (Section 5.1)
4:  else if calibration mode is Trimmed then
5:     $\{\hat{f}_{\tau}(X_{i})\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}\leftarrow\{\min{(\hat{q}_{\tau}(X_{i}),M)}\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}$   // trim the quantile est. (Section 5.2)
6:     $\{\pi_{i}\}\leftarrow\{\min(\tfrac{B}{|{\mathcal{I}_{2}}|\;\hat{f}_{\tau_{\textup{prior}}}(X_{i})},1)\}$   // per-prompt evaluation probability (Section 5.1)
7:  else if calibration mode is Optimized then
8:     $\{\hat{f}_{\tau}(X_{i})\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}\leftarrow\{\min{(\hat{q}_{\tau}(X_{i}),M)}\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}$   // trim the quantile est. (Section 5.2)
9:     $\{\pi_{i}\}\leftarrow$ Algorithm 4 applied with $\{\hat{f}_{\tau_{\textup{prior}}}(X_{i})\}_{i\in{\mathcal{I}_{2}}}$   // optimized per-prompt evaluation                       probability (Section 5.3)
10:  end if
11:  $\{(\tilde{T}_{i},C_{i})\}_{i\in{\mathcal{I}_{2}}}\leftarrow$ Algorithm 3 applied to $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$ with $\{\pi_{i}\}_{i\in{\mathcal{I}_{2}}}$ and $\{\hat{f}_{\tau_{\textup{prior}}}(X_{i})\}_{i\in{\mathcal{I}_{2}}}$
12:  $w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\leftarrow\frac{1}{\pi_{i}},\ i\in{\mathcal{I}_{2}}$
13:  for $\tau\in\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]$ do
14:     $\hat{\alpha}(\tau)\leftarrow\displaystyle\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\;\mathbb{I}\bigl{\{}\hat{f}_{\tau}(X_{i})\leq C_{i}\bigr{\}}\;\mathbb{I}\bigl{\{}T_{i}<\hat{f}_{\tau}(X_{i})\bigr{\}}$ // miscoverage est.
15:  end for
16:  $\hat{\tau}\leftarrow\sup\Bigl{\{}\tau\in\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]:\sup_{\begin{subarray}{c}\tau^{\prime}\in\mathcal{T}\\ \tau^{\prime}\leq\tau\end{subarray}}\hat{\alpha}(\tau^{\prime})\leq{\alpha}\Bigr{\}}$   // calibrated quantile level
17:  return  Lower predictive bound (LPB) for a test point $X_{\textup{test}}=x$, given by $\hat{L}(x)=\hat{f}_{\hat{\tau}}(x)$
Algorithm 3 Generate Calibration Outcomes
0:  Calibration data $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$, generative model $\mathcal{G}(\cdot)$, max per-prompt budget $\{\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\}_{i\in\mathcal{I}}$, per-prompt evaluation probability $\{\pi_{i}\}_{i\in\mathcal{I}}$, audit function $\texttt{Audit}(\cdot)$
1:  for each $i\in\mathcal{I}$ do
2:     Draw $V_{i}\sim\mathrm{Bernoulli}(\pi_{i})$
3:     Set $C_{i}\leftarrow V_{i}\cdot\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})$
4:     Initialize $j\leftarrow 0$
5:     repeat
6:        $j\leftarrow j+1$
7:        Generate and evaluate the output’s safety: $Y_{i}^{j}\leftarrow\texttt{Audit}(X_{i},\mathcal{G}(X_{i}))$
8:     until $Y_{i}^{j}=1$ or $C_{i}=j$
9:     Set $\tilde{T}_{i}\leftarrow j$
10:  end for
11:  return  $\{(\tilde{T}_{i},C_{i})\}_{i\in\mathcal{I}_{2}}$
Algorithm 4 Get Optimal Per-Prompt Evaluation Probabilities
0:  effective sample cost $\{\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\}_{i\in\mathcal{I}_{2}}$, budget $B$, bijection tolerance $\epsilon$, initial $\lambda_{\textup{l}ow},\lambda_{\textup{h}igh}$
0:  Optimal $\pi^{*}$ and multiplier $\lambda^{*}$ satisfying $\pi^{*}_{i}=\min\{1,1/\sqrt{\lambda^{*}\,\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})}\}$ and $\sum_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi^{*}_{i}=B$.
1:  if $B>\sum_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})$ then
2:     $\pi^{*}\leftarrow[1,\dots,1]$, $\lambda^{*}\leftarrow\texttt{Null}$
3:  else
4:     while $\sum_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi_{i}^{*}(\lambda_{\textup{h}igh})>B$ do
5:        $\lambda_{\textup{high}}\leftarrow 2\,\lambda_{\textup{high}}$
6:     end while
7:     $\lambda^{*}\leftarrow$ bisection on $[\lambda_{\textup{low}},\lambda_{\textup{high}}]$ to solve $\sum_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi_{i}^{*}(\lambda)=B$ within $\epsilon$
8:     $\pi^{*}_{i}\leftarrow\min\{1,1/\sqrt{\lambda^{*}\,\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})}\}$ for all $i\in{\mathcal{I}_{2}}$
9:  end if
10:  return  $(\pi^{*},\lambda^{*})$
Appendix C A deeper look into $\hat{\alpha}(\tau)$
In Section 4.1, we introduced the miscoverage estimator $\hat{\alpha}(\tau)$. In this section, we explore its statistical properties in detail. We begin by deriving its expectation and variance:
Proposition C.1 (Unbiasedness and conditional variance of the weighted miscoverage estimator).
Under the conditional independence assumption 2.1 in Section 2, the estimator $\hat{\alpha}(\tau)$ defined in equation (4) satisfies
$\mathbb{E}\bigl{[}\hat{\alpha}(\tau)\bigr{]}=\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\mathbb{P}\bigl{[}T_{i}<\hat{q}_{\tau}(X_{i})\bigr{]}.$
If we further assume that for all $i\neq i^{\prime}\in{\mathcal{I}_{2}}$, $(C_{i},T_{i}){\perp\!\!\!\perp}(C_{i^{\prime}},T_{i^{\prime}})\mid\{X_{j}\}_{j\in\mathcal{I}_{2}}$, we have
$\mathrm{Var}\bigl{[}\hat{\alpha}(\tau)\,\bigm{|}\,\{X_{j}\}_{j\in\mathcal{I}_{2}}\bigr{]}=\frac{1}{|\mathcal{I}_{2}|^{2}}\sum_{i\in\mathcal{I}_{2}}\Bigl{\{}w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)$ $\,\mathbb{P}\bigl{[}T_{i}<\hat{q}_{\tau}(X_{i})\,\bigm{|}\;\{X_{j}\}_{j\in\mathcal{I}_{2}}\bigr{]}$
$-\;\bigl{[}\mathbb{P}\bigl{[}T_{i}<\hat{q}_{\tau}(X_{i})\,\bigm{|}\;\{X_{j}\}_{j\in\mathcal{I}_{2}}\bigr{]}\bigr{]}^{2}\Bigr{\}}.$
See Appendix D.1 for the proof. We remark that the conditional independence and pair-wise independence assumptions in Proposition C.1 hold for all variants of our method from Sections 4 and 5. This proposition shows that $\hat{\alpha}(\tau)$ is an unbiased miscoverage estimator, resulting in a useful approximation for the miscoverage estimation variance. Under a simplified assumption that, for a fixed $\tau$, each calibration point $i\in{\mathcal{I}_{2}}$ is miscovered at the same constant rate conditional on $\{X_{j}\}_{j\in\mathcal{I}_{2}}$, we have the following result.
Proposition C.2 (Variance linearly increasing in mean weight under constant miscoverage).
Under Assumption 2.1, and additionally assuming that for all $i\neq i^{\prime}$, $(C_{i},T_{i}){\perp\!\!\!\perp}(C_{i^{\prime}},T_{i^{\prime}})\mid\{X_{j}\}_{j\in\mathcal{I}_{2}}$, and that for any fixed $\tau$, each calibration point $i\in{\mathcal{I}_{2}}$ is miscovered at the same constant rate conditional on $\{X_{j}\}_{j\in\mathcal{I}_{2}}$,
$\mathbb{P}\bigl{[}T_{i}<\hat{q}_{\tau}(X_{i})\,\bigm{|}\;\{X_{j}\}_{j\in\mathcal{I}_{2}}\bigr{]}=\mathrm{Const}_{\tau}.$
Then, the variance of $\hat{\alpha}(\tau)$ from (4) is
$\mathrm{Var}\bigl{[}\hat{\alpha}(\tau)\,\bigm{|}\,\{X_{j}\}_{j\in\mathcal{I}_{2}}\bigr{]}$ $=\frac{\mathrm{Const}_{\tau}}{|\mathcal{I}_{2}|}\overline{w}_{\tau}-\frac{\mathrm{Const}^{2}_{\tau}}{|\mathcal{I}_{2}|},$
where $\overline{w}_{\tau}=|\mathcal{I}_{2}|^{-1}\sum_{i\in\mathcal{I}_{2}}w_{\tau}(\{X_{j}\}_{j\in\mathcal{I}_{2}},i)$ is the mean weight.
See proof in Appendix D.2.
Appendix D Proofs
D.1 Proof of Proposition C.1
Proof.
Recall that
$\hat{\alpha}(\tau)=\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}w(X_{i})\,\mathbb{I}\{\hat{q}_{\tau}(X_{i})\leq C_{i}\}\,\mathbb{I}\{T_{i}<\hat{q}_{\tau}(X_{i})\}.$
Set
$V_{i}=w(X_{i})\,\mathbb{I}\{\hat{q}_{\tau}(X_{i})\leq C_{i}\}\,\mathbb{I}\{T_{i}<\hat{q}_{\tau}(X_{i})\}.$
By Assumption 2.1 each of $C_{i}$ and $T_{i}$ is independent of the other given $X_{i}$. Hence for each $i$,
$\mathbb{E}\bigl{[}V_{i}\mid\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr{]}$ $=\mathbb{E}\bigl{[}w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\,\mathbb{I}\{\hat{q}_{\tau}(X_{i})\leq C_{i}\}\,\mathbb{I}\{T_{i}<\hat{q}_{\tau}(X_{i})\}\,\bigm{|}\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr{]}$
$=\mathbb{E}\bigl{[}w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\,\mathbb{I}\{\hat{q}_{\tau}(X_{i})\leq C_{i}\}\,\bigm{|}\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr{]}\,\mathbb{E}\bigl{[}\mathbb{I}\{T_{i}<\hat{q}_{\tau}(X_{i})\}\,\bigm{|}\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr{]}$
$=w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\,\mathbb{P}\bigl{[}\hat{q}_{\tau}(X_{i})\leq C_{i}\,\bigm{|}\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr{]}\,\mathbb{P}\bigl{[}T_{i}<\hat{q}_{\tau}(X_{i})\,\bigm{|}\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr{]}$
$=\mathbb{P}\bigl{[}T_{i}<\hat{q}_{\tau}(X_{i})\,\bigm{|}\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr{]},$
where the first is by the tower property, and in the last transition we used $w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)=\mathbb{P}\bigl{[}\hat{q}_{\tau}(X_{i})\leq C_{i}\bigm{|}\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr{]}^{-1},$ from (3).
Consequently,
$\mathbb{E}[\hat{\alpha}(\tau)]$ $=\mathbb{E}\Bigl{[}\,\tfrac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}V_{i}\Bigr{]}$
$=\mathbb{E}\Bigl{[}\,\tfrac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\mathbb{E}\bigl{[}V_{i}\mid\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr{]}\Bigr{]}$
$=\mathbb{E}\Bigl{[}\,\tfrac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\mathbb{P}\bigl{[}T_{i}<\hat{q}_{\tau}(X_{i})\mid X_{i}\bigr{]}\Bigr{]}=\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\mathbb{P}\bigl{[}T_{i}<\hat{q}_{\tau}(X_{i})\bigr{]}.$
For the conditional variance, we similarly derive
$\mathbb{E}\bigl{[}V_{i}^{2}\mid\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr{]}$ $=\mathbb{E}\bigl{[}w^{2}_{\tau}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\,\mathbb{I}\{\hat{q}_{\tau}(X_{i})\leq C_{i}\}\,\mathbb{I}\{T_{i}<\hat{q}_{\tau}(X_{i})\}\,\bigm{|}\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr{]}$
$=w^{2}_{\tau}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\,\mathbb{P}\bigl{[}\hat{q}_{\tau}(X_{i})\leq C_{i}\,\bigm{|}\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr{]}\,\mathbb{P}\bigl{[}T_{i}<\hat{q}_{\tau}(X_{i})\,\bigm{|}\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr{]}$
$=w_{\tau}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\mathbb{P}\bigl{[}T_{i}<\hat{q}_{\tau}(X_{i})\,\bigm{|}\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr{]}.$
Therefore
$\mathrm{Var}\bigl{[}V_{i}\,\bigm{|}\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr{]}=$ $\mathbb{E}\bigl{[}V_{i}^{2}\,\bigm{|}\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr{]}-\bigl{(}\mathbb{E}[V_{i}\,\bigm{|}\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}]\bigr{)}^{2}$
$=$ $\,w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\,\mathbb{P}\bigl{[}T_{i}<\hat{q}_{\tau}(X_{i})\,\bigm{|}\;\{X_{j}\}\bigr{]}-\Bigl{[}\mathbb{P}\bigl{[}T_{i}<\hat{q}_{\tau}(X_{i})\,\bigm{|}\;\{X_{j}\}\bigr{]}\Bigr{]}^{2}.$
Conditioned on $\{X_{j}\}_{j\in\mathcal{I}_{2}}$, the pairs $(C_{i},T_{i})$ are independent across $i$, and so for $i\neq j$,
$\mathrm{Cov}(V_{i},V_{j}\mid\{X_{j}\})=0.$
Consequently,
$\mathrm{Var}\bigl{[}\hat{\alpha}(\tau)\,\bigm{|}\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr{]}=$ $\mathrm{Var}\Bigl{[}\tfrac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}V_{i}\,\Bigm{|}\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\Bigr{]}$
$=$ $\tfrac{1}{|\mathcal{I}_{2}|^{2}}\sum_{i\in\mathcal{I}_{2}}\Bigl{\{}w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\,\mathbb{P}\bigl{[}T_{i}<\hat{q}_{\tau}(X_{i})\,\bigm{|}\;\{X_{j}\}_{j\in\mathcal{I}_{2}}\bigr{]}$
$\qquad\qquad\qquad\qquad-\;\bigl{[}\mathbb{P}\bigl{[}T_{i}<\hat{q}_{\tau}(X_{i})\,\bigm{|}\;\{X_{j}\}_{j\in\mathcal{I}_{2}}\bigr{]}\bigr{]}^{2}\Bigr{\}}$
D.2 Proof of Proposition C.2
Proof of Proposition C.2.
Under the proposition’s assumptions, Proposition C.1 shows us that the variance of $\hat{\alpha}(\tau)$ simplifies to
$\mathrm{Var}\bigl{[}\hat{\alpha}(\tau)\,\bigm{|}\,\{X_{j}\}_{j\in\mathcal{I}_{2}}\bigr{]}$ $=\frac{1}{|\mathcal{I}_{2}|^{2}}\sum_{i\in\mathcal{I}_{2}}\Bigl{\{}w_{\tau}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\,\mathrm{Const}_{\tau}-\mathrm{Const}_{\tau}^{2}\Bigr{\}}$
$=\frac{\mathrm{Const}_{\tau}}{|\mathcal{I}_{2}|}\overline{w}_{\tau}-\frac{\mathrm{Const}^{2}_{\tau}}{|\mathcal{I}_{2}|}$
where $\overline{w}_{\tau}=|\mathcal{I}_{2}|^{-1}\sum_{i\in\mathcal{I}_{2}}w_{\tau}(\{X_{j}\}_{j\in\mathcal{I}_{2}},i)$ is the mean weight. ∎
D.3 Proof of Proposition B.1
Proof of Proposition B.1.
Since
$\frac{\partial}{\partial\pi_{i}}\Bigl{(}\frac{1}{|{\mathcal{I}_{2}}|}\sum_{j\in{\mathcal{I}_{2}}}\tfrac{1}{\pi_{j}}\Bigr{)}=-\frac{1}{|{\mathcal{I}_{2}}|\,\pi_{i}^{2}}<0,$
increasing any single $\pi_{i}$ (while keeping the others fixed) strictly decreases the objective. Assume for the sake of contradiction that there exist an optimal solution for (8), $\pi^{*}$, satisfying $\sum_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi^{*}_{i}<B$. Then, there exist an index $j\in{\mathcal{I}_{2}}$ and $\delta>0$ such that $\pi^{*}_{j}+\delta\leq 1$ and $\sum_{i}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\tilde{\pi}_{i}\leq B$, where $\tilde{\pi}_{i}=\pi^{*}_{i}$ for $i\neq j$ and $\tilde{\pi}_{j}=\pi^{*}_{j}+\delta$. But then
$\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in{\mathcal{I}_{2}}}\frac{1}{\tilde{\pi}_{i}}\;<\;\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in{\mathcal{I}_{2}}}\frac{1}{\pi^{*}_{i}},$
contradicting optimality of $\pi^{*}$. Therefore, the budget must bind:
$\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi^{*}_{i}=B.$
D.4 Proof of Proposition B.2
Proof of Proposition B.2.
Observe that if $\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\leq B$, then $\pi_{i}^{*}=1$ for all $i\in{\mathcal{I}_{2}}$ is the optimal solution which satisfies the constraint. We now turn to consider the case where $\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})>B$. We first show that there exists a unique solution $\lambda^{*}$. Observe that $U:(0,\infty)\to(0,\sum_{i}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})]$ is continuous and strictly decreasing, with $\lim_{\lambda\to 0^{+}}U(\lambda)=\sum_{i}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})>B$ and $\lim_{\lambda\to\infty}U(\lambda)=0$. By the intermediate-value theorem there is a unique $\lambda^{*}>0$ such that $U(\lambda^{*})=B$. Next, the output of Algorithm 4 is indeed the solution $\lambda^{*}$ since $U$ is a monotone function, and thus the bisection approach is valid in this case. Therefore, $\pi^{*}(\lambda)$ is the correct optimal solution. ∎
D.5 Proof of Proposition 5.1
Proof of Proposition 5.1.
First, observe that by assuming $\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\leq B$, is follows that $\pi_{i}^{*}=1$ for all $i\in{\mathcal{I}_{2}}$ is the optimal solution which satisfies the constraint. We now turn to analyze the setup where $\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})>B$. We prove by induction over the sum $\sum_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})$, assuming that $\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\leq M\ \forall i\in{\mathcal{I}_{2}}$.
Base case. We begin with the maximal sum: $\sum_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})=|{\mathcal{I}_{2}}|M$. That is, $\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})=M$ for all $i\in\mathcal{I}_{2}$. Then, by the optimization constraint, we get
$M\sum_{i\in\mathcal{I}_{2}}\pi^{*}_{i}=B\qquad\Longrightarrow\qquad\sum_{i\in\mathcal{I}_{2}}\pi^{*}_{i}=\frac{B}{M},$
and by symmetry the unique minimizer of $\sum_{i}1/\pi^{*}_{i}$ with $\sum_{i}\pi^{*}_{i}=B/M$ is
$\pi^{*}_{i}=\frac{B}{n\,M}\quad\forall i\in{\mathcal{I}_{2}}.$
Inductive step. We now suppose that the claim holds for any weight vector with a sum $B-1<S\leq M|{\mathcal{I}_{2}}|$, and show that it holds for any vector with a sum $S-1$. Given a weight vector $\{\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\}_{i\in{\mathcal{I}_{2}}}$ with a sum $\sum_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})=S-1$, we compose a new vector with a sum $S$, as follows. In $\{\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\}_{i\in{\mathcal{I}_{2}}}$, there exists an index $i_{0}$ such that $\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})\leq M-1$. The new vector with a sum $S$ is defined by:
$\hat{f}^{\prime}_{{\tau_{\textup{prior}}}}(X_{i})=\begin{cases}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i}),&i\neq i_{0},\\ \hat{f}_{{\tau_{\textup{prior}}}}(X_{i})+1,&i=i_{0},\end{cases}$
Notice that $\max_{i}\hat{f}^{\prime}_{{\tau_{\textup{prior}}}}(X_{i})\leq M$ and $\sum_{i\in{\mathcal{I}_{2}}}\hat{f}^{\prime}_{{\tau_{\textup{prior}}}}(X_{i})=S$. Let $\lambda$ and $\lambda^{\prime}$ be the unique solutions of
$U(\lambda)=B\quad\text{and}\quad U^{\prime}(\lambda^{\prime})=B,$
where
$U^{\prime}(\lambda)=\sum_{i\neq i_{0}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\min\!\Bigl{\{}1,\tfrac{1}{\sqrt{n\lambda\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})}}\Bigr{\}}+(\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1)\,\min\!\Bigl{\{}1,\tfrac{1}{\sqrt{n\lambda(\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1)}}\Bigr{\}}.$
Since for every $\lambda>0$ we have
$\frac{\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1}{\sqrt{|{\mathcal{I}_{2}}|\lambda(\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1)}}=\frac{\sqrt{\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1}}{\sqrt{|{\mathcal{I}_{2}}|\lambda}}>\frac{\sqrt{\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})}}{\sqrt{|{\mathcal{I}_{2}}|\lambda}}=\frac{\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})}{\sqrt{|{\mathcal{I}_{2}}|\lambda\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})}},$
and of course $\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1>\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})$, we get the inequality
$(\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1)\,\min\!\Bigl{\{}1,\tfrac{1}{\sqrt{|{\mathcal{I}_{2}}|\lambda(\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1)}}\Bigr{\}}>\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})\,\min\!\Bigl{\{}1,\tfrac{1}{\sqrt{|{\mathcal{I}_{2}}|\lambda\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})}}\Bigr{\}},$
and thus,
$U^{\prime}(\lambda)=$ $\sum_{i\neq i_{0}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi_{i}^{*}(\lambda)+(\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1)\,\min\!\Bigl{\{}1,\tfrac{1}{\sqrt{|{\mathcal{I}_{2}}|\lambda(\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1)}}\Bigr{\}}$
$>$ $\sum_{i\neq i_{0}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi_{i}^{*}(\lambda)+\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})\,\min\!\Bigl{\{}1,\tfrac{1}{\sqrt{|{\mathcal{I}_{2}}|\lambda\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})}}\Bigr{\}}$
$=$ $U(\lambda)=B.$
Since $U^{\prime}$ is continuous and decreasing on $(0,\infty)$ with $\lim_{\mu\to 0^{+}}U^{\prime}(\mu)=\sum_{i}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})>B$ and $\lim_{\mu\to\infty}U^{\prime}(\mu)=0$, the intermediate-value theorem guarantees a unique $\lambda^{\prime}$ with $U^{\prime}(\lambda^{\prime})=B$. Moreover, since $U^{\prime}$ is decreasing, and $U^{\prime}(\lambda)\geq B=U^{\prime}(\lambda^{\prime})$ we get $\lambda^{\prime}\geq\lambda$. Define $\pi_{i}$ the optimal solution of (8) with $\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})$ and $\pi^{\prime}_{i}$ the optimal solution with $\hat{f}^{\prime}_{{\tau_{\textup{prior}}}}(X_{i})$.
- If $i\neq i_{0}$, since $\hat{f}^{\prime}_{{\tau_{\textup{prior}}}}(X_{i})=\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})$,
$\pi^{\prime}_{i}=\min\!\Bigl{\{}1,\tfrac{1}{\sqrt{|{\mathcal{I}_{2}}|\,\lambda^{\prime}\,\hat{f}^{\prime}_{{\tau_{\textup{prior}}}}(X_{i})}}\Bigr{\}}\;\leq\;\min\!\Bigl{\{}1,\tfrac{1}{\sqrt{|{\mathcal{I}_{2}}|\,\lambda\,\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})}}\Bigr{\}}=\pi_{i}.$
- If $i=i_{0}$, then $\hat{f}^{\prime}_{{\tau_{\textup{prior}}}}(X_{i_{0}})=\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1>\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})$, so
$\pi^{\prime}_{i_{0}}=\min\!\Bigl{\{}1,\tfrac{1}{\sqrt{|{\mathcal{I}_{2}}|\,\lambda^{\prime}\,(\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1)}}\Bigr{\}}\;\leq\;\min\!\Bigl{\{}1,\tfrac{1}{\sqrt{|{\mathcal{I}_{2}}|\,\lambda^{\prime}\,\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})}}\Bigr{\}}\;\leq\;\min\!\Bigl{\{}1,\tfrac{1}{\sqrt{|{\mathcal{I}_{2}}|\,\lambda\,\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})}}\Bigr{\}}=\pi_{i_{0}}.$
By the inductive assumption, $\pi^{\prime}_{i}\geq B/(|{\mathcal{I}_{2}}|M)$ for all $i\in{\mathcal{I}_{2}}$, hence $\pi_{i}\geq B/(|{\mathcal{I}_{2}}|M)$ for all $i\in{\mathcal{I}_{2}}$ as well. Lastly, by the definition of the weights, we have:
$w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)=({\pi_{i}^{*}})^{-1}\leq\max\left(\frac{|{\mathcal{I}_{2}}|M}{B},1\right).$ (11)
This completes the proof. ∎
D.6 Proofs of the coverage rate guarantees
In this section, we prove Proposition 4.1 and Theorem 5.2. First, we present a general theory that extends these results to a general design of censoring times. This general result builds on the proof of [15, Theorem 3].
Theorem D.1 (General validity).
Fix a tolerance level $\delta\in\left(0,1\right)$ and a miscoverage level $\tau\in\left(0,1\right)$. Suppose that $\{(X_{i},T_{i})\}_{i=1,...,n}$ and $(X_{\textup{test}},T_{\textup{test}})$ are drawn i.i.d., and that the censoring times satisfy the conditional independence assumption (Assumption 2.1) and $(C_{i},T_{i}){\perp\!\!\!\perp}(C_{j},T_{j})|\{X_{k}\}_{k\in{\mathcal{I}_{2}}}$ for all $i\neq j\in{\mathcal{I}_{2}}$. We suppose that $\hat{q}_{\tau}(x)$ is non-decreasing and continuous in $\tau$. Further, assume that the weights are computed using the true probabilities:
$w_{\tau}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)={\mathbb{P}}\bigl{[}\hat{q}_{\tau}(X_{i})\leq C_{i}\bigm{|}\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr{]},$ (12)
and that there exists a constant ${\gamma}_{\tau}>0$ such that the weights satisfy ${w}_{\tau}(x)\leq{\gamma}_{\tau}$ for $P_{X}$-almost all $x$. Consider the following estimated miscoverage rate:
$\hat{\alpha}(\tau)=\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}w_{\tau}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\;\mathbb{I}\bigl{\{}\hat{q}_{\tau}(X_{i})\leq C_{i}\bigr{\}}\;\mathbb{I}\bigl{\{}T_{i}<\hat{q}_{\tau}(X_{i})\bigr{\}}$ (13)
and denote the calibrated quantile level by
$\hat{\tau}=\sup\Bigl{\{}\tau\in\mathcal{T}:\sup_{\begin{subarray}{c}\tau^{\prime}\in\mathcal{T}\\ \tau^{\prime}\leq\tau\end{subarray}}\hat{\alpha}(\tau^{\prime})\leq{\alpha}\Bigr{\}}.$ (14)
Above, the search space $\mathcal{T}$ is formulated as in Section 4, or using $\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]$, as in the proposed Adaptive method from Section 5. We remark that for $\hat{\tau}$ to be well defined, we assume there exists $\tau^{\prime}\in\mathcal{T}$ such that $\hat{\alpha}(\tau^{\prime})\leq\alpha$ This assumption can be trivially satisfied by setting $\hat{q}_{0}(X_{i})=0$. Then, with probability at least $1-\delta$ over the draws of $\mathcal{D}$, the LPB $\hat{L}(x)=\hat{q}_{\hat{\tau}}(x)$ satisfies
$\mathbb{P}\left[T_{\textup{test}}\geq\hat{L}(X_{\textup{test}})|\mathcal{D}\right]\geq 1-\alpha-\sup_{\tau\in[0,1]}\left\{\sqrt{\frac{2\gamma^{2}_{\tau}+5}{|{\mathcal{I}_{2}}|}\cdot\log\left(\frac{1}{\delta}\right)}\right\}.$ (15)
Proof.
For ease of notation, we define the coverage gap by:
$\Delta:=\sup_{\tau\in[0,1]}\left\{\sqrt{\frac{2\gamma^{2}_{\lambda}+5}{|{\mathcal{I}_{2}}|}\cdot\log\left(\frac{1}{\delta}\right)}\right\}.$ (16)
We define the oracle miscoverage level by:
$\tau(\alpha+\Delta)=\sup\left\{\lambda\in[0,1]:\mathbb{P}\left(T<\hat{q}_{\lambda}(X)\mid{\mathcal{I}_{1}}\right)\leq\alpha+\Delta\right\}.$ (17)
Observe that by assuming $1-\delta\leq\mathbb{P}(\hat{\tau}\leq\tau(\alpha+\Delta)\mid{\mathcal{I}_{1}})$, we get that the event $\{\hat{\tau}\leq\tau(\alpha+\Delta)\}$ holds with probability at least $1-\delta$. Under the monotonicity of $\hat{q}_{\tau}$, and following the the left-continuity of $\mathbb{P}(T\geq\hat{q}_{\tau}(X)\mid\mathcal{D})$ in $\tau$ we obtain that with probability at least $1-\delta$:
$\begin{split}&\mathbb{P}(T\geq\hat{q}_{\hat{\tau}}(X)\mid\mathcal{D})\\ \geq&\mathbb{P}(T\geq\hat{q}_{\tau(\alpha+\Delta)}(X)\mid\mathcal{D})\\ \geq&1-\alpha-\Delta.\end{split}$ (18)
Now, we turn to show that $1-\delta\leq\mathbb{P}(\hat{\tau}\leq\tau(\alpha+\Delta)\mid{\mathcal{I}_{1}})$. We begin by fixing $\varepsilon>0$, and denoting $\lambda:=\tau(\alpha+\Delta)+\varepsilon$. By the definition of $\hat{\alpha}(\tau)$, we get:
$\begin{split}&\mathbb{P}(\hat{\alpha}(\tau(\alpha+\Delta)+\varepsilon)\leq\alpha\mid{\mathcal{I}_{1}})\\ =&\mathbb{P}(\hat{\alpha}(\lambda)\leq\alpha\mid{\mathcal{I}_{1}})\\ =&\mathbb{P}\left(\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}w_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\;\mathbb{I}\bigl{\{}\hat{q}_{\lambda}(X_{i})\leq C_{i}\bigr{\}}\;\mathbb{I}\bigl{\{}T_{i}<\hat{q}_{\lambda}(X_{i})\bigr{\}}\leq\alpha\bigg{|}{\mathcal{I}_{1}}\bigg{.}\right)\\ =&\mathbb{P}\left(\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}w_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\;\mathbb{I}\bigl{\{}\hat{q}_{\lambda}(X_{i})\leq C_{i}\bigr{\}}\;\mathbb{I}\bigl{\{}T_{i}<\hat{q}_{\lambda}(X_{i})\bigr{\}}-\alpha\leq 0\bigg{|}{\mathcal{I}_{1}}\bigg{.}\right)\end{split}$ (19)
Next, we apply Markov’s inequality for any $t>0$, and obtain:
(19)$&\leq\mathbb{E}\left(\exp\left(\alpha-\frac{t}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}w_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\;\mathbb{I}\bigl{\{}\hat{q}_{\lambda}(X_{i})\leq C_{i}\bigr{\}}\;\mathbb{I}\bigl{\{}T_{i}<\hat{q}_{\lambda}(X_{i})\bigr{\}}\right)\bigg{|}{\mathcal{I}_{1}}\bigg{.}\right)\end{split}$ (20)
By conditioning on $\{(X_{i},T_{i})\}_{i\in{\mathcal{I}_{2}}}$ and following the $\frac{1}{4}$ sub-gaussianity of $w_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)^{-1}-\mathbb{I}\{\hat{q}_{\lambda}(X_{i})\leq C_{i}\}$, the conditionally independent censoring assumption (Assumption 2.1) and the bounded weights, we have:
$\begin{split}\mathbb{E}&\left(\exp\left(\frac{t}{|{\mathcal{I}_{2}}|}\cdot\sum_{i\in{\mathcal{I}_{2}}}w_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\mathbb{I}\bigl{\{}T_{i}<\hat{q}_{\lambda}(X_{i})\bigr{\}}\left(w_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)^{-1}-\mathbb{I}\{\hat{q}_{\lambda}(X_{i})\leq C_{i}\}\right)\right)\bigg{|}\{(X_{i},T_{i})\}_{i\in{\mathcal{I}_{2}}},{\mathcal{I}_{1}}\bigg{.}\right)\\ \leq&\exp\left(\frac{t^{2}}{8|{\mathcal{I}_{2}}|^{2}}\cdot\sum_{i\in{\mathcal{I}_{2}}}w_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)^{2}\mathbb{I}\bigl{\{}T_{i}<\hat{q}_{\lambda}(X_{i})\bigr{\}}^{2}\right)\\ \leq&\exp\left(\frac{t^{2}}{8|{\mathcal{I}_{2}}|^{2}}\cdot\sum_{i\in{\mathcal{I}_{2}}}w_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)^{2}\right)\\ \leq&\exp\left(\frac{\gamma^{2}_{\lambda}t^{2}}{8|{\mathcal{I}_{2}}|}\right).\end{split}$ (21)
Therefore,
(20)$&\leq\exp\left(\frac{\gamma^{2}_{\lambda}t^{2}}{8|{\mathcal{I}_{2}}|}\right)\cdot\mathbb{E}\left(\exp\left(\alpha-\frac{t}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}w_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\;w_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)^{-1}\;\mathbb{I}\bigl{\{}T_{i}<\hat{q}_{\lambda}(X_{i})\bigr{\}}\right)\bigg{|}{\mathcal{I}_{1}}\bigg{.}\right)\\ &=\exp\left(\frac{\gamma^{2}_{\lambda}t^{2}}{8|{\mathcal{I}_{2}}|}\right)\cdot\mathbb{E}\left(\exp\left(\alpha-\frac{t}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}\mathbb{I}\bigl{\{}T_{i}<\hat{q}_{\lambda}(X_{i})\bigr{\}}\right)\bigg{|}{\mathcal{I}_{1}}\bigg{.}\right)\end{split}$ (22)
Further, defining $p_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i):=\mathbb{P}(T_{i}<\hat{q}_{\lambda}(X_{i})\mid\{X_{j}\}_{j\in{\mathcal{I}_{2}}},{\mathcal{I}_{1}})$, we condition on $\{X_{j}\}_{j\in{\mathcal{I}_{2}}}$ and get, using the $\frac{1}{4}$ subgaussianity of $p_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)-\mathbb{I}\bigl{\{}T_{i}<\hat{q}_{\lambda}(X_{i})\bigr{\}}$:
$\begin{split}\mathbb{E}&\left(\exp\left(p_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i))-\frac{t}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}(\mathbb{I}\bigl{\{}T_{i}<\hat{q}_{\lambda}(X_{i})\bigr{\}}\right)\Bigg{|}\{X_{j}\}_{j\in{\mathcal{I}_{2}}},{\mathcal{I}_{1}}\bigg{.}\right)\\ \leq&\exp\left(\frac{t^{2}}{8|{\mathcal{I}_{2}}|^{2}}\cdot|{\mathcal{I}_{2}}|\right)\\ =&\exp\left(\frac{t^{2}}{8|{\mathcal{I}_{2}}|}\right)\end{split}$ (23)
We plug this into (22) to obtain:
(22)$\leq\exp\left(\frac{(\gamma^{2}_{\lambda}+1)t^{2}}{8|{\mathcal{I}_{2}}|}\right)\mathbb{E}\left(\exp\left(\frac{t}{|{\mathcal{I}_{2}}|}\cdot\sum_{i\in{\mathcal{I}_{2}}}(\alpha-p_{\lambda}(X_{i}))\right)\bigg{|}{\mathcal{I}_{1}}\bigg{.}\right)\end{split}$ (24)
We now apply the Cauchy-Schwarz inequality:
$\begin{split}&\mathbb{E}\left(\exp\left(\frac{t}{|{\mathcal{I}_{2}}|}\cdot\sum_{i\in{\mathcal{I}_{2}}}(\alpha-p_{\lambda}(X_{i}))\right)\bigg{|}{\mathcal{I}_{1}}\bigg{.}\right)\\ \leq&\mathbb{E}\left(\exp\left(\frac{2t}{|{\mathcal{I}_{2}}|}\cdot\sum_{i\in{\mathcal{I}_{2}}}(\alpha-p_{\lambda}(X_{i}))\right)\bigg{|}{\mathcal{I}_{1}}\bigg{.}\right)^{1/2}\end{split}$ (25)
Observe that by the definition of $\tau(\alpha+\Delta)$ we have:
$\mathbb{P}(T<\hat{q}_{\lambda}(X)\mid{\mathcal{I}_{1}})\geq\alpha+\Delta.$ (26)
By plugging it into the above, we get:
(25)$&=\mathbb{E}\left(\exp\left(\frac{2t}{|{\mathcal{I}_{2}}|}\cdot\sum_{i\in{\mathcal{I}_{2}}}(\alpha-p_{\lambda}(X_{i}))\right)\bigg{|}{\mathcal{I}_{1}}\bigg{.}\right)\\ \leq&\mathbb{E}\left(\exp\left(\frac{2t}{|{\mathcal{I}_{2}}|}\cdot\sum_{i\in{\mathcal{I}_{2}}}(\mathbb{P}(T<\hat{q}_{\lambda}(X)\mid{\mathcal{I}_{1}})-\Delta-p_{\lambda}(X_{i}))\right)\bigg{|}{\mathcal{I}_{1}}\bigg{.}\right)\\ =&\exp(-2t\Delta)\mathbb{E}\left(\exp\left(\frac{2t}{|{\mathcal{I}_{2}}|}\cdot\sum_{i\in{\mathcal{I}_{2}}}(\mathbb{P}(T<\hat{q}_{\lambda}(X)\mid{\mathcal{I}_{1}})-p_{\lambda}(X_{i}))\right)\bigg{|}{\mathcal{I}_{1}}\bigg{.}\right)\\ \leq&\exp\left(\frac{t^{2}}{2|{\mathcal{I}_{2}}|}-2t\Delta\right).\end{split}$ (27)
Above, we used the $\frac{1}{4}$-sub-gaussianity of $\mathbb{P}(T<\hat{q}_{\lambda}(X)\mid{\mathcal{I}_{1}})-p_{\lambda}(X_{i})$. By combining it all, we obtain:
(19)$\leq&\exp\left(\frac{t^{2}}{2|{\mathcal{I}_{2}}|}-2t\Delta+\frac{(\gamma^{2}_{\lambda}+1)t^{2}}{8|{\mathcal{I}_{2}}|}+\frac{\gamma^{2}_{\lambda}t^{2}}{8|{\mathcal{I}_{2}}|}\right)\\ =&\exp\left(\frac{(2\gamma^{2}_{\lambda}+5)t^{2}}{8|{\mathcal{I}_{2}}|}-2t\Delta\right)\end{split}$ (28)
We define
$t=\frac{(8+2\sqrt{14})|{\mathcal{I}_{2}}|\Delta}{2\gamma^{2}_{\lambda}+5}$ (29)
and plug it into (24):
(24)$\leq&\exp\left(\frac{(2\gamma^{2}_{\lambda}+5)(\frac{(8+2\sqrt{14})|{\mathcal{I}_{2}}|\Delta}{2\gamma^{2}_{\lambda}+5})^{2}}{8|{\mathcal{I}_{2}}|}-\frac{(16+4\sqrt{14})|{\mathcal{I}_{2}}|\Delta}{2\gamma^{2}_{\lambda}+5}\Delta\right)\\ =&\exp\left(\frac{(15+4\sqrt{14})|{\mathcal{I}_{2}}|\Delta^{2}}{(2\gamma^{2}_{\lambda}+5)}-\frac{(16+4\sqrt{14})}{2\gamma^{2}_{\lambda}+5}|{\mathcal{I}_{2}}|\Delta^{2}\right)\\ =&\exp\left(-\frac{1|{\mathcal{I}_{2}}|\Delta^{2}}{2\gamma^{2}_{\lambda}+5}\right)\\ \leq&\exp\left(-\log\left(\frac{1}{\delta}\right)\right)\\ =&\delta\\ \end{split}$ (30)
That is, we just showed that:
$1-\delta\leq\mathbb{P}(\hat{\alpha}(\tau(\alpha+\Delta)+\varepsilon)>\alpha\mid{\mathcal{I}_{1}})\leq\mathbb{P}(\hat{\tau}<\tau(\alpha+\Delta)+\varepsilon\mid{\mathcal{I}_{1}}).$ (31)
The above equation holds for every $\varepsilon>0$, and thus by taking $\varepsilon\rightarrow 0$, and following the continuity of the probability measure, we obtain that $1-\delta\leq\mathbb{P}(\hat{\tau}\leq\tau(\alpha+\Delta)\mid{\mathcal{I}_{1}})$, as required. ∎
Since the proposed methods construct censoring times that satisfy the conditional independence assumption of Theorem D.1, the proofs of Proposition 4.1 and Theorem 5.2 follow immediately from this theorem with the corresponding definitions of the censoring times and the weights.
Appendix E Experimental details
This section provides additional information about the experimental setup and methodology described in Section 6. Section E.1 details the synthetic experiments, including data generation and model architecture. Section E.2 presents experiments conducted on real-world data. Finally, Section E.3 outlines implementation details that are shared across both experimental settings.
E.1 Synthetic experiments
Synthetic covariate generation
As described in Section 6.1 of the main text, we use synthetic data to evaluate the informativity and coverage of the LPBs constructed by our calibration procedures. Each synthetic data point consists of covariates $X_{i}\in\mathbb{R}^{d}$ with $d=10$, and a prompt-specific unsafe-output probability $p_{i}$. The dataset is designed to simulate a setting in which most prompts have a high likelihood of producing unsafe outputs, while a smaller subset is relatively safe. Notably, our experiments show that fitting a model on this synthetic dataset results in a large prediction error, which turns the calibration step more challenging.
To this end, we sample the covariates from a Gaussian distribution:
$X_{i}\mid p_{i}\sim\mathcal{N}\left(\mu(p_{i}),\sigma^{2}I_{d}\right),\quad\text{with }\sigma=0.1.$ (32)
where the mean vector $\mu(p_{i})\in\mathbb{R}^{d}$ is designed as follows. First, we construct a pool of $p_{i}$ values, where (i) 90% of which are drawn uniformly in the range $\log_{10}(p_{i})\sim\mathcal{U}[-4,-3]$, and (ii) the remaining 10% are drawn uniformly in the range $\log_{10}(p_{i})\sim\mathcal{U}[-6,-5]$. We then assign each $p_{i}$ to a data point by sampling without replacement from this pool. Then, we define $d$ quantile levels as follows
$\tau_{j}=0.1+0.8\,\frac{j-1}{d-1}\ \ \ j=1,\ldots,d.$
Next, for a given probability $p_{i}$, we compute:
$\tilde{\mu}(p_{i})=\bigl{[}F^{-1}_{\mathrm{Geom}(p_{i})}(\tau_{1})^{1/4},\,\ldots,\,F^{-1}_{\mathrm{Geom}(p_{i})}(\tau_{d})^{1/4}\bigr{]}^{\top},$
where $F^{-1}_{\mathrm{Geom}(p_{i})}$ is the quantile function of the Geometric distribution with success probability $p_{i}$. The $1/4$ power transformation reduces the size of the often extremely high raw quantiles. Lastly, to improve numerical stability, we normalize this vector by the average magnitude across all training examples:
$\bar{\mu}=\frac{1}{n\,d}\sum_{i=1}^{n}\sum_{j=1}^{d}F^{-1}_{\mathrm{Geom}(p_{i})}(\tau_{j})^{1/4},\quad\mu(p_{i})=\frac{\tilde{\mu}(p_{i})}{\bar{\mu}}.$
Model architecture and training
To estimate the unsafe generation probabilities, we train a neural network with four hidden layers of size 32 each. We used ReLU for non-linearity and a sigmoid for output. We optimize the BCE loss (see Section E.3) using AdamW with learning rate $10^{-4}$, weight decay $10^{-5}$, batch size $100$, for $10$ epochs.
E.1.1 Additional experiments
To confirm that our findings on synthetic data do not depend on a particular choice of calibration/test split, we repeated the experiment from Section 6.1 for 20 independent splits of calibration and test sets. Specifically, in each trial, we fixed the original training set of 45,000 prompts, then randomly partitioned the remaining pool into a calibration set of size 45000 and a test set of size 10,000. All other settings follow the experimental setup described in Section E.1.
We repeated the experiment for the following values of average sampling budget per prompt
$B/|I_{2}|:{10,25,50,100,200,300,600,1200}.$
In all experiments, we used the same prior quantile $\tau_{\mathrm{prior}}=10^{-1/4}$, with trimming threshold $M$ set so that $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right)=10$, as in Section 6.1.
Figure 3 presents the performance of the methods introduced in this work for 20 data splits. Across every budget level, the performances closely match those from our original fixed-split experiment from Figure 1 in Section 6.1.
Figure 3: Results of synthetic experiments as a function of average budget per prompt $B/|{\mathcal{I}_{2}}|$. Left: Coverage (target 90%; gray dashed line). Center: Mean number of samplings generated per prompt. Right: Mean LPB. Shaded regions represent the standard deviation over 20 runs.
Next, we evaluate how the target coverage level affects the constructed LPBs and their informativeness by repeating the experiment from Section 6.1 across different nominal coverage levels. We present the results in Figure 4. This figure shows that the Optimized consistently achieves its target coverage across all levels.
Figure 4: Results of synthetic experiments as a function of average budget per prompt $B/|{\mathcal{I}_{2}}|$. Top: Coverage (target $100\cdot(1-\tau)$%; gray dashed line). Bottom: Mean LPB. Shaded regions represent the standard deviation over 20 runs.
Finally, we turn to study the effect of the maximal weight size $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right)$ on the stability and power of the Trimmed and Optimized calibration methods. Figure 5 compares the two methods under a fixed per-prompt budget of ${B}/{|{\mathcal{I}_{2}}|}=1000$. For each weight size, we set the threshold $M={B\gamma}/{|{\mathcal{I}_{2}}|}$. As can be seen in this figure, the size of the LPBs we construct increases with $\gamma$. However, this comes at the cost of increasing the coverage variance. Observe how the variance of Optimized method for $\gamma=100$ is similar to that of Trimmed method for $\gamma=10$, but the former yields much more informative LPBs under these corresponding $\gamma$ values. This is a result of the variance-reducing optimization objective in (8).
Figure 5: Results of synthetic experiments as a function of the maximum weight $w_{\text{max}}$ of the Trimmed and Optimized methods. Left: Coverage (target 90%; gray dashed line). Right: Average LPB. Shaded regions present the standard deviation across $20$ runs.
E.2 Real-Data Experiments
Efficient prompt-parallel sampling
To efficiently generate outputs during our calibration procedures, we implement a prompt-parallel generation algorithm on top of the vLLM package [33]. In each iteration, we gather a batch of $1500$ prompts ${X_{i}}$ along with their corresponding censoring budgets ${C_{i}}$, and distribute them across multiple GPUs. The pipeline iteratively performs the following steps in parallel for a batch of prompts, until all prompts in the calibration set are processed:
  1. 1.
    Generate an output $\mathcal{G}(X_{i})$ using the LLM.
  2. 2.
    Apply the audit function $\mathrm{Audit}\bigl{(}X_{i},\mathcal{G}(X_{i})\bigr{)}$ to identify unsafe outputs.
  3. 3.
    Check whether $X_{i}$ has either: (i) reached the $C_{i}$ generations limit; or (ii) produced an output that fails the audit. If at least one of these two conditions is met, remove the prompt from the pipeline and replace it with a new prompt.
By promptly removing any prompt that yields an unsafe output, the processing maintains high throughput and avoids idle GPU time.
Runtime evaluation
Table 1 presents average runtime for the Naive and Optimized calibration procedures across varying prompt budgets. While the Naive method can have a long runtime due to its unbounded censorship times $C_{i}$, it is observed to be faster than the Optimized method; see Table 1. We explain this by two key factors. First, our implementation of the Naive method includes an efficient runtime optimization: instead of always generating all $C_{i}$ samples, we terminate generation early when additional samples no longer improve the estimate of $\hat{\alpha}(\tau)$. As detailed in Section E.3.1, this strategy yields LPBs that are equivalent to those produced by the original Naive procedure, while significantly reducing computation time. Second, the base model used in these experiments tends to produce low predicted quantiles $\hat{q}_{\tau}(X_{i})$, as shown in Figure 2. Consequently, the effective censorship budgets $C_{i}$ are also low for most prompts, further limiting runtime.
In sum, these two factors—runtime-efficient implementation and tendency to produce low predicted quantiles—explain why the reported runtime of the Naive is lower than the Optimized method in this setting.
Budget per prompt Naive runtime (hours) Optimized runtime (hours) # GPUs
10 0.032 $\pm$ 0.005 0.063 $\pm$ 0.001 6
25 0.042 $\pm$ 0.001 0.117 $\pm$ 0.001 6
50 0.070 $\pm$ 0.006 0.209 $\pm$ 0.001 6
100 0.124 $\pm$ 0.003 0.388 $\pm$ 0.003 6
200 0.215 $\pm$ 0.010 0.589 $\pm$ 0.007 6
300 0.316 $\pm$ 0.012 0.718 $\pm$ 0.002 6
600 0.561 $\pm$ 0.016 0.979 $\pm$ 0.006 6
1200 0.918 $\pm$ 0.051 1.792 $\pm$ 0.041 4
Table 1: Average runtime (hours) by method and budget per prompt (mean $\pm$ SD).
Model architecture and training
We employ the ModernBERT-base [61] model. We use Adam optimizer [29] to minimize the loss function in (33), with a learning rate of 3e-5, and a batch size of $1500$. We train the model for 10 epochs and use early stopping based on the validation set.
E.3 Shared implementation details for both real and synthetic experiments
Our calibration framework applies to any machine learning model that outputs an estimated conditional quantile of the time-to-unsafe-sampling, denoted by $\hat{q}_{\tau}(x)$, given the covariates $x$. This is motivated by the fact that $T$, the time-to-unsafe-sampling, is a Geometric random variable with parameter $p(x)$, the probability of unsafe generation. Consequently, the conditional quantile function of $T$ is given analytically in terms of $p(x)$, and vice versa:
$q_{\tau}(x)=\left\lceil\frac{\log(1-\tau)}{\log(1-p(x))}\right\rceil\quad\text{and}\quad p(x)=1-(1-\tau)^{1/q_{\tau}(x)}.$
The above relationship allows us to estimate either $p(x)$ or $q_{\tau}(x)$ and recover the other via a closed-form transformation.
Loss function
In our experiments, we estimate $p(x)$ by fitting a model that minimizes the binary cross-entropy (BCE) loss over aggregate unsafe proportions. Specifically, for each prompt $X_{i}$, we define the empirical success rate as $\bar{Y}_{i}=\frac{1}{N}\sum_{j=1}^{N}Y_{i}^{j}$, where $Y_{i}^{j}\in\{0,1\}$ indicates whether the $j$-th sample is unsafe. We then use the BCE loss:
$\mathrm{BCE}(\bar{Y}_{i},\hat{p}(X_{i}))=-\bigl{[}\bar{Y}_{i}\log\hat{p}(X_{i})+(1-\bar{Y}_{i})\log(1-\hat{p}(X_{i}))\bigr{]},$ (33)
where $\hat{p}(X_{i})$ is the model’s prediction of the probability of unsafe generation. Notably, this loss is equivalent to the mean of the standard BCE loss across individual samples:
$\frac{1}{N}\sum_{j=1}^{N}\mathrm{BCE}(Y_{i}^{j},\hat{p}(X_{i}))=\mathrm{BCE}(\bar{Y}_{i},\hat{p}(X_{i})).$
This aggregation allows for a more computationally and memory-efficient implementation compared to the use of the vanilla BCE.
E.3.1 Implementation details of the Naive calibration procedure
In theory, the Naive calibration (Algorithm 1) can assign an unbounded censorship time per prompt, which leads to arbitrarily long running times. To avoid this, we restrict our search for the threshold $\tau$ to the compact set $\mathcal{T}\cap\tau_{\mathrm{prior}}$ (see Section 6.2). This restriction in search space does not violate the method’s validity, as it may only make it more conservative, as discussed in Section 5.1. This section also shows that assigning a censorship time $C_{i}\geq\hat{q}_{\tau}(X_{i})$ is equivalent to drawing exactly $\hat{q}_{\tau}(X_{i})$ generations. Conversely, setting $C_{i}<\hat{q}_{\tau}(X_{i})$ amounts to not drawing any samples. Based on this observation, we implement a run-time efficient version of the Naive method that draws censorship times according to
$C_{i}:=\text{Ber}(g(X_{i}))\cdot\hat{q}_{\tau_{\text{prior}}}(X_{i}),$
where $g(x)=\mathbb{P}(\mathrm{Geom}(\min(|{\mathcal{I}_{2}}|/B,1)\geq\hat{q}_{\tau_{\text{prior}}}(X_{i}))$. It is important to note at this point that the average number of generations per prompt reported in Figure 1 is the simulated number of samples that the equivalent, non-optimized Naive solution would use. This implementation also implies that the Naive method has a much smaller overall number of generations, leading to a faster runtime.
E.4 Machine specifications
The hardware and OS used for the experiments are as follows.
  • CPU: AMD EPYC 7443 24-Core Processor
  • GPU: NVIDIA RTX A6000
  • OS: Ubuntu 20.04