### v4

# Calibrated Predictive Lower Bounds on Time-to-Unsafe-Sampling in LLMs

Abstract
We introduce time-to-unsafe-sampling, a novel safety measure for generative models, defined as the number of generations required by a large language model (LLM) to trigger an unsafe (e.g., toxic) response. While providing a new dimension for prompt-adaptive safety evaluation, quantifying time-to-unsafe-sampling is challenging: unsafe outputs are often rare in well-aligned models and thus may not be observed under any feasible sampling budget. To address this challenge, we frame this estimation problem as one of survival analysis. We build on recent developments in conformal prediction and propose a novel calibration technique to construct a lower predictive bound (LPB) on the time-to-unsafe-sampling of a given prompt with rigorous coverage guarantees. Our key technical innovation is an optimized sampling-budget allocation scheme that improves sample efficiency while maintaining distribution-free guarantees. Experiments on both synthetic and real data support our theoretical results and demonstrate the practical utility of our method for safety risk assessment in generative AI models.
1 Introduction
Despite substantial progress in aligning LLMs to human intent, current systems remain vulnerable to rare but impactful unsafe outputs that emerge through repeated sampling (Bianchi et al., 2024, Vidgen et al., 2023, Ray, 2023). For example, a seemingly benign prompt like “Tell me a joke that would make my grandfather laugh” can eventually elicit outputs that marginalize minorities—not necessarily on the first try, but after many samplings. Although the cumulative risk of unsafe outputs from repeated sampling is a realistic concern at scale, existing safety evaluation methods often overlook it. This oversight sparks the question we pursue in this paper: how can we provide rigorous foresight into these rare yet impactful unsafe generations, before they occur?
We define an unsafe output as content such as toxic language (Vidgen et al., 2023), the disclosure of private or sensitive information (Kaddour et al., 2023), or deepfakes that defame individuals (Westerlund, 2019). A standard retrospective approach to protect users from such outputs is to audit the LLM’s response at inference time using an Audit function (Inan et al., 2023, Hu et al., 2024). This mechanism can either trigger refusal or allow the model to generate new responses until a safe one is observed (Ahmed et al., 2025, Bhatt et al., 2025, Stroebl et al., 2024).
While effective, this retrospective approach introduces additional computational overhead and latency, as it requires invoking the Audit function during inference. More critically, it offers no foresight for making proactive safety decisions, such as whether to deploy a more conservative system prompt, restrict the use of certain prompts, or selectively apply the Audit function when needed.
In this paper, we complement retrospective safety auditing with a more proactive approach. Specifically, for a given prompt and an LLM, we ask: “How many safe responses can we expect before encountering an unsafe one?” We refer to this quantity as the time-to-unsafe-sampling. This measure indicates when stronger safety checks are necessary, provides a practical basis for setting usage limits, and serves as a safety evaluation metric across different models and prompt.
A direct way to estimate time-to-unsafe-sampling is to train a regression model, ideally using data annotated by generating multiple responses per prompt until an unsafe outcome is observed. However, with well-aligned LLMs, the number of generations required to trigger an unsafe output can be exceedingly large. As a result, under any realistic sampling budget $B\in\mathbb{N}$, the value of the time-to-unsafe-sampling can be censored—i.e., unobserved—for some prompts. This censoring problem makes the prediction task challenging: naively applying regression to such partially observed data can result in forecasts that systematically over- or under-estimate the true time-to-unsafe-sampling.
Recognizing the difficulty of accurately predicting the time-to-unsafe-sampling, we instead focus on calibrating the output of the regression model to provide a reliable lower bound on this quantity; an objective we show to be attainable. Denote by $\hat{L}(X_{\textup{test}})$ a lower predictive bound (LPB) on the unknown time-to-unsafe-sampling $T_{\textup{test}}$ of a new test prompt $X_{\textup{test}}$. We define the coverage rate of $\hat{L}(X_{\textup{test}})$ as the probability that the true (unknown) time-to-unsafe-sampling $T_{\textup{test}}$ exceeds the LPB, i.e., $\mathbb{P}(T_{\textup{test}}\geq\hat{L}(X_{\textup{test}}))$. Our goal is to construct a calibrated LPB with a formal, user-specified coverage rate guarantee, such as 90%. Informally, this means that, with high probability, one can expect to generate at least $\hat{L}(X_{\textup{test}})$ safe responses to the test prompt $X_{\textup{test}}$ before encountering an unsafe one, on average across future test prompts.
To construct an LPB for time-to-unsafe-sampling, we formulate the LLM prompt risk assessment problem as a survival analysis task. In survival analysis, the goal is to predict a time-to-event—e.g., a patient’s time to mortality—given covariates, such as clinical lab results. Similar to our setting, an inherent challenge in survival data is that the time-to-event can be censored due to the patient’s withdrawal of consent or the end of the clinical trial (Machin et al., 2006). This structure mirrors our setting, where time-to-unsafe-sampling is censored due to a finite LLM-sampling budget. Leveraging this analogy, we adopt recent conformalized survival analysis methods, which provide a principled way to construct LPBs with rigorous, finite-sample coverage guarantees under censoring (Gui et al., 2024, Candès et al., 2023).
However, in existing conformalized survival analysis works, the censoring mechanism for each subject is externally defined, e.g., by study design, and treated as fixed. Put simply, these methods do not intervene in how the data is collected. In contrast, our time-to-unsafe-sampling setting presents a unique opportunity: the censoring times—i.e., how many LLM samplings are allocated per prompt—can be actively designed under a sampling budget $B$. Indeed, our key technical innovation is an optimized sampling-budget allocation scheme that improves the sample-size efficiency of our conformal-based LPB construction method, while maintaining its coverage guarantees.
In sum, our main contributions are as follows:
  1. (1)
    We propose time-to-unsafe-sampling: a new safety measure that quantifies the risk of temporal, probabilistic failures at a fine-grained, per-prompt level (Section 2).
  2. (2)
    We refine the method of Gui et al. (2024), simplifying it to construct LPBs with a formal coverage guarantee under a known censoring mechanism (Section 4). Notably, this coverage guarantee is LLM-agnostic and holds in finite samples, regardless of the prompt distribution and the sampling budget. Moreover, obtaining an LPB from our method at inference time requires only a single query to the calibrated regression model, without invoking the Audit model or generating multiple LLM samples.
  3. (3)
    We analyze how the sampling-budget allocation influences the variance and coverage rate achieved by our LPBs (Section 5). Building on these insights, we formulate an optimization objective for budget allocation. Solving this objective yields a censoring mechanism that offers tighter control over coverage compared to non-optimized baselines (Section 5.4).
  4. (4)
    We validate our theory on simulated and real-data experiments. Specifically, our experiments with the “RealToxicityPrompts” dataset (Gehman et al., 2020) using Llama 3.2 (Meta AI, 2024, Grattafiori et al., 2024), demonstrate that the proposed method yields more informative bounds and more reliable risk estimates than baseline approaches (Section 6). Code is provided for reproducibility at http://github.com/giladfrid009/llm-survival/.
Remark on the structure of the paper. We progressively build our sampling-budget allocation mechanism, which enables us not only to introduce our flagship Optimized approach, but also to develop and evaluate strong baseline methods for comparison. We begin with a Naive baseline that illustrates the core challenges in sampling-budget allocation. We then introduce two increasingly refined per-prompt allocation strategies: Basic and Trimmed. Finally, we present our flagship Optimized method, which we show to outperform the above baselines.
2 Problem setup
Let $\mathcal{D}\!=\!\{X_{i}\}_{i=1}^{n}$ be a dataset of user prompts, drawn i.i.d. from some distribution $P_{X}$. Let $\mathcal{G}(x)$ denote an LLM. For each prompt $X_{i}$, let $\{\mathcal{G}^{j}(X_{i})\}_{j=1}^{\infty}$ be a sequence of responses generated by repeated calls to $\mathcal{G}(X_{i})$. We define the binary indicator $Y_{i}^{j}\!=\!\texttt{Audit}\bigl(X_{i},\mathcal{G}^{j}(X_{i})\bigr),$ where $Y_{i}^{j}=1$ if and only if the $j$-th response is labeled unsafe by the Audit function. The uncensored time-to-unsafe-sampling for prompt $X_{i}$ is then
$T_{i}\!=\!\min\bigl\{\,j\geq 1:Y_{i}^{j}=1\bigr\}.$
In the simplest case, conditional on $X_{i}$, the sequence $\{Y_{i}^{j}\}$ is i.i.d. Bernoulli with unsafe sampling probability $p(X_{i})$, implying $T_{i}\mid X_{i}\sim\mathrm{Geom}\!\bigl(p(X_{i})\bigr)$. In practice, however, successive generations may be dependent; for instance, through iterative self-reflective prompting (Madaan et al., 2023, Shinn et al., 2023) or LLM tool use where tool availability or behavior changes over time (Wang et al., 2023). Crucially, the method we propose does not assume that $\{Y_{i}^{j}\}_{j}$ are independent conditional on $X_{i}$; our guarantees remain valid under arbitrary dependence within each sequence. This should not be confused with the assumption that the prompts $\{X_{i}\}_{i}$ are i.i.d., which is required for our theoretical analysis.
Given a user-specified coverage level $1-\alpha\in(0,1)$, and a tolerance level $\delta\in(0,1)$, our goal is to construct an LPB $\hat{L}$ that satisfies
$\mathbb{P}\big(T_{\textup{test}}\geq\hat{L}(X_{\textup{test}})\mid\mathcal{D}\big)\geq 1-\alpha,$ (1)
with probability at least $1-\delta$ over the randomness of the training data $\mathcal{D}$. The probability in (1) is taken over the distribution of a test point $(X_{\textup{test}},T_{\textup{test}})\sim P_{X,T}$. An LPB $\hat{L}(X_{\textup{test}})$ satisfying this requirement is called a Probably Approximately Correct (PAC) LPB at level $\alpha$ with tolerance $\delta$.
We follow standard split-conformal methodology (Vovk et al., 2005), and partition the index set $\{1,\dots,n\}$ of the training prompts $\mathcal{D}$ into a proper training set $\mathcal{I}_{1}$ and a holdout calibration set $\mathcal{I}_{2}$. The set $\mathcal{I}_{1}$ is used to train a regression model to predict $T$, which, in general, does not have formal validity guarantees, particularly under model misspecification or when using complex deep learning algorithms. The calibration set $\mathcal{I}_{2}$ is used to calibrate the output of this “black-box” regression model to form a valid LPB. Our focus in this work is on the calibration stage.
We operate under a global sampling budget $B\in\mathbb{N}$, reflecting a realistic setting in which computational and monetary resources are limited. To respect this budget during calibration, we introduce a per-prompt censoring time $C_{i}\in\mathbb{N}$, which limits the number of response–audit rounds performed for each prompt $X_{i}$. Formally, we require that
$\mathbb{E}\Bigl[\sum_{i\in{\mathcal{I}_{2}}}C_{i}\;\Big|\;\{X_{i}\}_{i\in{\mathcal{I}_{2}}}\Bigr]\!\leq\!B.$ (2)
In words, the above ensures that the expected number of response–audit rounds does not exceed the overall budget BB. [^1] We remark that in our setup, the training procedure has a separate sampling budget.
In Section 5, we propose specific schemes for designing the censoring mechanism under the budget constraint (2). This results in a calibration dataset of the form $\{X_{i},C_{i},\tilde{T}_{i}\}_{i\in\mathcal{I}_{2}}$, where $\tilde{T}_{i}\!=\!\min\bigl(T_{i},C_{i}\bigr)$ denotes the censored time-to-unsafe-sampling. That is, if an unsafe response for $X_{i}$ is triggered within the first $C_{i}$ generations—i.e., among $Y_{i}^{1},Y_{i}^{2},\dots,Y_{i}^{C_{i}}$—then $\tilde{T}_{i}=T_{i}$; otherwise, the time-to-unsafe-sampling is censored at $C_{i}$, and we have $\tilde{T}_{i}=C_{i}$.
In addition to satisfying the global budget constraint (2), we design the censoring mechanism to support valid statistical inference. As is standard in survival analysis, additional assumptions are required to make inference about the (potentially censored) time-to-unsafe-sampling variable $T$ feasible. Specifically, our censoring mechanism is constructed to satisfy a widely adopted assumption in the survival literature that the censoring time $C$ is conditionally independent of the survival time $T$ given the covariates $X$:
Assumption 2.1 (Conditionally Independent Censoring (Kalbfleisch & Prentice, 2011)).
[…]
3 Related work
In this section, we present the conformalized survival analysis algorithm of Gui et al. (2024), called Adaptive-T, which constructs a valid LPB in the sense of (1) under Assumption 2.1. Additional related work is discussed in Appendix A.
Given calibration data $\{X_{i},C_{i},\tilde{T}_{i}\}_{i\in\mathcal{I}_{2}}$, Adaptive-T constructs a PAC-type LPB $\hat{L}$ by calibrating a pre-trained quantile regression model. To motivate the use of quantile regression, consider an oracle setting in which the conditional distribution of $T\mid X=x$ is known. Let $q_{\tau}(x)$ denote the true $\tau$-th conditional quantile of $T\mid X=x$. In this case, the oracle LPB at coverage level $1-\alpha$ for a new test prompt $X_{\textup{test}}$ is simply $\hat{L}(X_{\textup{test}})=q_{\alpha}(X_{\textup{test}})$, which achieves the desired coverage by definition.
Unfortunately, since $q_{\alpha}(x)$ is unknown in practice, we cannot compute this oracle bound. Instead, we can estimate the conditional quantile function, denoted by $\hat{q}_{\tau}(x)$, and then obtain a naive plug-in LPB: $\hat{L}(X_{\textup{test}})\!=\!\hat{q}_{\alpha}(X_{\textup{test}}).$ However, this naive LPB might fail to attain valid coverage if $\hat{q}_{\alpha}(x)$ is not sufficiently accurate due to model misspecification, overfitting, or limited data, highlighting the need for calibration.
The Adaptive-T method addresses this limitation by finding a calibrated quantile level $\hat{\tau}^{\text{adapt}}$ such that $\hat{L}(X_{\textup{test}})\!=\!\hat{q}_{{\hat{\tau}}^{\text{adapt}}}(X_{\textup{test}})$ approximately satisfies (1) in a PAC sense. At a high level, this calibration is done by estimating the miscoverage probability $\mathbb{P}\bigl(T<\hat{q}_{\tau}(X)\bigr)$ for each candidate $\tau$ using the held-out calibration set $\mathcal{I}_{2}$}. Then, this method proceeds by choosing the largest $\tau$ whose estimated miscoverage is less than or equal to $\alpha$.
In detail, for each level $\tau$, the miscoverage estimator $\hat{\alpha}^{\text{adapt}}(\tau)$ proposed by Gui et al. (2024) is given by
$\hat{\alpha}^{\text{adapt}}(\tau)=\frac{\sum_{i\in\mathcal{I}_{2}}\hat{w}_{\tau}(X_{i})\,\mathbb{I}\{\tilde{T}_{i}<\hat{q}_{\tau}(X_{i})\leq C_{i}\}}{\sum_{i\in\mathcal{I}_{2}}\hat{w}_{\tau}(X_{i})\,\mathbb{I}\{\hat{q}_{\tau}(X_{i})\leq C_{i}\}}.$ (3)
We pause to parse the above expression. First, observe that the indicator $\mathbb{I}\{\tilde{T}_{i}<\hat{q}_{\tau}(X_{i})\leq C_{i}\}$ is a function of observed quantities. Second, under Assumption 2.1, Gui et al. (2024) show that this indicator introduce a covariate shift. To account for this shift, each selected example is re-weighted by an “inverse-censoring” weight $\hat{w}_{\tau}(X_{i})$ that approximates $1/\mathbb{P}\bigl[\hat{q}_{\tau}(X_{i})\leq C_{i}\bigm|X_{i}\bigr]$. Lastly, the denominator in (3) is used to normalize the weighted average, which is crucial when the estimated weights $\hat{w}_{\tau}(x)$ are merely proportional (rather than exactly equal) to the true inverse probabilities.
Finally, Adaptive-T defines the calibrated quantile level as $\hat{L}(x)=\hat{q}_{\hat{\tau}^{\text{adapt}}}(x)$, where
$\hat{\tau}^{\text{adapt}}=\sup\Bigl\{\tau\in\mathcal{T}:\sup_{\tau^{\prime}<\tau}\hat{\alpha}^{\text{adapt}}(\tau^{\prime})\leq\alpha\Bigr\},$ (4)
with $\mathcal{T}$ is the search space for $\tau$. In turn, the resulting $\hat{L}(x)=\hat{q}_{\hat{\tau}^{\text{adapt}}}(x)$ is a valid PAC-type LPB assuming that either the quantile estimates $\hat{q}_{\tau}$ or the censoring‐weight estimates $\hat{w}_{\tau}$ are sufficiently accurate.
4 Calibration with a known censoring mechanism
Recall that in our LLM prompt risk assessment setting, we design the sampling-budget allocation scheme. As a result, we have full knowledge of the inverse-censoring weights required to implement the method of Gui et al. (2024) presented in the previous section. In fact, this knowledge allows us to further simplify their procedure by leveraging the fact that the censoring distribution is known by construction.
In this section, we first introduce the calibration algorithm tailored to our setting. We then analyze the factors that influence its coverage guarantee and tightness. This analysis not only underpins our calibration procedure, but also guides our sampling-budget allocation strategies developed in Section 5.
Formally, the inverse-censoring weight corresponding to each $\hat{q}_{\tau}$ is defined as
$w_{\tau}(X_{i})=1/\mathbb{P}\bigl[\hat{q}_{\tau}(X_{i})\leq C_{i}\bigm|X_{i}\bigr].$ (5)
Using these weights, we estimate the empirical miscoverage rate at quantile level $\tau$ as a weighted average over the calibration set:
$\hat{\alpha}(\tau)\;=\;\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}w_{\tau}(X_{i})\;\mathbb{I}\bigl\{\tilde{T}_{i}<\hat{q}_{\tau}(X_{i})\leq C_{i}\bigr\}.$ (6)
Note that this estimator differs from the one used in (3): it does not require additional normalization since our proposed censoring mechanisms provide the true weights $w_{\tau}(X_{i})$, as explained in Section 5.
Armed with $\hat{\alpha}(\tau)$, we obtain calibrated LPB by following (4) and selecting the largest $\tau$ such that $\hat{\alpha}(\tau)\leq\alpha$:
$\hat{L}(x)=\hat{q}_{\hat{\tau}}(x),\ \ \hat{\tau}=\sup\bigl\{\tau\in\mathcal{T}:\sup_{\tau^{\prime}<\tau}\hat{\alpha}(\tau^{\prime})\leq\alpha\bigr\},\ \ \$(7)
where $\mathcal{T}=\bigl\{\sup_{\tau\in[0,1]}\{\hat{q}_{\tau}(X_{i})\leq\tilde{T}_{i}\}:i\in{\mathcal{I}_{2}}\}\cup\{\sup_{\tau\in[0,1]}\{\hat{q}_{\tau}(X_{i})\leq C_{i}\}:{i\in{\mathcal{I}_{2}}}\bigr\}\cup\{0\}$.
The following informal theorem shows that the LPB $\hat{L}$ formulated in (7) holds a PAC-type coverage guarantee. In Appendix D.6 we state the formal version of this theorem and prove it by building on the theoretical framework developed by Gui et al. (2024).
Theorem 4.1 (General validity, informal).
Fix a tolerance level $\delta\in\left(0,1\right)$ and a miscoverage level $\tau\in\left(0,1\right)$. Suppose that $\{(X_{i},T_{i})\}_{i=1}^{n}$ and $(X_{\textup{test}},T_{\textup{test}})$ are drawn i.i.d and that the censoring times satisfy the conditional independence assumption (Assumption 2.1). Further, assume that there exists a constant ${\gamma}_{\tau}>0$ such that the weights satisfy ${w}_{\tau}(x)\leq{\gamma}_{\tau}$ for almost all $x$. Then, with probability at least $1-\delta$ over the draws of $\mathcal{D}$, the LPB $\hat{L}(x)=\hat{q}_{\hat{\tau}}(x)$ from (7) satisfies
$\mathbb{P}\left[T_{\textup{test}}\geq\hat{L}(X_{\textup{test}})|\mathcal{D}\right]\geq 1-\alpha-\sup_{\tau\in[0,1]}\left\{\sqrt{\frac{2\gamma^{2}_{\tau}+5}{|{\mathcal{I}_{2}}|}\cdot\log\left(\frac{1}{\delta}\right)}\right\}.$
Importantly, the above guarantee holds in finite samples, for any LLM $\mathcal{G}(x)$, any prompts distribution $P_{X}$, sampling-budget $B$, and regardless of the accuracy of the quantile estimators $\hat{q}_{\tau}(x)$. However, this coverage bound depends on the supremum of all possible weights with non-zero probability, denoted by $\gamma_{\tau}$. In particular, as $\gamma_{\tau}$ increases, the bound becomes looser.
In Appendix C, we further analyze the calibration method in greater depth by studying the properties of $\hat{\alpha}(\tau)$ defined in (6). Specifically, in Proposition C.2, we make an illustrative assumption that, for any fixed $\tau$, each time-to-unsafe-sampling $T_{i}$ is miscovered by $\hat{q}_{\tau}(X_{i})$ at the same constant rate, conditional on the calibration prompts $\{X_{j}\}_{j\in\mathcal{I}_{2}}$. This assumption is satisfied by the oracle conditional quantile $q_{\tau}(\cdot)$ because, by definition, $\mathbb{P}\left[T_{i}<q_{\tau}(X_{i})\mid X_{i}\right]=\tau$ for every $i$, and hence also conditional on $\{X_{j}\}_{j\in\mathcal{I}_{2}}$. Under this assumption, we prove that the conditional variance $\mathrm{Var}\!\bigl[\hat{\alpha}(\tau)\,\bigm|\,\{X_{j}\}_{j\in\mathcal{I}_{2}}\bigr]$ is linearly monotone with regard to the mean calibration weight
$\overline{w}_{\tau}=\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in{\mathcal{I}_{2}}}w_{\tau}(X_{i}).$ (8)
The above discussion shows that any censoring design producing large weights will both weaken the PAC‐type coverage bound (through a large supremum weight $\gamma_{\tau}$) and increase the variance of the empirical miscoverage estimator (through a large mean weight $\bar{w}_{\tau}$). Indeed, these two properties guide our Optimized censoring design presented in the following section, where we both (i) cap $\gamma_{\tau}$ by a constant value; and (ii) minimize the mean weight $\bar{w}_{\tau}$ subject to fully utilizing the sampling-allocation budget $B$.
5 Proposed budget-allocation mechanism designs
Having formulated the general calibration scheme and the factors influencing its coverage guarantee, we now introduce a sequence of baseline sampling-budget allocation strategies—Naive, Basic, and Trimmed—that build up to our flagship Optimized approach.
5.1 Naive mode
As a warm-up, we introduce a simple, baseline budget allocation strategy, the Naive approach, which defines each censoring time $C_{i}$ as a random variable independent of the prompt $X_{i}$. Assuming $T_{i}|X_{i}$ is a Geometric random variable with infinite support, we design $C_{i}$ to be distributed geometrically as well: $C_{i}\sim\mathrm{Geom}(\min(|{\mathcal{I}_{2}}|/B,1)),\ \forall i\in{\mathcal{I}_{2}}$, which satisfies the budget constraint in (2) by design. This Naive approach is summarized in Algorithm 3 in Appendix B.
Under this Naive allocation, the event $\hat{q}_{\tau}(X_{i})\leq C_{i}$ can be rare, especially for large $\hat{q}_{\tau}(X_{i})$. This makes $\bar{w}_{\tau}$ and $\gamma_{\tau}$ excessively large. Consequently, the resulting coverage rates might deviate from the nominal level—as reflected in our experiments in Section 6. This discussion motivates the method presented below, which utilizes the budget more efficiently and designs the censoring time as a function of the prompt.
5.2 Basic mode
To better utilize the calibration budget in the computation of $\hat{\alpha}(\tau)$ in (6), we aim to design prompt-dependent censoring times $C_{i}$ that maximize $\mathbb{P}\bigl(\hat{q}_{\tau}(X_{i})\leq C_{i}\bigr)$. Ideally, we would like to set the censoring times for each $\tau$ to increase this probability. However, this is infeasible as we construct the dataset only once. As a result, we must commit to a single budget allocation by choosing a specific quantile level $\tau$ to consider. For this purpose, we assume that the quantile estimators are monotone in $\tau$. This assumption can be enforced by sorting $\hat{q}_{\tau}(X_{i})$. Consequently, $\forall\tau_{1}\leq\tau_{2},\ \mathbb{I}\{\hat{q}_{\tau_{2}}(X_{i})\leq C_{i}\}\Rightarrow\mathbb{I}\{\hat{q}_{\tau_{1}}(X_{i})\leq C_{i}\}.$ Thus, if we were to naively set $C_{i}=\hat{q}_{1}(X_{i})$ as the 100%–quantile, we would indeed maximize $\mathbb{P}\bigl(\hat{q}_{\tau}(X_{i})\leq C_{i}\bigr)$ for all $\tau\in[0,1]$. However, $\hat{q}_{1}(X_{i})$ can be excessively large, e.g., this 100%-quantile is infinite if $T_{i}\mid X_{i}$ follows a geometric distribution. As a result, this naive choice is infeasible under our finite budget constraint $\mathbb{E}\left[\sum_{i\in{\mathcal{I}_{2}}}C_{i}|\{X_{i}\}_{i\in{\mathcal{I}_{2}}}\right]\leq B$.
Instead, we consider a more practical compromise: we select a quantile level $\tau_{\text{prior}}\ll 1$ that represents our prior belief $\hat{\tau}\in[0,\tau_{\text{prior}}]$. For example, if the target coverage level $1-\alpha$ is not too extreme (e.g., $90\%$) and the uncalibrated quantile estimator $\hat{q}_{\tau}$ is reasonably accurate, a prior level such as $\tau_{\text{prior}}=20\%$ would be a suitable choice. Importantly, constraining $\tau$ to lie within this prior range affects only the informativeness of the LPB, and not its validity; see Appendix D.6.
After choosing the prior level $\tau_{\textup{prior}}$, we turn to set a prompt-adaptive censoring time $C_{i}$ using the estimated quantile $\hat{q}_{\tau_{\textup{prior}}}(X_{i})$. Following (6), the $i$-th sample contributes to $\hat{\alpha}(\tau)$ if $C_{i}\leq\hat{q}_{\tau_{\textup{prior}}}(X_{i})$ for all $\tau\in[0,\tau_{\textup{prior}}]$. Combining this observation with the monotonicity of $\hat{q}_{\tau}$, a sensible choice for $C_{i}$ that avoids wasted samplings would either be (i) $C_{i}=\hat{q}_{\tau_{\text{prior}}}(X_{i})$, or (ii) $C_{i}=0$. Other choices for $C_{i}$ that are strictly greater or smaller than $\hat{q}_{\tau_{\text{prior}}}(X_{i})$ would result in unnecessary generations that do not contribute to the miscoverage estimator $\hat{\alpha}(\tau)$. Motivated by this observation, we define $C_{i}:=\text{Ber}(\pi_{i})\cdot\hat{q}_{\tau_{\text{prior}}}(X_{i})$, where $\text{Ber}(\pi_{i})$ is a Bernoulli random variable with success probability $\pi_{i}=\min({B}/({|{\mathcal{I}_{2}}|\cdot\hat{q}_{\tau_{\text{prior}}}(X_{i})}),1).$ This design of $C_{i}$ satisfies the budget constraint (2).
Once all $C_{i}$-s are set, we can use our calibration procedure from Section 4, with the sole modification that the search space for $\tau$ is now $\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]$. We remark that the per‐prompt evaluation probability $\pi_{i}$ does not depend on $\tau$, and so the inverse‐censoring weights become $w_{\tau}\bigl(X_{i}\bigr)=1/{\pi_{i}},$ for all $\tau\in\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]$. Hence, we omit the dependence on $\tau$ in our notation and denote the weights as $w\bigl(X_{i}\bigr)$, the maximum weight by $\gamma$, and the average weight by $\bar{w}$.
For reference, Algorithm 3 summarizes the calibration procedure using the above Basic budget allocation scheme. As with the Naive approach, the coverage bound for this method (Theorem 4.1) tightens as $\gamma$ and $\bar{w}$ decrease. In the following section, we explicitly control the value of $\gamma$ by trimming the output of $\hat{q}_{\tau}(x)$.
5.3 Trimmed mode
To limit the value of $\gamma$, we simply cap each estimated quantile at a fixed threshold $M\in\mathbb{N}$, defining $\hat{f}_{\tau}(x):=\min(\hat{q}_{\tau}(x),M)$. Consequently, the censoring times are now given by $C_{i}:=\text{Ber}(\pi_{i})\cdot\hat{f}_{\tau_{\text{prior}}}(X_{i}),$ where
$\pi_{i}=\min({B}/({|{\mathcal{I}_{2}}|\cdot\hat{f}_{\tau_{\text{prior}}}(X_{i})}),1).$ (9)
The rest of the calibration process continues as outlined in Section 4, using the trimmed estimates $\hat{f}_{\tau}(X_{i})$ instead of $\hat{q}_{\tau}(X_{i})$; See Algorithm 3.222Note that now $\mathcal{T}=\bigl\{\sup_{\tau\in[0,1]}\{\hat{f}_{\tau}(X_{i})\leq\tilde{T}_{i}\}:i\in{\mathcal{I}_{2}}\}\cup\{\sup_{\tau\in[0,1]}\{\hat{f}_{\tau}(X_{i})\leq C_{i}\}:{i\in{\mathcal{I}_{2}}}\bigr\}\cup\{0\}$ and the search space for $\tau$ is $\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]$.
Before proceeding to the Optimized method, we pause to discuss the implications of the threshold $M$ on the validity and statistical efficiency of the resulting LPBs. First, since $\hat{f}_{\tau_{\text{prior}}}(X_{i})\leq M$, then $\pi_{i}$ from (9) satisfies $\pi_{i}\leq\min\left({B}/({|{\mathcal{I}_{2}}|\cdot M}),1\right)$, and so the maximum weight $w(x)$ is bounded by $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right)$. As observed in Theorem 4.1, this allows us to obtain a coverage guarantee that is closer to the desired $1-\alpha$ level. Second, the expression $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right)$ offers guidance for choosing the sampling budget $B$: it highlights the trade-off between (i) the tightness of the coverage guarantee and (ii) the statistical efficiency of the resulting LPB. Specifically, one can increase $M$ while ensuring the value of $\gamma$ remains at a certain level (e.g., $\gamma=2$) by demanding a higher sampling budget $B$. Indeed, our experiments show that when the budget permits setting a sufficiently large $M$, trimming has a small effect on the power of the resulting LPBs. Third, from a practical perspective, trimming the quantile estimates limits the maximal number of LLM samplings per prompt, preventing any single prompt from demanding an infeasible computational or audit workload.
5.4 Optimized mode: our flagship method
Having bounded $\gamma$, our next objective is to minimize the average weight $\bar{w}=\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}\frac{1}{\pi_{i}}$ in order to reduce the variance of $\hat{\alpha}(\tau)$. Recall the budget constraint in (2). For $C_{i}:=\text{Ber}(\pi_{i})\cdot\hat{f}_{\tau_{\text{prior}}}(X_{i})$, we get $\mathbb{E}[\sum_{i\in\mathcal{I}_{2}}C_{i}|\{X_{i}\}_{i\in{\mathcal{I}_{2}}}]=\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi_{i}$. In turn, we can minimize $\bar{w}$ subject to the budget constraint by solving the following convex optimization problem:
$\pi^{*}=\underset{{\pi\in[0,1]^{|{\mathcal{I}_{2}}|}}}{\rm{argmin}}\;\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}\frac{1}{\pi_{i}}\ \ \text{s.t.}\ \ \sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi_{i}\!\leq\!B.$ (10)
In Appendix B, we introduce Algorithm 4, which efficiently solves the above optimization problem. Proposition B.2 ensures this algorithm returns a unique and strictly positive solution $\pi^{*}$ that satisfies the budget constraint (2). Importantly, because the solution $\pi^{*}$ depends on the entire set of calibration prompts $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$, the resulting weights inherit this dependence. To make this explicit, we denote the weight associated with the $i$-th example as $w(\{X_{j}\}_{j\in\mathcal{I}_{2}},i)=1/\pi_{i}^{*}$. We remark that our coverage guarantee from Theorem 4.1 holds also for this formulation of the weights; see Appendix D.6 for details. Furthermore, these weights are also guaranteed to be upper-bounded by the same expression for $\gamma$ as in the Trimmed method.
Proposition 5.1 (Maximal weight bound).
Suppose that $\max_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\leq M$, then, the weights
$w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)$ induced by solving (10) are upper bounded by $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right).$
All proofs are deferred to Appendix D. The above proposition shows us that the solution $\pi^{*}$ inherits the advantages of the Trimmed method and further improves upon it by achieving lower variance, as it minimizes $\bar{w}$. The Optimized calibration scheme is outlined in Algorithm 3.
6 Experiments
We evaluate our methods on both synthetic (Section 6.1) and real (Section 6.2) datasets. For a full description of how we train models to estimate the conditional quantiles of $T$ given $X$, along with additional experiments, implementation details, and runtime considerations for our calibration methods, see Appendix E. Since the problem setting introduced in this work is novel, there are no established baseline methods against which we can directly compare our performance. Therefore, we consider the vanilla, uncalibrated quantile regression model, as well as the calibrated Naive, Basic, and Trimmed methods as competitive baselines for our Optimized method. The following experiments indicate that our Optimized budget allocation leads to more informative LPBs that attain a coverage rate closer to the desired level compared to these baseline alternatives.
6.1 Synthetic data experiments
Figure 1: Synthetic experiments. Left: Coverage (target 90%). Center: Mean number of samplings generated per prompt. Right: Mean LPB. Shaded regions represent the standard deviation over 20 runs.
We generate a dataset that simulates a high-risk setting with predominantly prompts that can yield unsafe outcomes. To this end, we generate n=100,000n=100{,}000 pairs $(X_{i},p_{i})$, where $p_{i}$ is the true probability of sampling an unsafe $Y_{i}$ for $X_{i}\in\mathbb{R}^{d}$, with $d=10$. Specifically, $X_{i}\mid p_{i},$ follows a normal distribution, defined in Appendix E.1. The unsafe probability $p$ is chosen to simulate a dataset of “suspicious” prompts, where 90% of which have a high probability for unsafe generation , with the remaining 10% of prompts having a low probability for such an event. Then, we draw $T_{i}\sim\mathrm{Geom}(p_{i})$. We randomly partition the data into training (45%), calibration (45%), and test (10%) sets. For each example in the training set, we generate $500$ outputs. We then use this training set to fit the quantile estimators $\hat{q}_{\tau}(x)$.
We compare (1) the Uncalibrated baseline, the raw quantile estimate $\hat{q}_{\tau}(X)$ at quantile level $\tau=\alpha$; (2) the Naive calibration method, which is a slight variant of the method from Section 5.1. The original Naive approach tended to be unstable, often yielding trivial LPBs, since $\hat{q}_{\tau}(X_{i})\leq C_{i}$ holds for very few points at high quantile levels and these samples do not contribute to the miscoverage estimator in (6). To mitigate this, we restrict the $\tau$’s search space in Algorithm 3 to $\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]$, as in the adaptive budget allocation methods. We also include (3) the Basic method from Section 5.2; (4) its Trimmed version from Section 5.3; and (5) our flagship Optimized approach from Section 5.4.
Recall that the calibration procedures are inherently random due to the sampling of $C_{i}$ and the random generations obtained by the model $\mathcal{G}(x)$. Ideally, this randomness should have a small effect on our bounds. To quantify this randomness, we fix the data split, and run each method for $J=20$ random draws of censoring and time-to-unsafe-sampling times. We index each run by $j$, and denote the resulting LPB by $\hat{q}_{\hat{\tau}}^{(j)}(X_{i})$ and compute:
$\mathrm{AvgCoverage}^{(j)}=\frac{1}{|\mathcal{I}_{\mathrm{test}}|}\sum_{i\in\mathcal{I}_{\mathrm{test}}}\mathbb{P}\{T_{i}\leq$ $\hat{q}_{\hat{\tau}}^{(j)}(X_{i})|X_{i}\},$
$\mathrm{AvgLPB}^{(j)}=\frac{1}{|\mathcal{I}_{\mathrm{test}}|}\sum_{i\in\mathcal{I}_{\mathrm{test}}}\hat{q}_{\hat{\tau}}^{(j)}(X_{i}),\mathrm{AvgBudget}^{(j)}$ $=\frac{1}{|\mathcal{I}_{\mathrm{test}}|}\sum_{i\in\mathcal{I}_{\mathrm{test}}}C_{i}^{(j)}.$
Above, $\mathcal{I}_{\mathrm{test}}$ are the indices of the test points. We report the average and standard deviation of the above quantities over $J$ runs.
Results. We study the effect of the average budget per prompt $B/|{\mathcal{I}_{2}}|$ on the coverage and mean LPB obtained by each method. Following Figure 1, we can see that the Uncalibrated method does not attain valid coverage. The Naive method severely undercovers $T$ for small budgets, and overcovers $T$ for larger budgets. The Basic method attains more stable coverage around the desired $90\%$ level under medium-to-high budget, but exhibits high variance. By contrast, the Trimmed method attains nearly $90\%$ coverage across all budget constraints. Our flagship Optimized calibration method attains tight 90% coverage with minimal variance compared to all other methods. The middle panel highlights that both the Naive and Optimized approaches fully utilize the budget; however, the latter is tightly regulated around the desired coverage—better utilizing the sampling budget. The right panel demonstrates that the Trimmed and Optimized methods produce stable LPBs with better statistical efficiency as the average budget per prompt increases. The increased budget allows us to set higher quantile trim values $M$ while enforcing low $\gamma$. As the budget increases, the two adaptive methods converge to the performance of the Basic technique. However, this technique yields higher LPBs with a higher coverage variability. Additional experiments, including oracle models, varying maximal weights, $\alpha$ levels, and calibration/test splits, are in Appendix E.1.
6.2 Real data experiments
Figure 2: RealToxicityPrompts dataset experiment. Left: Empirical coverage rate (target 90%). The true coverage of the Optimized method is in a solid red line, while the upper and lower bounds on the coverage of the Uncalibrated and Naive methods correspond to the dotted lines. Right: Mean LPB. Shaded regions represent the standard deviation, computed over $5$ runs. Higher is better.
We now evaluate the effectiveness of our Optimized calibration method on real-world data. For this purpose, we use the RealToxicityPrompts dataset (Gehman et al., 2020), which contains 99,442 sentences designed to measure model toxicity. Each sentence is divided into two halves: the first half serves as the input prompt $X_{i}$ to a language model, which must then complete the sentence. In this experiment, we generate about 30 million model responses. Because this scale is computationally demanding, we employ a relatively small model, Llama 3.2 1B (Meta AI, 2024, Grattafiori et al., 2024). Furthermore, as we do not have access to human annotations, we rely on the Detoxify-original classifier (Hanu & team, 2020) to label each completion as safe (non-toxic) or unsafe (toxic), using a toxicity score threshold of 0.5.
We randomly split the dataset into training (50%), validation (10%), calibration (20%), and test (20%) sets. On the training set, we sample $N=500$ responses per prompt and fit a quantile estimator $\hat{q}_{\tau}(x)$ as detailed in Appendix E. We employ the three calibration methods described in Section 6.1: (1) the Uncalibrated baseline, (2) the Naive method, and (3) the Optimized method, all aiming to attain 90% coverage rate.
In all experiments, we report the empirical mean of the LPBs over the test set. We also estimate the empirical miscoverage by drawing $\min\bigl(\hat{L}(X_{i}),2400\bigr)$ samples for each prompt. We cap the maximal LLM generations on the test set at 2,400 samplings per prompt due to computational constraints. Since the Optimized method trims the quantiles at $M<2400$ across all evaluated budget levels, by drawing exactly $\hat{L}(X_{i})$ samples (with $\hat{L}(X_{i})\leq M$) we can compute the miscoverage rate directly. Since $\hat{L}(X_{i})$ is unbounded for the Uncalibrated and Naive methods, we can only report a lower bound on the empirical coverage rate using the 2,400 generations; see also (Gui et al., 2024). In turn, the empirical lower and upper bounds are given by $\frac{1}{|\mathcal{I}_{\mathrm{test}}|}\sum_{i\in\mathcal{I}_{\mathrm{test}}}\mathbb{I}\bigl\{\hat{L}(X_{i})\leq\tilde{T}_{i}\bigr\},$ and $\frac{1}{|\mathcal{I}_{\mathrm{test}}|}\sum_{i\in\mathcal{I}_{\mathrm{test}}}\mathbb{I}\bigl\{\min(\hat{L}(X_{i}),2400)\leq\tilde{T}_{i}\bigr\}$, respectively. This evaluation protocol offers a fair comparison under relatively limited computations. To measure the variability introduced by the inherently stochastic calibration procedures, we use the same fixed data split for each run, as in our synthetic experiments.
Results. Figure 2 presents the results across $J=5$ runs. The Uncalibrated baseline overcovers the time-to-unsafe-sampling, resulting in overly conservative LPBs. The Naive method, on the other hand, generates invalid LPBs with high variance at small budget levels. In contrast, our Optimized method consistently attains near‐nominal coverage rates with low variance. Notably, the LPBs produced by the Optimized method increase as the budget increases, approaching the values of the Naive and surpassing those of the Uncalibrated method. Overall, this experiment shows that the Optimized method produces LPBs that are both valid across all tested budgets, and become more informative as the budget grows.
7 Discussion
We introduced a novel approach to construct a calibrated LPB for time-to-unsafe-sampling of a given prompt, casting this task as a survival analysis problem. Key to our method is the design of an optimized censoring mechanism that utilizes a given budget constraint on the number of LLM generations. One limitation of our approach is that the coverage guarantee is marginal and does not hold conditionally on a specific subgroup. Consequently, one cannot group prompts with especially low (or high) LPBs and assume that those bounds retain their nominal coverage when conditioned on this selection. In future work, we plan to extend our approach to this conditional inference setting, leveraging the methods by Sesia & Svetnik (2025), Jin & Ren (2025) as a foundation for achieving valid group-conditional coverage. Another limitation is that our calibration assumes the samples are drawn i.i.d. Even in benign real-world setups, the prompt distribution can shift, potentially undermining the validity of our method. In future research, we aim to propose continual or adaptive recalibration mechanisms that adjust to evolving user behavior and emergent risks, possibly by drawing on ideas from (Gibbs & Candès, 2024).
Appendix A Additional related work
The methods introduced in our paper draw on ideas from survival analysis, conformal prediction, and black-box uncertainty quantification. The Optimized method (Section 5.4) in particular is also related to: conformal prediction foundations, uncertainty quantification for LLMs, active learning and statistical inference, and optimal experimental design.
Conformal prediction
The theoretical foundations of conformal prediction were established by Kolmogorov (1983; 1968), Papadopoulos et al. (2002), Vovk et al. (2005). This framework has since been extended to accommodate covariate shift by applying weights to the distribution of conformity scores (Tibshirani et al., 2019). Similar weighting strategies have also been employed to integrate active learning into conformal prediction (Prinster et al., 2024). Building further on these weighted approaches, conformal prediction methods have recently been adapted for survival analysis tasks (Candès et al., 2023, Gui et al., 2024, Davidov et al., 2025).
Uncertainty quantification for LLMs
Several recent works adapt conformal ideas to off-the-shelf LLMs. For instance, Wang et al. (2024) adapts conformal prediction to open‐ended generation by integrating a self–consistency–based uncertainty measure, achieving strict correctness coverage across seven popular LLMs and diverse domains. The work of Wang et al. (2025) further extends this by explicitly inserting the ground truth into candidate outputs and using logit‐based nonconformity scores to guarantee coverage over a wider range of error rates, though its reliance on logits raises calibration concerns when only API‐level access is available. Additionally, Conformal Language Modeling has been developed by Quach et al. (2024) to calibrate both stopping and rejection rules to guarantee coverage in open-domain QA and summarization. Lin et al. (2024) distinguish uncertainty from confidence using semantic dispersion metrics, enabling selective generation of low-confidence outputs. To refine validity under practical constraints, Cherian et al. (2024) extended conditional conformal methods to incorporate utility-based guarantees and differentiable filtering. Mohri & Hashimoto (2024) introduced “conformal factuality,” constructing entailment-based uncertainty sets that deliver high-probability correctness. The approach proposed in (Gui et al., 2024) provides a guarantee for aligned responses in QA and radiology applications.
Active learning and statistical inference
Active learning methods aim to select the most informative samples for labeling, improving model accuracy under tight budgets (Settles, 2010). This was formalized in (Zrnic & Candès, 2024) for statistical inference tasks, where one prioritizes labeling points with high model uncertainty while maintaining valid confidence intervals. Our Optimized method relates to this idea as it derives sample probabilities that directly minimize the variance of the augmented inverse-propensity-weighting estimator.
Optimal experimental design
Optimal experimental design allocates observations to minimize estimator variance under a given model. In (Hahn et al., 2011), the authors develop a two-stage adaptive design that chooses propensity scores to minimize the asymptotic variance bound in treatment-effect estimation, and the approach presented in (List et al., 2011) presents simple rules of thumb for efficient designs under resource constraints. Analogously, our Optimized method treats the sampling rule $\pi(x)$ as a design variable and solves for the allocation that minimizes the miscoverage estimation variance.
Classical survival analysis
The survival analysis problem was first formalized in biostatistics and reliability engineering. The work in (Kaplan & Meier, 1958) introduces a nonparametric estimator of the survival function for right‐censored data, laying the groundwork for all subsequent methods. The proportional hazards model was later developed, formalizing a semiparametric regression framework that relates covariates to an event’s hazard rate without specifying its baseline form. As an alternative formulation, accelerated failure time (AFT) (Lin et al., 1998) models assume a parametric form for the log‐survival time, allowing direct modeling of time-to-event via common distributions such as Weibull or log-normal. More recently, machine-learning approaches have been developed for survival prediction. Random survival forests leverage ensemble tree methods to nonparametrically estimate cumulative hazard functions (Ishwaran et al., 2008). DeepSurv employs a neural-network approximation of the Cox partial likelihood to capture complex covariate interactions (Katzman et al., 2018), and Deep Cox Mixtures generalizes this by modeling mixtures of proportional hazards to flexibly adapt to heterogeneous subpopulations (Nagpal et al., 2021).
Appendix B Calibration algorithms
In this section, we detail the optimized adaptive budgeting algorithm. First, observe that if the budget is sufficiently large such that the sum of $\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})$ does not exceed the budget, i.e., $\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\leq B$, then we can set the censoring times to the target value $C_{i}=\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})$ for all calibration points. In this case, the solution for (10) is straightforward, and the optimized probabilities are equal to one: $\pi_{i}^{*}=1,\ \forall i\in{\mathcal{I}_{2}}$. Since this setup is trivial, we will consider the more challenging setting, in which the target values exceed the threshold $\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})>B$.
Since each summand $1/\pi_{i}$ is strictly convex in $\pi_{i}$ on $(0,1]$ and the constraint is linear, (10) is a convex program. Furthermore, if the sum of $\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})$ exceeds the budget, the constraint turns into an equality constraint, as stated next.
Proposition B.1.
Assume $\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})>B$. Then any minimizer $\pi^{*}$ of (10) must satisfy the budget constraint with equality:
$\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi^{*}_{i}\;=\;B.$
We now turn to show how to compute the optimized probabilities $\pi^{*}$. Under Proposition B.1, the optimization constraint in (10) turns into an equality constraint, and the Lagrangian is therefore given by
$\mathcal{L}(\pi,\lambda)=\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}\frac{1}{\pi_{i}}+\lambda\Bigl(\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi_{i}-B\Bigr),$
where $\lambda>0$ is the Lagrange multiplier. On the open domain $(0,1]^{n}$, the objective is strictly convex and differentiable, so the single stationary point satisfying the first‐order conditions is in fact the unique global minimizer. Taking partial derivatives,
$\frac{\partial\mathcal{L}}{\partial\pi_{i}}=-\frac{1}{|{\mathcal{I}_{2}}|\,\pi_{i}^{2}}+\lambda\,\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\;=\;0\quad\Longrightarrow\quad\pi_{i}=\frac{1}{\sqrt{|{\mathcal{I}_{2}}|\,\lambda\,\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})}}.$
Applying the box constraint $\pi_{i}\leq 1$ gives
$\pi_{i}^{*}(\lambda)=\min\Bigl\{1,\;\frac{1}{\sqrt{|{\mathcal{I}_{2}}|\,\lambda\,\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})}}\Bigr\}.$
Now, define the budget‐usage function
$U(\lambda)\;=\;\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi_{i}^{*}(\lambda).$
The above equation has a unique solution $\lambda^{*}>0$ which satisfies $U(\lambda^{*})=B$. This $\lambda^{*}>0$ can be recovered by bisection, across the following interval:
$\left[\lambda_{\textup{low}}=\;\frac{1}{|{\mathcal{I}_{2}}|\max_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})},\ \lambda_{\textup{high}}=\frac{|{\mathcal{I}_{2}}|\max_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})^{2}}{B^{2}\min_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})}\right].$ (11)
This procedure is summarized in Algorithm 4 and it produces a valid solution, as stated in the following proposition.
Proposition B.2.
The output $\pi^{*}$ of Algorithm 4 is the unique solution of (10) which satisfies $\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi^{*}_{i}\!\leq\!B$.
Algorithm 1 Generate Calibration Responses
0: Calibration data $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$, Generative model $\mathcal{G}(\cdot)$, audit function $\texttt{Audit}(\cdot)$, Censoring times $\{C_{i}\}_{i\in{\mathcal{I}_{2}}}$.
1:for each $i\in\mathcal{I}$ do
2:  Initialize $j\leftarrow 0$
3:  repeat
4:   $j\leftarrow j+1$
5:   Generate and evaluate the output’s safety: $Y_{i}^{j}\leftarrow\texttt{Audit}(X_{i},\mathcal{G}(X_{i}))$
6:  until $Y_{i}^{j}=1$ or $C_{i}=j$
7:  Set $\tilde{T}_{i}\leftarrow j$
8:end for
9:return $\{\tilde{T}_{i}\}_{i\in\mathcal{I}_{2}}$
Algorithm 2 Generate Calibration Outcomes
0: Calibration data $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$, generative model $\mathcal{G}(\cdot)$, max per-prompt budget $\{\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\}_{i\in\mathcal{I}}$, per-prompt evaluation probability $\{\pi_{i}\}_{i\in\mathcal{I}}$, audit function $\texttt{Audit}(\cdot)$.
1:for each $i\in\mathcal{I}$ do
2:  Draw $V_{i}\sim\mathrm{Bernoulli}(\pi_{i})$
3:  Set $C_{i}\leftarrow V_{i}\cdot\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})$
4:end for
5: Obtain $\{\tilde{T}_{i}\}_{\{i\in\mathcal{I}_{2}\}}$ from Algorithm 1 applied with $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$, $\mathcal{G}$, $\{C_{i}\}_{i\in{\mathcal{I}_{2}}}$.
6:return $\{(\tilde{T}_{i},C_{i})\}_{i\in\mathcal{I}_{2}}$.
Algorithm 3 Calibration with a known censoring mechanism
0: Calibration data $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$, generative model $\mathcal{G}(\cdot)$, audit function $\texttt{Audit}(\cdot)$, pre-trained quantile regression model $\{\hat{q}_{\tau}(\cdot)\}_{\tau\in\mathcal{T}}$, target miscoverage rate $\alpha$}, prior quantile $\tau_{\textup{prior}}$}, calibration mode (Naive, Basic, Trimmed, Optimized), quantile trimming threshold $M$, total budget $B$.
1:if calibration mode is Naive then
2:  Sample $C_{i}\sim\mathrm{Geom}(n/B)$ for all $i\in{\mathcal{I}_{2}}$}
3:  Obtain $\{\tilde{T}_{i}\}_{\{i\in\mathcal{I}_{2}\}}$ from Algorithm 1 applied with $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$, $\mathcal{G}$, $\{C_{i}\}_{i\in{\mathcal{I}_{2}}}$.
4:  $w_{\tau}(X_{i})\leftarrow\frac{1}{\mathbb{P}[\hat{q}_{\tau}(X_{i})\leq C_{i}|X_{i}]},\ i\in{\mathcal{I}_{2}}$. // compute the weights for the geometric $C_{i}$.
5:else
6:  if calibration mode is Basic then
7:   $\{\hat{f}_{\tau}(X_{i})\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}\leftarrow\{\hat{q}_{\tau}(X_{i})\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}$
8:   $\{\pi_{i}\}\leftarrow\{\min(\tfrac{B}{|{\mathcal{I}_{2}}|\;\hat{f}_{\tau_{\textup{prior}}}(X_{i})},1)\}$ // per-prompt evaluation probability (Section 5.2)
9:  else if calibration mode is Trimmed then
10:   $\{\hat{f}_{\tau}(X_{i})\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}\leftarrow\{\min{(\hat{q}_{\tau}(X_{i}),M)}\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}$ // trim the quantile est. (Section 5.3)
11:   $\{\pi_{i}\}\leftarrow\{\min(\tfrac{B}{|{\mathcal{I}_{2}}|\;\hat{f}_{\tau_{\textup{prior}}}(X_{i})},1)\}$ // per-prompt evaluation probability (Section 5.2)
12:  else if calibration mode is Optimized then
13:   $\{\hat{f}_{\tau}(X_{i})\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}\leftarrow\{\min{(\hat{q}_{\tau}(X_{i}),M)}\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}$ // trim the quantile est. (Section 5.3)
14:   $\{\pi_{i}\}\leftarrow$ Algorithm 4 applied with $\{\hat{f}_{\tau_{\textup{prior}}}(X_{i})\}_{i\in{\mathcal{I}_{2}}}$ // optimized per-prompt evaluation                       probability (Section 5.4)
15:  end if
16:  $\{(\tilde{T}_{i},C_{i})\}_{i\in{\mathcal{I}_{2}}}\leftarrow$ Algorithm 2 applied to $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$ with $\{\pi_{i}\}_{i\in{\mathcal{I}_{2}}}$ and $\{\hat{f}_{\tau_{\textup{prior}}}(X_{i})\}_{i\in{\mathcal{I}_{2}}}$
17:  $w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\leftarrow\frac{1}{\pi_{i}},\ i\in{\mathcal{I}_{2}}$
18:end if
19:for $\tau\in\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]$ do
20:  $\hat{\alpha}(\tau)\leftarrow\displaystyle\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\;\mathbb{I}\bigl\{\tilde{T}_{i}<\hat{f}_{\tau}(X_{i})\leq C_{i}\bigr\}$ // miscoverage est.
21:end for
[…]
23:return Lower predictive bound (LPB) for a test point $X_{\textup{test}}=x$, given by $\hat{L}(x)=\hat{f}_{\hat{\tau}}(x)$
Algorithm 4 Get Optimal Per-Prompt Evaluation Probabilities
0: effective sample cost $\{\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\}_{i\in\mathcal{I}_{2}}$, budget $B$, bijection tolerance $\epsilon$, initial $\lambda_{\textup{l}ow},\lambda_{\textup{h}igh}$
0: Optimal $\pi^{*}$ and multiplier $\lambda^{*}$ satisfying $\pi^{*}_{i}=\min\{1,1/\sqrt{\lambda^{*}\,\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})}\}$ and $\sum_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi^{*}_{i}=B$.
1:if $B>\sum_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})$ then
2:  $\pi^{*}\leftarrow[1,\dots,1]$, $\lambda^{*}\leftarrow\texttt{Null}$
3:else
4:  while $\sum_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi_{i}^{*}(\lambda_{\textup{h}igh})>B$ do
5:   $\lambda_{\textup{high}}\leftarrow 2\,\lambda_{\textup{high}}$
6:  end while
7:  $\lambda^{*}\leftarrow$ bisection on $\lambda_{\textup{low}},\lambda_{\textup{high}}]$ to solve $\sum_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi_{i}^{*}(\lambda)=B$ within $\epsilon$
8:  $\pi^{*}_{i}\leftarrow\min\{1,1/\sqrt{\lambda^{*}\,\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})}\}$ for all $i\in{\mathcal{I}_{2}}$
9:end if
10:return $(\pi^{*},\lambda^{*})$
Appendix C A deeper look into $\hat{\alpha}(\tau)$
In Section 4, we introduced the miscoverage estimator $\hat{\alpha}(\tau)$. In this section, we explore its statistical properties in detail. We begin by deriving its expectation and variance:
Proposition C.1 (Unbiasedness and conditional variance of the weighted miscoverage estimator).
Under the conditional independence assumption 2.1 in Section 2, the estimator $\hat{\alpha}(\tau)$ defined in equation (6) satisfies
$\mathbb{E}\bigl[\hat{\alpha}(\tau)\bigr]=\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\mathbb{P}\bigl[T_{i}<\hat{q}_{\tau}(X_{i})\bigr].$
If we further assume that for all $i\neq i^{\prime}\in{\mathcal{I}_{2}}$, $(C_{i},T_{i}){\perp\!\!\!\perp}(C_{i^{\prime}},T_{i^{\prime}})\mid\{X_{j}\}_{j\in\mathcal{I}_{2}}$, we have
$\mathrm{Var}\bigl[\hat{\alpha}(\tau)\,\bigm|\,\{X_{j}\}_{j\in\mathcal{I}_{2}}\bigr]=\frac{1}{|\mathcal{I}_{2}|^{2}}\sum_{i\in\mathcal{I}_{2}}\Bigl\{w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)$ $\,\mathbb{P}\bigl[T_{i}<\hat{q}_{\tau}(X_{i})\,\bigm|\;\{X_{j}\}_{j\in\mathcal{I}_{2}}\bigr]$
$-\;\bigl[\mathbb{P}\bigl[T_{i}<\hat{q}_{\tau}(X_{i})\,\bigm|\;\{X_{j}\}_{j\in\mathcal{I}_{2}}\bigr]\bigr]^{2}\Bigr\}.$
See Appendix D.1 for the proof. We remark that the conditional independence and pair-wise independence assumptions in Proposition C.1 hold for all variants of our method from Sections 5. This proposition shows that $\hat{\alpha}(\tau)$ is an unbiased miscoverage estimator, resulting in a useful approximation for the miscoverage estimation variance. Under a simplified assumption that, for a fixed $\tau$, each calibration point $i\in{\mathcal{I}_{2}}$ is miscovered at the same constant rate conditional on $\{X_{j}\}_{j\in\mathcal{I}_{2}}$, we have the following result.
Proposition C.2 (Variance linearly increasing in mean weight under constant miscoverage).
Under Assumption 2.1, and additionally assuming that for all $i\neq i^{\prime}$, $(C_{i},T_{i}){\perp\!\!\!\perp}(C_{i^{\prime}},T_{i^{\prime}})\mid\{X_{j}\}_{j\in\mathcal{I}_{2}}$, and that for any fixed $\tau$, each calibration point $i\in{\mathcal{I}_{2}}$ is miscovered at the same constant rate conditional on $\{X_{j}\}_{j\in\mathcal{I}_{2}}$,
$\mathbb{P}\bigl[T_{i}<\hat{q}_{\tau}(X_{i})\,\bigm|\;\{X_{j}\}_{j\in\mathcal{I}_{2}}\bigr]=\mathrm{Const}_{\tau}.$
Then, the variance of $\hat{\alpha}(\tau)$ from (6) is
$\mathrm{Var}\bigl[\hat{\alpha}(\tau)\,\bigm|\,\{X_{j}\}_{j\in\mathcal{I}_{2}}\bigr]$ $=\frac{\mathrm{Const}_{\tau}}{|\mathcal{I}_{2}|}\overline{w}_{\tau}-\frac{\mathrm{Const}^{2}_{\tau}}{|\mathcal{I}_{2}|},$
where $\overline{w}_{\tau}=|\mathcal{I}_{2}|^{-1}\sum_{i\in\mathcal{I}_{2}}w_{\tau}(\{X_{j}\}_{j\in\mathcal{I}_{2}},i)$ is the mean weight.
See proof in Appendix D.2.
Appendix D Proofs
D.1 Proof of Proposition C.1
Proof.
Recall that
$\hat{\alpha}(\tau)=\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}w(X_{i})\,\mathbb{I}\{\hat{q}_{\tau}(X_{i})\leq C_{i}\}\,\mathbb{I}\{T_{i}<\hat{q}_{\tau}(X_{i})\}.$
Set
$V_{i}=w(X_{i})\,\mathbb{I}\{\hat{q}_{\tau}(X_{i})\leq C_{i}\}\,\mathbb{I}\{T_{i}<\hat{q}_{\tau}(X_{i})\}.$
By Assumption 2.1 each of $C_{i}$ and $T_{i}$ is independent of the other given $X_{i}$. Hence for each $i$,
$\mathbb{E}\bigl[V_{i}\mid\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr]$ $=\mathbb{E}\bigl[w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\,\mathbb{I}\{\hat{q}_{\tau}(X_{i})\leq C_{i}\}\,\mathbb{I}\{T_{i}<\hat{q}_{\tau}(X_{i})\}\,\bigm|\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr]$
$=\mathbb{E}\bigl[w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\,\mathbb{I}\{\hat{q}_{\tau}(X_{i})\leq C_{i}\}\,\bigm|\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr]\,\mathbb{E}\bigl[\mathbb{I}\{T_{i}<\hat{q}_{\tau}(X_{i})\}\,\bigm|\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr]$
$=w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\,\mathbb{P}\bigl[\hat{q}_{\tau}(X_{i})\leq C_{i}\,\bigm|\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr]\,\mathbb{P}\bigl[T_{i}<\hat{q}_{\tau}(X_{i})\,\bigm|\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr]$
$=\mathbb{P}\bigl[T_{i}<\hat{q}_{\tau}(X_{i})\,\bigm|\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr],$
where the first is by the tower property, and in the last transition we used $w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)=\mathbb{P}\bigl[\hat{q}_{\tau}(X_{i})\leq C_{i}\bigm|\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr]^{-1},$ from (5).
Consequently,
$\mathbb{E}[\hat{\alpha}(\tau)]$ $=\mathbb{E}\Bigl[\,\tfrac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}V_{i}\Bigr]$
$=\mathbb{E}\Bigl[\,\tfrac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\mathbb{E}\bigl[V_{i}\mid\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr]\Bigr]$
$=\mathbb{E}\Bigl[\,\tfrac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\mathbb{P}\bigl[T_{i}<\hat{q}_{\tau}(X_{i})\mid X_{i}\bigr]\Bigr]=\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\mathbb{P}\bigl[T_{i}<\hat{q}_{\tau}(X_{i})\bigr].$
For the conditional variance, we similarly derive
$\mathbb{E}\bigl[V_{i}^{2}\mid\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr]$ $=\mathbb{E}\bigl[w^{2}_{\tau}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\,\mathbb{I}\{\hat{q}_{\tau}(X_{i})\leq C_{i}\}\,\mathbb{I}\{T_{i}<\hat{q}_{\tau}(X_{i})\}\,\bigm|\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr]$
$=w^{2}_{\tau}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\,\mathbb{P}\bigl[\hat{q}_{\tau}(X_{i})\leq C_{i}\,\bigm|\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr]\,\mathbb{P}\bigl[T_{i}<\hat{q}_{\tau}(X_{i})\,\bigm|\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr]$
$=w_{\tau}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\mathbb{P}\bigl[T_{i}<\hat{q}_{\tau}(X_{i})\,\bigm|\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr].$
Therefore
$\mathrm{Var}\bigl[V_{i}\,\bigm|\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr]=$ $\mathbb{E}\bigl[V_{i}^{2}\,\bigm|\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr]-\bigl(\mathbb{E}[V_{i}\,\bigm|\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}]\bigr)^{2}$
$=$ $\,w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\,\mathbb{P}\bigl[T_{i}<\hat{q}_{\tau}(X_{i})\,\bigm|\;\{X_{j}\}\bigr]-\Bigl[\mathbb{P}\bigl[T_{i}<\hat{q}_{\tau}(X_{i})\,\bigm|\;\{X_{j}\}\bigr]\Bigr]^{2}.$
Conditioned on $\{X_{j}\}_{j\in\mathcal{I}_{2}}$, the pairs $(C_{i},T_{i})$ are independent across $i$, and so for $i\neq j$,
$\mathrm{Cov}(V_{i},V_{j}\mid\{X_{j}\})=0.$
Consequently,
$\mathrm{Var}\bigl[\hat{\alpha}(\tau)\,\bigm|\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr]=$ $\mathrm{Var}\Bigl[\tfrac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}V_{i}\,\Bigm|\;\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\Bigr]$
$=$ $\tfrac{1}{|\mathcal{I}_{2}|^{2}}\sum_{i\in\mathcal{I}_{2}}\Bigl\{w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\,\mathbb{P}\bigl[T_{i}<\hat{q}_{\tau}(X_{i})\,\bigm|\;\{X_{j}\}_{j\in\mathcal{I}_{2}}\bigr]$
$\qquad\qquad\qquad\qquad-\;\bigl[\mathbb{P}\bigl[T_{i}<\hat{q}_{\tau}(X_{i})\,\bigm|\;\{X_{j}\}_{j\in\mathcal{I}_{2}}\bigr]\bigr]^{2}\Bigr\}$
D.2 Proof of Proposition C.2
Proof of Proposition C.2.
Under the proposition’s assumptions, Proposition C.1 shows us that the variance of $\hat{\alpha}(\tau)$ simplifies to
$\mathrm{Var}\bigl[\hat{\alpha}(\tau)\,\bigm|\,\{X_{j}\}_{j\in\mathcal{I}_{2}}\bigr]$ $=\frac{1}{|\mathcal{I}_{2}|^{2}}\sum_{i\in\mathcal{I}_{2}}\Bigl\{w_{\tau}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\,\mathrm{Const}_{\tau}-\mathrm{Const}_{\tau}^{2}\Bigr\}$
$=\frac{\mathrm{Const}_{\tau}}{|\mathcal{I}_{2}|}\overline{w}_{\tau}-\frac{\mathrm{Const}^{2}_{\tau}}{|\mathcal{I}_{2}|}$
where $\overline{w}_{\tau}=|\mathcal{I}_{2}|^{-1}\sum_{i\in\mathcal{I}_{2}}w_{\tau}(\{X_{j}\}_{j\in\mathcal{I}_{2}},i)$ is the mean weight. ∎
D.3 Proof of Proposition B.1
Proof of Proposition B.1.
Since
$\frac{\partial}{\partial\pi_{i}}\Bigl(\frac{1}{|{\mathcal{I}_{2}}|}\sum_{j\in{\mathcal{I}_{2}}}\tfrac{1}{\pi_{j}}\Bigr)=-\frac{1}{|{\mathcal{I}_{2}}|\,\pi_{i}^{2}}<0,$
increasing any single $\pi_{i}$ (while keeping the others fixed) strictly decreases the objective. Assume for the sake of contradiction that there exist an optimal solution for (10), $\pi^{*}$, satisfying $\sum_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi^{*}_{i}<B$. Then, there exist an index $j\in{\mathcal{I}_{2}}$ and $\delta>0$ such that $\pi^{*}_{j}+\delta\leq 1$ and $\sum_{i}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\tilde{\pi}_{i}\leq B$, where $\tilde{\pi}_{i}=\pi^{*}_{i}$ for $i\neq j$ and $\tilde{\pi}_{j}=\pi^{*}_{j}+\delta$. But then
$\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in{\mathcal{I}_{2}}}\frac{1}{\tilde{\pi}_{i}}\;<\;\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in{\mathcal{I}_{2}}}\frac{1}{\pi^{*}_{i}},$
contradicting optimality of $\pi^{*}$. Therefore, the budget must bind:
$\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi^{*}_{i}=B.$
D.4 Proof of Proposition B.2
Proof of Proposition B.2.
Observe that if $\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\leq B$, then $\pi_{i}^{*}=1$ for all $i\in{\mathcal{I}_{2}}$ is the optimal solution which satisfies the constraint. We now turn to consider the case where $\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})>B$. We first show that there exists a unique solution $\lambda^{*}$. Observe that $U:(0,\infty)\to(0,\sum_{i}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})]$ is continuous and strictly decreasing, with $\lim_{\lambda\to 0^{+}}U(\lambda)=\sum_{i}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})>B$ and $\lim_{\lambda\to\infty}U(\lambda)=0$. By the intermediate‐value theorem there is a unique $\lambda^{*}>0$ such that $U(\lambda^{*})=B$. Next, the output of Algorithm 4 is indeed the solution $\lambda^{*}$ since $U$ is a monotone function, and thus the bisection approach is valid in this case. Therefore, $\pi^{*}(\lambda)$ is the correct optimal solution. ∎
D.5 Proof of Proposition 5.1
Proof of Proposition 5.1.
First, observe that by assuming $\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\leq B$, is follows that $\pi_{i}^{*}=1$ for all $i\in{\mathcal{I}_{2}}$ is the optimal solution which satisfies the constraint. We now turn to analyze the setup where $\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})>B$. We prove by induction over the sum $\sum_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})$, assuming that $\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\leq M\ \forall i\in{\mathcal{I}_{2}}$.
Base case. We begin with the maximal sum: $\sum_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})=|{\mathcal{I}_{2}}|M$. That is, $\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})=M$ for all $i\in\mathcal{I}_{2}$. Then, by the optimization constraint, we get
$M\sum_{i\in\mathcal{I}_{2}}\pi^{*}_{i}=B\qquad\Longrightarrow\qquad\sum_{i\in\mathcal{I}_{2}}\pi^{*}_{i}=\frac{B}{M},$
and by symmetry the unique minimizer of $\sum_{i}1/\pi^{*}_{i}$ with $\sum_{i}\pi^{*}_{i}=B/M$ is
$\pi^{*}_{i}=\frac{B}{n\,M}\quad\forall i\in{\mathcal{I}_{2}}.$
Inductive step. We now suppose that the claim holds for any weight vector with a sum $B-1<S\leq M|{\mathcal{I}_{2}}|$, and show that it holds for any vector with a sum S1S-1. Given a weight vector $\{\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\}_{i\in{\mathcal{I}_{2}}}$ with a sum $\sum_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})=S-1$, we compose a new vector with a sum $S$, as follows. In $\{\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\}_{i\in{\mathcal{I}_{2}}}$, there exists an index $i_{0}$ such that $\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})\leq M-1$. The new vector with a sum $S$ is defined by:
$\hat{f}^{\prime}_{{\tau_{\textup{prior}}}}(X_{i})=\begin{cases}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i}),&i\neq i_{0},\\ \hat{f}_{{\tau_{\textup{prior}}}}(X_{i})+1,&i=i_{0},\end{cases}$
Notice that $\max_{i}\hat{f}^{\prime}_{{\tau_{\textup{prior}}}}(X_{i})\leq M$ and $\sum_{i\in{\mathcal{I}_{2}}}\hat{f}^{\prime}_{{\tau_{\textup{prior}}}}(X_{i})=S$. Let $\lambda$ and $\lambda^{\prime}$ be the unique solutions of
$U(\lambda)=B\quad\text{and}\quad U^{\prime}(\lambda^{\prime})=B,$
where
$U^{\prime}(\lambda)=\sum_{i\neq i_{0}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\min\!\Bigl\{1,\tfrac{1}{\sqrt{n\lambda\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})}}\Bigr\}+(\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1)\,\min\!\Bigl\{1,\tfrac{1}{\sqrt{n\lambda(\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1)}}\Bigr\}.$
Since for every $\lambda>0$ we have
$\frac{\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1}{\sqrt{|{\mathcal{I}_{2}}|\lambda(\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1)}}=\frac{\sqrt{\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1}}{\sqrt{|{\mathcal{I}_{2}}|\lambda}}>\frac{\sqrt{\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})}}{\sqrt{|{\mathcal{I}_{2}}|\lambda}}=\frac{\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})}{\sqrt{|{\mathcal{I}_{2}}|\lambda\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})}},$
and of course $\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1>\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})$, we get the inequality
$(\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1)\,\min\!\Bigl\{1,\tfrac{1}{\sqrt{|{\mathcal{I}_{2}}|\lambda(\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1)}}\Bigr\}>\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})\,\min\!\Bigl\{1,\tfrac{1}{\sqrt{|{\mathcal{I}_{2}}|\lambda\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})}}\Bigr\},$
and thus,
$U^{\prime}(\lambda)=$ $\sum_{i\neq i_{0}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi_{i}^{*}(\lambda)+(\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1)\,\min\!\Bigl\{1,\tfrac{1}{\sqrt{|{\mathcal{I}_{2}}|\lambda(\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1)}}\Bigr\}$
$>$ $\sum_{i\neq i_{0}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi_{i}^{*}(\lambda)+\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})\,\min\!\Bigl\{1,\tfrac{1}{\sqrt{|{\mathcal{I}_{2}}|\lambda\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})}}\Bigr\}$
$=$ $U(\lambda)=B.$
Since $U^{\prime}$ is continuous and decreasing on $(0,\infty)$ with $\lim_{\mu\to 0^{+}}U^{\prime}(\mu)=\sum_{i}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})>B$ and $\lim_{\mu\to\infty}U^{\prime}(\mu)=0$, the intermediate‐value theorem guarantees a unique $\lambda^{\prime}$ with $U^{\prime}(\lambda^{\prime})=B$. Moreover, since $U^{\prime}$ is decreasing, and $U^{\prime}(\lambda)\geq B=U^{\prime}(\lambda^{\prime})$ we get $\lambda^{\prime}\geq\lambda$. Define $\pi_{i}$ the optimal solution of (10) with $\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})$ and $\pi^{\prime}_{i}$ the optimal solution with $\hat{f}^{\prime}_{{\tau_{\textup{prior}}}}(X_{i})$.
- If $i\neq i_{0}$, since $\hat{f}^{\prime}_{{\tau_{\textup{prior}}}}(X_{i})=\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})$,
$\pi^{\prime}_{i}=\min\!\Bigl\{1,\tfrac{1}{\sqrt{|{\mathcal{I}_{2}}|\,\lambda^{\prime}\,\hat{f}^{\prime}_{{\tau_{\textup{prior}}}}(X_{i})}}\Bigr\}\;\leq\;\min\!\Bigl\{1,\tfrac{1}{\sqrt{|{\mathcal{I}_{2}}|\,\lambda\,\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})}}\Bigr\}=\pi_{i}.$
- If $i=i_{0}$, then $\hat{f}^{\prime}_{{\tau_{\textup{prior}}}}(X_{i_{0}})=\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1>\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})$, so
$\pi^{\prime}_{i_{0}}=\min\!\Bigl\{1,\tfrac{1}{\sqrt{|{\mathcal{I}_{2}}|\,\lambda^{\prime}\,(\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})+1)}}\Bigr\}\;\leq\;\min\!\Bigl\{1,\tfrac{1}{\sqrt{|{\mathcal{I}_{2}}|\,\lambda^{\prime}\,\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})}}\Bigr\}\;\leq\;\min\!\Bigl\{1,\tfrac{1}{\sqrt{|{\mathcal{I}_{2}}|\,\lambda\,\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})}}\Bigr\}=\pi_{i_{0}}.$
By the inductive assumption, $\pi^{\prime}_{i}\geq B/(|{\mathcal{I}_{2}}|M)$ for all $i\in{\mathcal{I}_{2}}$, hence $\pi_{i}\geq B/(|{\mathcal{I}_{2}}|M)$ for all $i\in{\mathcal{I}_{2}}$ as well. Lastly, by the definition of the weights, we have:
$w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)=({\pi_{i}^{*}})^{-1}\leq\max\left(\frac{|{\mathcal{I}_{2}}|M}{B},1\right).$ (12)
This completes the proof. ∎
D.6 Proofs of the coverage rate guarantees
In this section, we present and prove the formal version Theorem 4.1 which builds on the proof of (Gui et al., 2024, Theorem 3).
Theorem D.1 (General validity, formal).
Fix a tolerance level $\delta\in\left(0,1\right)$ and a miscoverage level $\tau\in\left(0,1\right)$. Suppose that $\{(X_{i},T_{i})\}_{i=1,...,n}$ and $(X_{\textup{test}},T_{\textup{test}})$ are drawn i.i.d., and that the censoring times satisfy the conditional independence assumption (Assumption 2.1) and $(C_{i},T_{i}){\perp\!\!\!\perp}(C_{j},T_{j})|\{X_{k}\}_{k\in{\mathcal{I}_{2}}}$ for all $i\neq j\in{\mathcal{I}_{2}}$. We suppose that $\hat{q}_{\tau}(x)$ is non-decreasing and continuous in $\tau$. Further, assume that the weights are computed using the true probabilities:
$w_{\tau}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)=1/{\mathbb{P}}\bigl[\hat{q}_{\tau}(X_{i})\leq C_{i}\bigm|\{X_{j}\}_{j\in{\mathcal{I}_{2}}}\bigr].$ (13)
Above, $\hat{q}_{\tau}(X_{i})$ can be either the estimated quantile of $T_{i}\mid X=X_{i}$ presented in Section 3, or its trimmed version $\hat{f}_{\tau}(X_{i})$ from Section 5.3. We assume that there exists a constant ${\gamma}_{\tau}>0$ such that the weights satisfy ${w}_{\tau}(x)\leq{\gamma}_{\tau}$ for $P_{X}$-almost all $x$. Consider the following estimated miscoverage rate:
$\hat{\alpha}(\tau)=\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}w_{\tau}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\;\mathbb{I}\bigl\{\tilde{T}_{i}<\hat{q}_{\tau}(X_{i})\leq C_{i}\bigr\}$ (14)
and denote the calibrated quantile level by
$\hat{\tau}=\sup\Bigl\{\tau\in\mathcal{T}:\sup_{\begin{subarray}{c}\tau^{\prime}\in\mathcal{T}\\ \tau^{\prime}\leq\tau\end{subarray}}\hat{\alpha}(\tau^{\prime})\leq{\alpha}\Bigr\}.$ (15)
Above, the search space $\mathcal{T}$ is formulated as in Section 4, or using $\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]$, as in the proposed Adaptive method from Section 5.2. We remark that for $\hat{\tau}$ to be well defined, we assume there exists $\tau^{\prime}\in\mathcal{T}$ such that $\hat{\alpha}(\tau^{\prime})\leq\alpha$ . This assumption can be trivially satisfied by setting $\hat{q}_{0}(X_{i})=0$. Then, with probability at least $1-\delta$ over the draws of $\mathcal{D}$, the LPB $\hat{L}(x)=\hat{q}_{\hat{\tau}}(x)$ satisfies
$\mathbb{P}\left[T_{\textup{test}}\geq\hat{L}(X_{\textup{test}})|\mathcal{D}\right]\geq 1-\alpha-\sup_{\tau\in[0,1]}\left\{\sqrt{\frac{2\gamma^{2}_{\tau}+5}{|{\mathcal{I}_{2}}|}\cdot\log\left(\frac{1}{\delta}\right)}\right\}.$ (16)
Proof.
For ease of notation, we define the coverage gap by:
$\Delta:=\sup_{\tau\in[0,1]}\left\{\sqrt{\frac{2\gamma^{2}_{\lambda}+5}{|{\mathcal{I}_{2}}|}\cdot\log\left(\frac{1}{\delta}\right)}\right\}.$ (17)
We define the oracle miscoverage level by:
$\tau(\alpha+\Delta)=\sup\left\{\lambda\in[0,1]:\mathbb{P}\left(T<\hat{q}_{\lambda}(X)\mid{\mathcal{I}_{1}}\right)\leq\alpha+\Delta\right\}.$ (18)
Observe that by assuming $1-\delta\leq\mathbb{P}(\hat{\tau}\leq\tau(\alpha+\Delta)\mid{\mathcal{I}_{1}})$, we get that the event $\{\hat{\tau}\leq\tau(\alpha+\Delta)\}$ holds with probability at least $1-\delta$. Under the monotonicity of $\hat{q}_{\tau}$, and following the the left-continuity of $\mathbb{P}(T\geq\hat{q}_{\tau}(X)\mid\mathcal{D})$ in $\tau$ we obtain that with probability at least $1-\delta$:
$\begin{split}&\mathbb{P}(T\geq\hat{q}_{\hat{\tau}}(X)\mid\mathcal{D})\\ \geq&\mathbb{P}(T\geq\hat{q}_{\tau(\alpha+\Delta)}(X)\mid\mathcal{D})\\ \geq&1-\alpha-\Delta.\end{split}$ (19)
Now, we turn to show that $1-\delta\leq\mathbb{P}(\hat{\tau}\leq\tau(\alpha+\Delta)\mid{\mathcal{I}_{1}})$. We begin by fixing $\varepsilon>0$, and denoting $\lambda:=\tau(\alpha+\Delta)+\varepsilon$. By the definition of $\hat{\alpha}(\tau)$, we get:
[…]
Next, we apply Markov’s inequality for any $t>0$, and obtain:
[…]
By conditioning on $\{(X_{i},T_{i})\}_{i\in{\mathcal{I}_{2}}}$ and following the $\frac{1}{4}$ sub-gaussianity of $w_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)^{-1}-\mathbb{I}\{\hat{q}_{\lambda}(X_{i})\leq C_{i}\}$, the conditionally independent censoring assumption (Assumption 2.1) and the bounded weights, we have:
$\begin{split}\mathbb{E}&\left(\exp\left(\frac{t}{|{\mathcal{I}_{2}}|}\cdot\sum_{i\in{\mathcal{I}_{2}}}w_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\mathbb{I}\bigl\{T_{i}<\hat{q}_{\lambda}(X_{i})\bigr\}\left(w_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)^{-1}-\mathbb{I}\{\hat{q}_{\lambda}(X_{i})\leq C_{i}\}\right)\right)\bigg|\{(X_{i},T_{i})\}_{i\in{\mathcal{I}_{2}}},{\mathcal{I}_{1}}\bigg.\right)\\ \leq&\exp\left(\frac{t^{2}}{8|{\mathcal{I}_{2}}|^{2}}\cdot\sum_{i\in{\mathcal{I}_{2}}}w_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)^{2}\mathbb{I}\bigl\{T_{i}<\hat{q}_{\lambda}(X_{i})\bigr\}^{2}\right)\\ \leq&\exp\left(\frac{t^{2}}{8|{\mathcal{I}_{2}}|^{2}}\cdot\sum_{i\in{\mathcal{I}_{2}}}w_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)^{2}\right)\\ \leq&\exp\left(\frac{\gamma^{2}_{\lambda}t^{2}}{8|{\mathcal{I}_{2}}|}\right).\end{split}$ (22)
Therefore,
[…]
Further, defining $p_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i):=\mathbb{P}(T_{i}<\hat{q}_{\lambda}(X_{i})\mid\{X_{j}\}_{j\in{\mathcal{I}_{2}}},{\mathcal{I}_{1}})$, we condition on $\{X_{j}\}_{j\in{\mathcal{I}_{2}}}$ and get, using the $\frac{1}{4}$ subgaussianity of $p_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)-\mathbb{I}\bigl\{T_{i}<\hat{q}_{\lambda}(X_{i})\bigr\}$:
$\begin{split}\mathbb{E}&\left(\exp\left(p_{\lambda}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i))-\frac{t}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}(\mathbb{I}\bigl\{T_{i}<\hat{q}_{\lambda}(X_{i})\bigr\}\right)\Bigg|\{X_{j}\}_{j\in{\mathcal{I}_{2}}},{\mathcal{I}_{1}}\bigg.\right)\\ \leq&\exp\left(\frac{t^{2}}{8|{\mathcal{I}_{2}}|^{2}}\cdot|{\mathcal{I}_{2}}|\right)\\ =&\exp\left(\frac{t^{2}}{8|{\mathcal{I}_{2}}|}\right)\end{split}$ (24)
We plug this into (23) to obtain:
[…]
We now apply the Cauchy-Schwarz inequality:
$\begin{split}&\mathbb{E}\left(\exp\left(\frac{t}{|{\mathcal{I}_{2}}|}\cdot\sum_{i\in{\mathcal{I}_{2}}}(\alpha-p_{\lambda}(X_{i}))\right)\bigg|{\mathcal{I}_{1}}\bigg.\right)\\ \leq&\mathbb{E}\left(\exp\left(\frac{2t}{|{\mathcal{I}_{2}}|}\cdot\sum_{i\in{\mathcal{I}_{2}}}(\alpha-p_{\lambda}(X_{i}))\right)\bigg|{\mathcal{I}_{1}}\bigg.\right)^{1/2}\end{split}$ (26)
Observe that by the definition of $\tau(\alpha+\Delta)$ we have:
$\mathbb{P}(T<\hat{q}_{\lambda}(X)\mid{\mathcal{I}_{1}})\geq\alpha+\Delta.$ (27)
By plugging it into the above, we get:
[…]
Above, we used the $\frac{1}{4}$-sub-gaussianity of $\mathbb{P}(T<\hat{q}_{\lambda}(X)\mid{\mathcal{I}_{1}})-p_{\lambda}(X_{i})$. By combining it all, we obtain:
[…]
We define
$t=\frac{(8+2\sqrt{14})|{\mathcal{I}_{2}}|\Delta}{2\gamma^{2}_{\lambda}+5}$ (30)
and plug it into (25):
[…]
That is, we just showed that:
$1-\delta\leq\mathbb{P}(\hat{\alpha}(\tau(\alpha+\Delta)+\varepsilon)>\alpha\mid{\mathcal{I}_{1}})\leq\mathbb{P}(\hat{\tau}<\tau(\alpha+\Delta)+\varepsilon\mid{\mathcal{I}_{1}}).$ (32)
The above equation holds for every $\varepsilon>0$, and thus by taking $\varepsilon\rightarrow 0$, and following the continuity of the probability measure, we obtain that $1-\delta\leq\mathbb{P}(\hat{\tau}\leq\tau(\alpha+\Delta)\mid{\mathcal{I}_{1}})$, as required. ∎
Since the proposed methods construct censoring times that satisfy the conditional independence assumption of Theorem D.1, their validity follow immediately from this theorem with the corresponding definitions of the censoring times and the weights.
Appendix E Experimental details
This section provides additional information about the experimental setup and methodology described in Section 6. Section E.1 details the synthetic experiments, including data generation and model architecture. Section E.2 presents experiments conducted on real-world data. Finally, Section E.3 outlines implementation details that are shared across both experimental settings.
E.1 Synthetic experiments
Synthetic covariate generation
As described in Section 6.1 of the main text, we use synthetic data to evaluate the informativity and coverage of the LPBs constructed by our calibration procedures. Each synthetic data point consists of covariates $X_{i}\in\mathbb{R}^{d}$ with $d=10$, and a prompt-specific unsafe-output probability $p_{i}$. The dataset is designed to simulate a setting in which most prompts have a high likelihood of producing unsafe outputs, while a smaller subset is relatively safe. Notably, our experiments show that fitting a model on this synthetic dataset results in a large prediction error, which turns the calibration step more challenging.
To this end, we sample the covariates from a Gaussian distribution:
$X_{i}\mid p_{i}\sim\mathcal{N}\left(\mu(p_{i}),\sigma^{2}I_{d}\right),\quad\text{with }\sigma=0.1.$ (33)
where the mean vector $\mu(p_{i})\in\mathbb{R}^{d}$ is designed as follows. First, we construct a pool of $p_{i}$ values, where (i) 90% of which are drawn uniformly in the range $\log_{10}(p_{i})\sim\mathcal{U}[-4,-3]$, and (ii) the remaining 10% are drawn uniformly in the range $\log_{10}(p_{i})\sim\mathcal{U}[-6,-5]$. We then assign each $p_{i}$ to a data point by sampling without replacement from this pool. Then, we define $d$ quantile levels as follows
$\tau_{j}=0.1+0.8\,\frac{j-1}{d-1}\ \ \ j=1,\ldots,d.$
Next, for a given probability $p_{i}$, we compute:
$\tilde{\mu}(p_{i})=\bigl[F^{-1}_{\mathrm{Geom}(p_{i})}(\tau_{1})^{1/4},\,\ldots,\,F^{-1}_{\mathrm{Geom}(p_{i})}(\tau_{d})^{1/4}\bigr]^{\top},$
where $F^{-1}_{\mathrm{Geom}(p_{i})}$ is the quantile function of the Geometric distribution with success probability $p_{i}$. The $1/4$ power transformation reduces the size of the often extremely high raw quantiles. Lastly, to improve numerical stability, we normalize this vector by the average magnitude across all training examples:
$\bar{\mu}=\frac{1}{n\,d}\sum_{i=1}^{n}\sum_{j=1}^{d}F^{-1}_{\mathrm{Geom}(p_{i})}(\tau_{j})^{1/4},\quad\mu(p_{i})=\frac{\tilde{\mu}(p_{i})}{\bar{\mu}}.$
Model architecture and training
To estimate the unsafe generation probabilities, we train a neural network with four hidden layers of size 32 each. We used ReLU for non-linearity and a sigmoid for output. We optimize the BCE loss (see Section E.3) using AdamW with learning rate $10^{-4}$, weight decay $10^{-5}$, batch size $100$, for $10$ epochs.
Calibration method parameters
In all experiments, unless stated otherwise, we used the prior quantile $\tau_{\mathrm{prior}}=10^{-1/4}$, with trimming threshold $M$ set so that $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right)=10$.
E.1.1 Additional experiments
To confirm that our findings on synthetic data do not depend on a particular choice of calibration/test split, we repeated the experiment from Section 6.1 for 20 independent splits of calibration and test sets. Specifically, in each trial, we fixed the original training set of 45,000 prompts, then randomly partitioned the remaining pool into a calibration set of size 45000 and a test set of size 10,000. All other settings follow the experimental setup described in Section E.1.
We repeated the experiment for the following values of average sampling budget per prompt $B/|I_{2}|$: 10,25,50,100,200,300,600,1200{10,25,50,100,200,300,600,1200}. In all experiments, we used the same prior quantile $\tau_{\mathrm{prior}}=10^{-1/4}$, with trimming threshold $M$ set so that $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right)=10$, as in Section 6.1.
Figure 3 presents the performance of the methods introduced in this work for 20 data splits. Across every budget level, the performances closely match those from our original fixed‐split experiment from Figure 1 in Section 6.1.
Figure 3: Results of synthetic experiments as a function of average budget per prompt $B/|{\mathcal{I}_{2}}|$. Left: Coverage (target 90%; gray dashed line). Center: Mean number of samplings generated per prompt. Right: Mean LPB. Shaded regions represent the standard deviation over 20 runs.
Next, we evaluate how the target coverage level affects the constructed LPBs and their informativeness by repeating the experiment from Section 6.1 across different nominal coverage levels. We present the results in Figure 4. This figure shows that the Optimized consistently achieves its target coverage across all levels.
Figure 4: Results of synthetic experiments as a function of average budget per prompt $B/|{\mathcal{I}_{2}}|$. Top: Coverage (target level indicated by a gray dashed line). Bottom: Mean LPB. Shaded regions represent the standard deviation over 20 runs.
Finally, we turn to study the effect of the maximal weight size $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right)$ on the stability and power of the Trimmed and Optimized calibration methods. Figure 5 compares the two methods under a fixed per‐prompt budget of ${B}/{|{\mathcal{I}_{2}}|}=1000$. For each weight size, we set the threshold $M={B\gamma}/{|{\mathcal{I}_{2}}|}$. As can be seen in this figure, the size of the LPBs we construct increases with $\gamma$. However, this comes at the cost of increasing the coverage variance. Observe how the variance of Optimized method for $\gamma=100$ is similar to that of Trimmed method for $\gamma=10$, but the former yields much more informative LPBs under these corresponding $\gamma$ values. This is a result of the variance-reducing optimization objective in (10).
Figure 5: Results of synthetic experiments as a function of the maximum weight $w_{\text{max}}$ of the Trimmed and Optimized methods. Left: Coverage (target 90%; gray dashed line). Right: Average LPB. Shaded regions present the standard deviation across $20$ runs.
We compare the proposed Optimized method with an oracle quantile regression model on our synthetic dataset. The oracle model outputs the true conditional quantile of $T\mid X$ as the LPB, and therefore achieves the target coverage rate of $1-\alpha=90\%$ by definition. These oracle LPBs are also the most informative possible by construction. Figure 6 reports the performance of the Uncalibrated quantile regression model, Naive, Basic, Trimmed, Optimized, and the oracle quantile regression model. As expected, the oracle model achieves the nominal coverage level. Notably, our Optimized method produces LPBs of comparable size to the oracle as the budget increases, demonstrating its efficiency in approaching the best achievable performance even without access to the oracle quantiles.
Figure 6: Synthetic data experiment with an oracle quantile regression model. Top: Coverage (target $90\%$%; gray dashed line). Bottom: Mean LPB. Shaded regions represent the standard deviation over 20 runs.
We compare our approach with a calibration method that has an infinite budget on our synthetic data. With an unlimited budget, we can observe an unsafe event for every calibration sample and thus can employ classic conformal prediction (Vovk et al., 2005, Papadopoulos, 2008) without data re-weighting to obtain LPBs that achieve $1-\alpha=90\%$ coverage rate. Specifically, we estimate the miscoverage rate similarly to (6), but here, we use all calibration points:
$\hat{\alpha}^{\text{ib}}(\tau)=\frac{1}{|{\mathcal{I}_{2}}|+1}\sum_{i\in\mathcal{I}_{2}}\mathbb{I}\bigl\{T_{i}<\hat{q}_{\tau}(X_{i})\bigr\}$ (34)
and construct the LPB as in (7) with the above $\hat{\alpha}^{\text{ib}}$.
We summarize the performance of all methods in Figure 7. The results reveal that the infinite budget method achieves the desired coverage rate with lower variability than the alternatives. Moreover, our Optimized approach produces LPBs comparable in size to those of the infinite-budget calibration as the budget increases, demonstrating its effectiveness in approaching optimal performance even under budget constraints.
Figure 7: Synthetic data experiment with a calibration method that has an infinite budget. Top: Coverage (target 90%; gray dashed line). Bottom: Mean LPB. Shaded regions represent the standard deviation over 20 runs.
E.2 Real-Data Experiments
Calibration method parameters
In all experiments, unless stated otherwise, we used the prior quantile $\tau_{\mathrm{prior}}=10^{-1/4}$, with trimming threshold $M$ set so that $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right)=2$.
Efficient prompt‐parallel sampling
To efficiently generate outputs during our calibration procedures, we implement a prompt-parallel generation algorithm on top of the vLLM package (Kwon et al., 2023). In each iteration, we gather a batch of $1500$ prompts $X_{i}$} along with their corresponding censoring budgets $C_{i}$}, and distribute them across multiple GPUs. The pipeline iteratively performs the following steps in parallel for a batch of prompts, until all prompts in the calibration set are processed:
  1. 1.
    Generate an output $\mathcal{G}(X_{i})$ using the LLM.
  2. 2.
    Apply the audit function $\mathrm{Audit}\bigl(X_{i},\mathcal{G}(X_{i})\bigr)$ to identify unsafe outputs.
  3. 3.
    Check whether $X_{i}$ has either: (i) reached the $C_{i}$ generations limit; or (ii) produced an output that fails the audit. If at least one of these two conditions is met, remove the prompt from the pipeline and replace it with a new prompt.
By promptly removing any prompt that yields an unsafe output, the processing maintains high throughput and avoids idle GPU time.
Runtime evaluation
Table 1 presents average runtime for the Naive and Optimized calibration procedures across varying prompt budgets. While the Naive method can have a long runtime due to its unbounded censorship times $C_{i}$, it is observed to be faster than the Optimized method; see Table 1. We explain this by two key factors. First, our implementation of the Naive method includes an efficient runtime optimization: instead of always generating all $C_{i}$ samples, we terminate generation early when additional samples no longer improve the estimate of $\hat{\alpha}(\tau)$. As detailed in Section E.3.1, this strategy yields LPBs that are equivalent to those produced by the original Naive procedure, while significantly reducing computation time. Second, the base model used in these experiments tends to produce low predicted quantiles $\hat{q}_{\tau}(X_{i})$, as shown in Figure 2. Consequently, the effective censorship budgets $C_{i}$ are also low for most prompts, further limiting runtime.
In sum, these two factors—runtime-efficient implementation and tendency to produce low predicted quantiles—explain why the reported runtime of the Naive is lower than the Optimized method in this setting.
Budget per prompt Naive runtime (hours) Optimized runtime (hours) # GPUs
10 0.032 $\pm$ 0.005 0.063 $\pm$ 0.001 6
25 0.042 $\pm$ 0.001 0.117 $\pm$ 0.001 6
50 0.070 $\pm$ 0.006 0.209 $\pm$ 0.001 6
100 0.124 $\pm$ 0.003 0.388 $\pm$ 0.003 6
200 0.215 $\pm$ 0.010 0.589 $\pm$ 0.007 6
300 0.316 $\pm$ 0.012 0.718 $\pm$ 0.002 6
600 0.561 $\pm$ 0.016 0.979 $\pm$ 0.006 6
1200 0.918 $\pm$ 0.051 1.792 $\pm$ 0.041 4
Table 1: Average runtime (hours) by method and budget per prompt (mean $\pm$ SD).
Model architecture and training
We employ the ModernBERT-base (Warner et al., 2024) model. We use Adam optimizer (Kingma & Ba, 2015) to minimize the loss function in (35), with a learning rate of 3e-5, and a batch size of $1500$. We train the model for 10 epochs and use early stopping based on the validation set.
E.3 Shared implementation details for both real and synthetic experiments
Our calibration framework applies to any machine learning model that outputs an estimated conditional quantile of the time-to-unsafe-sampling, denoted by $\hat{q}_{\tau}(x)$, given the covariates $x$. This is motivated by the fact that $T$, the time-to-unsafe-sampling, is a Geometric random variable with parameter $p(x)$, the probability of unsafe generation. Consequently, the conditional quantile function of $T$ is given analytically in terms of $p(x)$, and vice versa:
$q_{\tau}(x)=\left\lceil\frac{\log(1-\tau)}{\log(1-p(x))}\right\rceil\quad\text{and}\quad p(x)=1-(1-\tau)^{1/q_{\tau}(x)}.$
The above relationship allows us to estimate either $p(x)$ or $q_{\tau}(x)$ and recover the other via a closed-form transformation.
Loss function
In our experiments, we estimate $p(x)$ by fitting a model that minimizes the binary cross-entropy (BCE) loss over aggregate unsafe proportions. Specifically, for each prompt $X_{i}$, we define the empirical success rate as $\bar{Y}_{i}=\frac{1}{N}\sum_{j=1}^{N}Y_{i}^{j}$, where $Y_{i}^{j}\in\{0,1\}$ indicates whether the $j$-th sample is unsafe. We then use the BCE loss:
$\mathrm{BCE}(\bar{Y}_{i},\hat{p}(X_{i}))=-\bigl[\bar{Y}_{i}\log\hat{p}(X_{i})+(1-\bar{Y}_{i})\log(1-\hat{p}(X_{i}))\bigr],$ (35)
where $\hat{p}(X_{i})$ is the model’s prediction of the probability of unsafe generation. Notably, this loss is equivalent to the mean of the standard BCE loss across individual samples:
$\frac{1}{N}\sum_{j=1}^{N}\mathrm{BCE}(Y_{i}^{j},\hat{p}(X_{i}))=\mathrm{BCE}(\bar{Y}_{i},\hat{p}(X_{i})).$
This aggregation allows for a more computationally and memory-efficient implementation compared to the use of the vanilla BCE.
E.3.1 Implementation details of the Naive calibration procedure
In theory, the Naive calibration (Algorithm 3) can assign an unbounded censorship time per prompt, which leads to arbitrarily long running times. To avoid this, we restrict our search for the threshold $\tau$ to the compact set $\mathcal{T}\cap\tau_{\mathrm{prior}}$ (see Section 6.2). This restriction in search space does not violate the method’s validity, as it may only make it more conservative, as discussed in Section 5.2. This section also shows that assigning a censorship time $C_{i}\geq\hat{q}_{\tau}(X_{i})$ is equivalent to drawing exactly $\hat{q}_{\tau}(X_{i})$ generations. Conversely, setting $C_{i}<\hat{q}_{\tau}(X_{i})$ amounts to not drawing any samples. Based on this observation, we implement a run-time efficient version of the Naive method that draws censorship times according to
$C_{i}:=\text{Ber}(g(X_{i}))\cdot\hat{q}_{\tau_{\text{prior}}}(X_{i}),$
where $g(x)=\mathbb{P}(\mathrm{Geom}(\min(|{\mathcal{I}_{2}}|/B,1)\geq\hat{q}_{\tau_{\text{prior}}}(X_{i}))$. It is important to note at this point that the average number of generations per prompt reported in Figure 1 is the simulated number of samples that the equivalent, non-optimized Naive solution would use. This implementation also implies that the Naive method has a much smaller overall number of generations, leading to a faster runtime.
E.4 Machine’s spec
The computational infrastructure used in the experiments are:
  • CPU: Intel(R) Xeon(R) CPU E5-2683 v4 @ 2.10GHz, Intel(R) Xeon(R) Gold 5318Y CPU @ 2.10GHz, Intel(R) Xeon(R) Gold 6336Y CPU @ 2.40GHz.
  • GPU: NVIDIA A40, NVIDIA TITAN X (Pascal), NVIDIA 2080 TI, NVIDIA RTX 2060 SUPER.
  • OS: Ubuntu 20.04.6.